### v1

# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

Abstract.
Performance plays a critical role in ensuring the smooth operation of any mobile application, directly influencing user engagement and retention. Android applications are no exception. However, unlike functionality issues, performance issues are more challenging to discover as their root causes are sophisticated and typically emerge under specific payloads. To tackle this problem, researchers have dedicated substantial efforts to proposing automatic approaches for understanding, detecting, and resolving performance issues. Despite these endeavors, it still remains unknown what the status quo of Android performance analysis is, and whether existing approaches can indeed accurately reflect real performance issues. To fill this research gap, we conducted a systematic literature review followed by an explanatory study to explore relevant studies and real-world challenges. Our findings reveal that current tools have limited capabilities, covering only 17.50% of the performance issues. Additionally, existing datasets encompass only 27.50% of the issues and are very limited in size. We also show real-world issue patterns, underscoring the huge gap between the identified techniques and practical concerns. Furthermore, possible solutions are provided to guide future research towards achieving effective performance issue detection and resolution.
conference: ; ; copyright: none
1. Introduction
With the increasing popularity of smartphones, the Android system has become one of the largest mobile platforms in the world, holding a 70.69% market share worldwide (iPhoneAndroidStatistics, ). The exponential growth of Android usage has seen an unprecedented expansion within the mobile app industry, with millions competing for users’ attention. In such a fiercely competitive circumstance, developers are under pressure to create new, fancy apps in the quickest manner. However, this emphasis on rapid development can sometimes come at the expense of app quality, particularly in terms of performance.
Unlike functionality issues that directly cause app malfunctions or failures, performance issues occur when the app functions correctly but its efficiency is low, resource usage is unreasonable, or user interaction is poor, significantly reducing user satisfaction and software reliability (Li2019DetectingAD, ; Liu2014CharacterizingAD, ; Gao2017EveryPC, ; Liu2018NavyDroidAE, ). These issues include, but are not limited to, excessive energy consumption (Banerjee2018EnergyPatchRR, ; Pathak2012WhatIK, ; Carette2017InvestigatingTE, ), resource leaks (Wu2016LightWeightIA, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Yan2013SystematicTF, ), and responsiveness delays (Farooq2018RuntimeDroidRR, ; Kang2016DiagDroidAP, ; Lin2024AgingOG, ). For example, Liu et al. (Liu2014CharacterizingAD, ) demonstrate that wasteful energy consumption can occur in background threads that continue to update invisible GUIs or perform complex network computations, rapidly depleting battery life and degrading app performance. Similarly, Yan et al. (Yan2013SystematicTF, ) reveal resource leak defects associated with events such as screen rotation, app switching, pressing the back button, and file or database access.
However, due to the non-functional nature of performance issues, their problems do not immediately manifest. Additionally, performance issues can be caused by multiple factors, requiring in-depth analysis and debugging to diagnose and identify the root causes (Liu2016FixingRL, ; Habchi2018OnAL, ; Jabbarvand2019SearchBasedET, ; Li2019CharacterizingAD, ; Liu2014GreenDroidAD, ). Furthermore, some performance issues only appear under specific usage scenarios or loads (Behrouz2016EnergyawareTM, ; Jabbarvand2019SearchBasedET, ; Yan2013SystematicTF, ; Farooq2018RuntimeDroidRR, ). Consequently, manually identifying performance issues is time-consuming and labor-intensive, making it difficult for developers to address all problems during the testing phase. Therefore, it is crucial to use effective automated tools to analyze, detect, and fix performance issues (Habchi2018OnAL, ; Li2019DetectingAD, ; Li2022CombattingEI, ).
Unfortunately, at the current stage, researchers face difficulties in designing automated tools to comprehensively address Android performance issues due to: 1) Lack of clear classification of performance issues and their root causes, making it hard to fully understand the performance issues; 2) Uncertainty about the state quo of existing approaches, making it unclear to what extent these tools are reliable for detecting or fixing performance issues; 3) The absence of benchmark datasets in evaluating state-of-the-art tools; 4) Uncertainty about whether state-of-the-art tools and datasets accurately reflect real performance issues.
To guide future research, we first conducted a systematic literature review to provide a holistic taxonomy for existing performance issues and their underlying root causes. We then identified research gaps between state-of-the-art tools and datasets relative to the targeted research objectives (i.e., performance issues). Subsequently, we analyzed real-world performance issues to pinpoint differences between existing literature and practical challenges. Through these investigations, our goal is to offer a thorough understanding of Android performance issues and to guide the development of more effective automated tools and comprehensive datasets for evaluation. In this paper, we aim to explore the following research questions:
RQ1: What is the status quo of Android performance issues analysis in research field?
  • What performance issues are being targeted by researchers? The lack of a universally accepted taxonomy leads to many papers targeting the same performance problems using different terms. Additionally, since one factor can lead to multiple performance issues and one issue can be caused by multiple factors, creating a clear classification system is challenging. Our work filled this gap by distinguishing between the consequences and factors of performance issues. We identified five categories of Android performance issues and 21 distinct root causes.
  • What approaches have been developed for performance issues? To solve the above performance issues, researchers have proposed various kinds of techniques. Thus, in the second research question, we aim to understand the status quo of current techniques, seeking to comprehensively grasp the research gap between performance issues and automatic tools. In total, we classified them into three primary techniques: static-based, dynamic-based, and deep learning-based approaches. Static analysis is the most frequently adopted technique, while deep learning-based methods are much less commonly utilized for performance issues. Moreover, it is surprising to find that only 7 out of 21 factors are targeted by these tools. This evidence confirms a significant gap between performance issues and automated resolutions. Additionally, the majority of these tools focus on issue detection rather than fixing, highlighting another research gap concerning the need for proactive problem-solving.
  • What datasets have been used in Android performance analysis? High-quality datasets are essential for evaluating performance analysis tools. We found that current datasets cover only 11 out of 21 factors, addressing a limited range of issues. Furthermore, the number of existing datasets is very small. This highlights an urgent need for a comprehensive public dataset.
RQ2: What are gaps between cutting-edge studies and real-world scenarios?
  • What real-world performance issues in the wild? Given that performance is an inherently practical concern, we proceeded to conduct a deeper analysis of real-world issues to pinpoint differences between existing literature and practical challenges. Our examination revealed one previously unrecognized performance issue and 19 previously unacknowledged factors across real-world platforms. This suggests that numerous factors contributing to performance issues in practical settings have not received attention in research. Furthermore, we uncovered a misalignment between the primary concerns of researchers, developers, and users. These discoveries underscore a substantial gap in fully comprehending and mitigating performance issues in real-world scenarios.
  • What are the common patterns exhibited in real-world performance issues? Based on our systematic analysis for real-world data, we identified six common code patterns: API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns. These newly discovered patterns provide insights for designing automated tools to address performance issues and for creating datasets to evaluate the effectiveness of these tools.
2. Studies on Performance Issues
In this work, we are interested in how far we are towards automating performance issues analysis in Android apps. Thus, in this section, we first resort to a systematic literature review (SLR) to understand the status quo of Android app performance analysis.
2.1. Literature Review
To conduct systematic literature review, we followed the methodologies outlined by Brereton et al.  (brereton2007lessons, ) and Kitchenham et al.  (Kitchenham2004ProceduresFP, ). Fig. 1 illustrates the working process which comprises 5 steps.
2.1.1. Keywords Identification
To investigate the status quo of Android performance analysis, we identified a bunch of search keywords to locate relevant publications on this topic. Specifically, the keywords comprise two groups (i.e., G1 AND G2). Each group is made up of several keywords. The search string is then constructed by combining them, i.e., g1 AND g2, where g1 and g2 represent the disjunction of keywords from groups G1 and G2, respectively. Please note that these keywords are based on our initial understanding of performance issues and are intended to help us identify “seed” papers. As we review these papers, we could uncover new performance issues, prompting a snowballing process that could lead us to further relevant studies. The * symbol represents any string of characters. For example, *phone* can refer to smartphone, phones, or phone. Similarly, issue* can represent issues or issue. Keywords are as follows:
G1: android, mobile, *phone*
G2: performance, resource, energy, responsiveness, issue*
Figure 1. The working process of our systematic literature review.
Table 1. CORE A/A* ranked software engineering and mobile venues.
Type Source Research Codes Venues
Journals CORE2020 0803, 0805 TSE, TOSEM, JSS, IST, TMC
Conferences CORE2023 4606, 4612 MOBICOM, Mobisys, EuroSys, MOBIHOC, MSWIM, ASE, ESEC/FSE, ICSE, ISCA, EASE, ECSA, ESEM, FOSSACS, ICSA, ICSME, ICST, ISSRE, ISSTA, MSR, SANER, SEAMS
2.1.2. Repository Search
To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library.
From this initial search, we retrieved a total of 10,766 papers. However, not all of them were of high quality, and many were outside our area of focus. To narrow down to the most relevant papers, we concentrated on the fields of software engineering and mobile, selecting research codes 4606 and 4612 for conferences, and 0805 and 0803 for journals, specifically targeting publications containing the keywords “software” and “mobile”. In addition, to further guarantee the quality of papers, we exclusively chose those published from conferences rated CORE A*/A  [^1] CORE provides conference and journal ratings. https://www.core.edu.au/home. We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the title of relevant publications. As a result, we successfully identified 274 publications distributed across 22 venues as listed in Table 1 (i.e., there is no relevant paper identified in 4 of the venues).
2.1.3. Paper Exclusion
To gather relevant papers as possible, we initially considered all returned results from our search. However, not every paper pertains to Android performance issues. To refine our collection, we conducted a thorough review of the abstracts (and full content when necessary) of these papers, aiming to retain only those closely related to our research focus. Specifically, we applied the following exclusion criteria to streamline this process:
  1. (1)
    Short papers (i.e., less than six pages in double-column format or 11 pages in single-column format) were excluded.
  2. (2)
    Papers targeting on non-Android mobile devices (e.g., Windows Phone, IOS devices) were excluded.
  3. (3)
    Papers using Android performance issues as metrics/indicators for comparison were excluded. For example, Mantis (Kwon2013MantisAP, ) predicts app execution time based on app implementations.
  4. (4)
    Papers targeting on Android performance issues but not focused on the mobile app level were excluded. For example, Hypnos (Jindal2013HypnosUA, ) is used for detecting sleep conflicts in smartphone device drivers, which is outside our scope focused on mobile app issues stemming from poor coding practices by developers.
After applying these exclusion criteria, we retained 27 papers that are closely related to the topic of performance issues analysis at Android mobile level.
2.1.4. Back/Forward Snowballing
After initially identifying relevant papers (i.e., 27 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature. The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources. Additionally, we performed the forward snowballing method by examining papers citing the identified papers. We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified. Eventually, we were able to locate a total of 48 papers. Please note that during snowballing, we also collected 82 keywords related to performance issues for real-world data collections in Section 3.1.2.
2.1.5. Cross Check
Even with the relevant selection criteria, it is not guaranteed that the quality of all identified papers is desirable. Therefore, we evaluated the quality of each paper by reading its full text. Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies. Ultimately, we collected 48 papers to answer our research questions defined in Section 1.
2.2. Result
In the following paragraphs of Section 2, we summarize the state quo from thorough reading of these papers and attempt to understand: 1) Targeted performance issues, which involve key factors that lead to problems needing resolution, 2) Automatic techniques designed to address mobile performance issues, and 3) Datasets designed to evaluate the effectiveness of these automatic techniques.
Table 2. The top problems targeted by the examined papers.
Category Factor Problem Example
Energy Consumption Database Operation Frequently triggering database queries for auto-complete (Li2019DetectingAD, ; Lyu2018RemoveRF, ; Li2022CombattingEI, ; Lyu2017AnES, )
Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, )
Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, )
Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, )
HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, )
Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, )
Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Image Decoding identical image in multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, )
Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, )
Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
UI Rendering Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, )
View Drawing all items in a ListView without reusing exising data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
Responsiveness Obsolete Task Downloading obsolete content occurs when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Image 1) Decoding large images in the UI thread, 2) Loading large image, 3) dynamically adjusting content during activity-switching events (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
SQL Statement Retrieving more columns/rows than necessary in SQL queries  (Lyu2019QuantifyingTP, )
Message Backloging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
Main Thread Performing Heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, )
Activity Loading new resources during activity restarts may freeze the UI  (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, )
Memory Consumption Obsolete Task Retaining obsolete parameters after updating view layouts  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Service Failing to destroy a dead service  (Song2019ServDroidDS, )
Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
AsyncTask Failing to destroy AsyncTask when it holds active GUI components (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, )
HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
Data Loss Activity Failing to save and restore GUI data during activity restarts from runtime changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
AsyncTask Updating destroyed GUI components when it has been recreated using AsyncTask (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
Distorted UI Display Activity Overlapping fragments occur when a new activity is restarted with a new fragment attached while the old fragment is still retained by the system (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
2.2.1. Result - Targeted Problem
We now highlight the representative problems targeted by our fellow researchers (cf. Table 2). Based on the literature review results, we manually summarized and identified five major categories of Android mobile performance issues: Energy Consumption, Memory Consumption, Responsiveness, Data Loss, and Distorted UI Display. Then we reviewed the root causes (a.k.a factors) for each category thoroughly and demonstrate them in Table 2 along with the specific performance problems our researcher fellows are working on.
Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic. The lack of a universally accepted taxonomy results in many papers targeting the same performance problems but using different descriptive terms. Additionally, since one factor can lead to multiple performance issues (consequences), and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system. Therefore, our work addresses this gap by distinguishing between the consequences and factors of performance issues. Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors. To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner. We now elaborate on the five major categories, each with the corresponding factors along with an example presented in Table 2.
Energy Consumption. As shown in Fig. 2, energy consumption receives the most research attention, with 34 out of the examined papers (70.83%) and its increasing trend is expected to continue. This result is not surprising given that the most urgent task at the moment is to ensure that the app remains user-friendly and sustainable in the rapidly evolving app market, which is why many previous studies focused on this research topic. Specifically, we found that most of the energy consumption issues stem from a common pattern: frequent energy-consuming operations occurring together with loops/lifecycle functions. For example, Lyu et al. (Lyu2018RemoveRF, ) found that database writes within loops trigger inefficient autocommit behaviors, leading to significant energy waste. Liu et al. (Liu2014CharacterizingAD, ) discovered that even when the GUI is in the background and not visible, threads continue updating the UI, causing unnecessary energy waste. In addition, invoking APIs for requiring resources (e.g., wakelock, sensor) when they are not needed can also results in a substantial increase in energy consumption (Liu2014GreenDroidAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Banerjee2018EnergyPatchRR, ).
Responsiveness. Another important aspect of Android performance issues is responsiveness. Fig. 2 shows that 18 reviewed articles (37.50%) are crafted for poor responsiveness, leading to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc. Specifically, we identified 10 distinct factors that contribute to this category. Most primary studies (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ) discovered that heavy operations associated with UI threads can slow responsiveness. Additionally, Gao et al. (Gao2017EveryPC, ) and Sarra et al. (Habchi2019TheRO, ) show that improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
Memory Consumption. There are 14 papers (29.17%) related to memory consumption issues in the collected studies. These memory consumption issues have diverse factors, often stemming from inefficient coding that leads to excessive memory use. For example, decoding images without resizing them results in extensive memory consumption (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ). Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks (Lin2015StudyAR, ; Pan2020StaticAC, ).
Data Loss With regard to the data loss problem, Fig. 2 indicates that 7 papers (14.58%) attempt to analyze data loss issues, which significantly degrade user experience when data fails to be saved or restored in UI threads. For instance, studies have revealed that failing to save and restore GUI data during activity restarts can cause existing data to vanish instantly when the screen orientation changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ).
Distorted UI Display. Another aspect of performance issues relates to distorted UI display, which consists of only 2.08% (i.e., 1 study) of the collected studies. Specifically, Umar et al. (Farooq2018RuntimeDroidRR, ) shows that overlapping fragments occur when initializing a new fragment while retaining the old one, causing a poor user experience.
Figure 2. Annual publications on each performance issue.
Discussion. Among all the collected studies on performance issues for Android, it appears that energy consumption is the most frequently studied topic, followed by responsiveness analysis. Additionally, it is noteworthy that an increasing number of recent studies have started to address specific issues related to memory consumption and data loss (e.g., runtime change handling (Farooq2018RuntimeDroidRR, ; Chen2023TransparentRC, )), although the total number of relevant studies remains small. Moreover, the amount of literature focusing on distorted UI display is relatively small, indicating a need for further in-depth research. While this review categorizes the existing literature’s research objectives into five categories, the scope of Android performance issues can be much broader. Therefore, future research should go beyond these categories and propose new approaches tailored to emerging app scenarios.
Answers to RQ1.1: Current research studies focus on five categories of performance issues, prioritized by their prevalence: Energy Consumption (70.83%), Responsiveness (37.50%), Memory Consumption (29.17%), Data Loss (14.58%), and Distorted UI Display (2.08%). We provide a holistic taxonomy of Android performance issues, encompassing 5 major categories and 21 distinct root causes. Performance issues are sophisiticated, as a single issue may stem from multiple factors and a single factor can cause multiple issues.
2.2.2. Result - Technique
To address the above performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches. Table 3 presents the the list of available automatic approaches collected from our examined papers. Despite numerous claims of proposed automated tools, only a small fraction (i.e., 10 out of the 48 examined papers (20.83%)) works have made their tool publicly available. This indicates a significant gap between automatic tools with their practical usage to within our community. In addition, among the identified factors leading to performance issues, only 7 out of 21 are targeted by existing tools. The fact that the majority of the factors are overlooked by SOTA tools indicates a huge gap in addressing performance issues. Moreover, most automatic tools are designed with a primary focus on simply detecting performance issues rather than actively fixing them. This emphasizes a need for a shift towards developing tools that not only identify but also offer solutions to enhance performance. To better understand how these available tools are implemented, we now summarize the techniques adopted in these approaches.
Table 3. A summary of automatic approaches used for Android mobile performance analysis. Static denotes Static Analysis, Dynamic denotes Dynamic Analysis, and DL denotes Deep Learning methods.
Tool/Reference Approach Technique Stage Target Factor Link
DDLDroid (Zhou2023DDLDroidED, ) Static Reverse Engineering Taint analysis Detect Activity  (DDLDroidTool, )
DLD (Riganelli2020DataLD, ) Dynamic Runtime Monitoring Model-based Testing Detect Activity  (Riganelli2020DataLDTool, )
iFixDataloss (Guo2022DetectingAF, ) Static Dynamic Fuzzing Search-based Testing Static Program Repair Detect Fix Activity  (Guo2022DetectingAFTool, )
RCHDroid (Chen2023TransparentRC, ) Dynamic Runtime Management Garbage Collection Fix Activity  (Chen2023TransparentRCTool, )
VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Activity Resource  (Lu2022DetectingRUTool, )
ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Resource  (Jabbarvand2020AutomatedCOTool, )
AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect AsyncTask  (Pan2020StaticACTool, )
TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Image  (Li2019CharacterizingADTool, )
Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Database Loop  (Lyu2018RemoveRFTool, )
ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural- Code Analysis Detect Service  (Song2019ServDroidDSTool, )
Static Analysis. Static approaches are the art of reasoning about the behavior of programs without actually running them. It has been widely used (70.00%) to either detect or fix performance issues, as shown in Table 3. Specifically, taint analysis, symbolic execution, code pattern analysis and data flow analysis are the most commonly adopted techniques for rule-based static analysis as many performance issues arise from common code patterns. Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ). Moreover, they can also provide fundamental information for dynamic fuzzing tools to trigger performance bugs on the fly (Guo2022DetectingAF, ). Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Guo2022DetectingAF, ).
Dynamic Analysis. There are only four (i.e. 40.00%) primary studies that use dynamic analysis approaches to detect or fix performance bugs (especially for data loss issues). This result is align with expectation because dynamic analysis involves executing apps in a controlled environment and unable to cover all possible paths (dong2020timemachine, ). However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ). Table 3 shows that three dynamic tools are designed to address data loss issues in the context of activity restarts associated with sophisticated lifecycle methods, which are challenging to be detected statically. We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
Deep Learning. With the increasing popularity of deep learning (DL) techniques, we are surprised to find that only one paper adopted DL techniques for performance analysis. Specifically, Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) observed runtime features dynamically and leveraged DL models to locate energy anomalies. Although the number of DL-related studies is relatively small, DL techniques have been widely employed in various software engineering tasks, such as code analysis (Allamanis2017LearningTR, ), bug detection (Pradel2018DeepBugsAL, ), and program repair (zhao4254659autopatch, ). It is promising to see a garnered considerable research attention to employ ML/DL techniques to thwart Android performance issues.
Answers to RQ1.2: The majority of existing approaches primarily utilize static analysis (70.00%), with dynamic analysis (40.00%) coming next, and deep learning (10.00%). Only 10 out of the examined papers (20.83%) offer publicly accessible automated tools, highlighting a significant accessibility gap. Only 7 factors (33.33%) are addressed by existing automated tools, indicating incomplete coverage and potential oversight. Most tools prioritize issue detection over resolution, indicating a need for proactive problem-solving.
Table 4. A list of datasets used for Android mobile performance analysis.
Reference Year # APP Type Target Factor Link
Guo et al. (Guo2022DetectingAF, ) 2022 66 Real-World Activity  (Guo2022DetectingAFTool, )
Chen et al. (Chen2023TransparentRC, ) 2023 27 Real-World Activity  (Chen2023TransparentRCTool, )
Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Activity Resource  (Lu2022DetectingRUTool, )
Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) 2020 9 Micro Resource  (Jabbarvand2020AutomatedCOTool, )
AsyncBench (Pan2020StaticAC, ) 2020 69 Micro AsyncTask  (Pan2020StaticACTool, )
Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Image  (Li2019CharacterizingADTool, )
Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Database Operations Loop  (Lyu2018RemoveRFTool, )
Song et al. (Song2019ServDroidDS, ) 2019 45 Real-World Service  (Song2019ServDroidDSTool, )
Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World UI Rendering View Inefficient Method Hashmap  (Habchi2019TheROData, )
2.2.3. Result - Dataset
To evaluate the aforementioned automatic tools, researchers have proposed several datasets targeting unique performance factors. After examining all the public links or websites within the 48 identified studies, we were surprised to find only 9 (18.75%) unique open-sourced datasets, as shown in Table 4. Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison. Additionally, the number of factors covered by existing datasets is quite limited, w.r.t 11 factors in total. The shortage of dataset suggests an urgent need of comprehensive dataset construction. Moreover, there are a limited number (i.e., 89 apps per dataset on average) of apps in each dataset, making it unrealistic to cover all potential performance issues. In other words, even though some tools might effectively detect issues, they would still suffer from soundness issues due to the existence of false negatives. Overall, more effort should be dedicated to creating more representative and comprehensive datasets.
Answers to RQ1.3: Available datasets covers only 11 out of 21 performance factors, highlighting a significant gap in tool evaluation. Existing datasets lack the representation needed to cover all performance issues adequately.
3. A Large-scale Exploration of Real-world Apps
In RQ1 (Section 2), we systematically summarised the Android performance issues in academia. However, there may still be gaps between cutting-edge studies and real-world scenarios. Thus, we conducted an explanatory study by examining real-world performance issues discussions. Specifically, we collected reported performance issues from three popular platforms of Android ecosystem: Google Play, Stack Overflow, and GitHub. We 1) reviewed user feedback on Google Play to assess the impact on end-users; 2) explored the discussions on Stack Overflow to gain insights into common issues, discussions, and solutions shared within the developer community; and 3) analysed issues and commit histories on GitHub to understand the challenges developers are facing. While detailing research gaps, we also present code patterns that we believe could inspire future works, contributing to our broader community.
Figure 3. The process of real-world data collection.
3.1. Data Collection
Fig. 3 shows our three-step pipeline to obtain real-world performance issues, discussions, and solutions. Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub. Next, we filtered collected information by keywords obtained from RQ1. Lastly, we manually verified real-world performance issues, discussions, and their common patterns.
3.1.1. Data Crawling
To understand real-world performance issues, we collect data from three popular platforms in the Android development ecosystem as below.
[Google Play - User Reviews.] The Google Play Store is the biggest Android app store in the world. It constitutes an essential component of the Android ecosystem, facilitating the distribution and discovery of apps for millions of users worldwide. Users frequently complain about the performance issues of the Android apps by leaving reviews in the Google Play Store, providing useful information for researchers (Liu2014CharacterizingAD, ). We collected user reviews from the apps on Google Play based on the AppBrain popularity ranking [^2] AppBrain website https://www.appbrain.com/stats/google-play-rankings. To hold the representativeness, we collected the top 200 apps in five monetary categories (top free, top paid, top grossing, top new free, top new paid) across 30 countries, totaling 7,681 unique apps. We used the Google Play scraper library333google-play-scraper: https://pypi.org/project/google-play-scraper/ and successfully collected 909,430 user reviews from these apps. Please note that we excluded user reviews which only have a rating score (1-5) without comments. Intuitively, performance issues are embedded in the negative user reviews, so we used the sentiment analysis model SiEBERT (hartmann2023, ) to identify negative reviews. This model is a fine-tuned checkpoint of RoBERTa (Liu2019RoBERTaAR, ), with fine-tuning datasets from diverse text sources such as reviews and tweets. As a result, we roughly obtained 60,684 negative user reviews for further process.
[Stack Overflow - Question Threads.] Stack Overflow is a popular Q&A website for programmers and developers. It plays a crucial role in the Android ecosystem by providing a platform for knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ). Thus, we explored the performance issues developers consider and discuss in the development process. Each question threads on Stack Overflow normally reflects one real-world programming challenges, and the follow-up discussion in the thread provide additional thoughts, insights, and potential solutions. To understand developers’ opinions on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers, resulting in 448,977 questions. We further narrowed down to 360,372 threads containing the code tag for further refinement.
[Github - Issues & Commits.] GitHub is a critical platform in the Android ecosystem. Issues and commits are essential tools on GitHub that help manage the development process in Android development projects. Issues provide a structured way to handle reported performance issues, while commits ensure a clear and traceable history of code changes as solutions. Therefore, issues and commits can provide valuable insights about how developers actually solve those issues in the maintenance process. We used F-Droid (FDroid, ), an open-source Android app repository, to access each app’s corresponding GitHub repository. Here, we chose F-Droid because it is way more easier to inspect code patterns in open-source apps compared to closed-source ones. To gather representative Android apps, we crawled 17 categories of open-source Android apps and their GitHub links from F-Droid. Based on these links, we successfully identify 4,228 Android apps and their GitHub repositories, containing 35,278 issues and 550,973 commits.
3.1.2. Keyword Filtering
As we focus on Android performance issues, we adapted a keyword-based filtering to select relevant data only. Alongside of the systematical literature review of 48 academic research papers in Section 2.1.4, we also snowballed a list of keywords about Android performance issues. The keyword list includes terms like “energy waste”, “response time”, “responsive”, and so on. As a previously unknown performance issue might be caused by an existing factor and vice versa, keywords from existing literature do not restrict us from discovering new performance issues and new factors. Using these keywords, we filtered 60,684 user reviews to 132, 360,372 Stack Overflow discussions to 773, 35,278 GitHub issues down to 297, and 550,973 GitHub commits to 645.
3.1.3. Manual Checking.
After filtering data by performance-related keywords, we notice some data entries are false positive, meaning they contained keywords but were not actually related to performance issues or their root causes. Therefore, we further conducted a manual verification to ensure the data quality and relevance.
User Reviews. We removed those unrelated to performance issues, such as criticisms of developers. For example, “The app works well, but only supports SMBv1, which has been deprecated and known to be very insecure. There have been numerous complaints about this but the author has been stubbornly unresponsive about it…”
Question Threads. For Stack Overflow, we excluded question threads that are wrongly labelled on the code tag and those low quality threads that their negative votes outnumber positive votes.
GitHub - Issues. We first checked and removed issues from inactive (dead) projects, such as those where developers do not respond or have explicitly stated that the project is no longer maintained. Additionally, we excluded issues that do not involve performance issues, such as “Show me … ports mac address host name mDNS response time, perhaps option to ping again and option to traceroute.
GitHub - Commits. We manually checked all remaining commits, and removed those without clear commit messages. Additionally, we excluded commits with multiple changes, such as fixing various bugs or features, where performance was only a minor part, making it difficult to identify the root causes.
Above all, we obtained 104 user reviews, 332 stack overflow discussions, 78 GitHub issues, and 289 GitHub commits. Two of authors performed the aforementioned manual checking individually, for any disagreement, they discussed and reached the same answer. For persisting disagreement, a senior author will then involve and decide. We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement. The results indicate a Cohen’s kappa of 0.953 for user reviews, 0.897 for commits, 0.923 for GitHub issues, and 0.885 for Stack Overflow posts, all of which signify substantial agreement. We then manually identified the Android performance issues in these data and present our findings in the next section.
[…]
Figure 4. Distribution of performance issues and factors in user reviews, stack overflow, GitHub issues, and GitHub commits. Other Factors: Polling, Syntax, Hasmap, Cookies, Override, Thread Pool, Cache, Copy Text, Background Thread, HTTP Request.
3.2. Result Analysis
Fig. 4 illustrates the distribution of performance issues and factors across the three data sources. Overall, we identified six categories of performance issues (including one previously unseen) and 35 factors (including 19 previously unseen) from the three real-world data sources. Next, we analyzed our findings from the perspectives of User Reviews, Stack Overflow, and GitHub Issues & Commits.
3.2.1. Performance Issues and Factors in Real-world
We first identified performance issues in real-world apps and highlight the gaps between these issues and those discussed in research communities.
[Google Play - User Reviews.] User reviews typically focus on how performance issues negatively impact their experience without delving into the specific causes. Some user reviews mention more than one performance issue, so the percentages add up to more than 100%. We totally identified six performance issues, five of which were known from the literature review, while one performance issue was previously unseen. Responsiveness issues, such as slow responsiveness or UI freezes, have the most direct impact on user experience, appearing 79 times (75.96%). For example, one comment says: “… the software appears to be sluggish in map updating and does not properly locate my vehicle… I am hopeful these issues will be resolved in the future but if not then I will stop discontinue my premium subscription service and find another solution.” This indicates that unresolved performance issues can lead to users canceling subscriptions, causing companies to lose market share. Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries. These issues appeared 25 times (24.04%). E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes. drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app. Additionally, a performance issue not covered in existing literature is Data Consumption. E.g., one comment mentions: “.. Still cant even disable it, much less delete it…And worst of all, it still runs, using up data, memory and running down battery. Shame on You XXX…”
[Stack Overflow - Question Threads.] From the 332 question threads, we identified five categories of performance issues, all of which are already mentioned in the literature review (Section 2). Developers most frequently discussed Memory Consumption (80.72%), which is mainly caused by Image factor. However, the prevalence of Memory Consumption only ranks forth in existing research papers according to Fig. 2. Additionally, while Responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions. This discrepancy indicates that the performance issues developers focus on, those users are concerned about, and the ones researchers prioritize, do not align. Moreover, we identified a total of 30 factors contributing to these performance issues, and only 16 factors appeared in the literature review. Among the 14 newly discovered factors, the most frequent one is File Operation, which typically involves inefficiently downloading, uploading, or saving large files, leading to memory shortages. The next most frequent previously unseen factor is Data Type, which often lead to Memory Consumption and Responsiveness issues.
[Github - Issues & Commits.] Among the 289 commits, we identified four categories of performance issues, all of which are already mentioned in the literature review (Section 2). The most frequent performance issue is Memory Consumption (84.43%), followed by Responsiveness (11.07%). This distribution closely resembles that found in Stack Overflow discussions, indicating that developers are most concerned with memory consumption. In addition, we identified 22 factors, 11 of which are previously unseen. For example, the factor Parameter can lead to data loss or memory consumption. In Commit  [^4] https://github.com/ente-io/ente/commit/0165b7569b4acee1699de59c7c6ec1a8b91a7b4f, the change was to remove the fileUrl from console.log parameters to prevent a memory leak.
Regarding GitHub issues, we identified four performance issues and nine factors (including three unseen factors). The most frequently discussed issue is Data Loss (51.28%), followed by Memory Consumption (38.46%) and Responsiveness (8.97%). Interestingly, Data Loss is the most common issue in GitHub Issues, while Memory Consumption is the most frequently addressed in commits. This discrepancy indicates that the issues reported in GitHub issues are not always the ones developers prioritize for fixing.
Answers to RQ2.1 We identified an unseen performance issue related to Data Consumption and 19 previously unrecognized factors such as Data Type, File Operation, and Object Creation, highlighting gaps in research coverage. The primary concerns of researchers, developers, and users are not aligned: researchers mainly focus on energy consumption, developers prioritize memory consumption, and users are most concerned with responsiveness.
3.2.2. Common Code Patterns in Real-World
After discovering numerous new factors, we then wanted to delve deeper into the prevalent code patterns associated with these issues. Through analysis of Stack Overflow discussions, GitHub issues, and commits, we aim to identify the performance issues commonly faced by developers during development and examine their code patterns. In total, we summarized six common patterns as follows:
API Misuse. API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API order, and misuse of API parameters. These errors can result in severe performance consequences, such as ANR and delays. For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error. Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5). In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.
1const CopyTextOption = ({bottomSheetId, postMessage, sourceScreen}: Props) => {
2 const handleCopyText = useCallback(async () => {
3 ...
4 await dismissBottomSheet(bottomSheetId);
5 Clipboard.setString(postMessage);
6 };};
Listing 1: Code Example of API Order Misuse
Unreleased References Another significant pattern can be characterized by the failure to release references (or have them garbage collected). For example, if a Fragment is held as an implicit reference, it will not be garbage collected until the object holding the reference completes its execution. Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak. Listing 2 demonstrates a code snippet from Stack Overflow  [^5] https://stackoverflow.com/questions/56102382, where a singleton object holds an mContext reference (line 2). Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object. This prevents it from being garbage collected, leading to a memory leak.
1public class MySingleton() {
2 Private Context mContext;
3 Private static MySingleton mInstance;
4 ...
5 public void myMethod(Context context)
6 {this.mContext = context;...}}
Listing 2: Code Example of Unreleased Reference
Redundant Objects. The creation of redundant objects is a common pattern that causes performance issues. Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption. Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation. Moreover, creating new objects with the same functionality can also unnecessarily degrades performance. Listing 3 demonstrates a code snippet from Stack Overflow  [^6] https://stackoverflow.com/questions/20942671, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8). This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
1public class MyTextView extends TextView {
2 public void setTypeface(Typeface tf, int style) {
3 if (style == Typeface.BOLD) {
4 super.setTypeface(Typeface.createFromAsset(...));
5 } else if (style == Typeface.ITALIC) {
6 super.setTypeface(Typeface.createFromAsset(...));
7 } else {
8 super.setTypeface(Typeface.createFromAsset(...));
9 }}}
Listing 3: Code Example of Redundant Objects
Large-Scale Data. Large-scale data is also a culprit for performance issues. Loading large images without proper caching can cause excessive memory usage and slowdowns. Uploading large files without separating them into smaller chunks can lead to high memory consumption. As shown in a code snippet from Stack Overflow  [^7] https://stackoverflow.com/questions/8820837, where the buffer size is the same as the uploaded file. If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
1 ...
2 int length = 0;
3 newFile.createNewFile();
4 InputStream inputStream = ctx.getAssets().open("myBigFile.apk");
5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
6 byte[] buffer = new byte[inputStream.available()];
7 while ((length = inputStream.read(buffer)) > 0) {
8 fOutputStream.write(buffer, 0, length);
9 }
Listing 4: Code Example of Large-Scale Data
UI Operations. UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration. For example, in the code snippet 5 from Stack Overflow  [^8] https://stackoverflow.com/questions/42944376, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw in invoked, even though the UI content doesn’t change. This causes the application to slow down and increases the computational load.
1public class ConcentricCircularView extends View {...
2 protected void onDraw(Canvas canvas) {
3 super.onDraw(canvas);
4 ...
5 for (int i=0;i<10;i++){...}}}
Listing 5: Drawing the Same Content in the Loop
Other Patterns. Besides the previously mentioned patterns, there are several other types of patterns to be aware of, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager in the main thread, etc.
Answers to RQ2.2 We have meticulously identified six common code patterns that contribute to performance issues, which encompass API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
Figure 5. Performance issues and corresponding factors from literature review and real-world data. Colors represent the number of automated tools capable of analyzing each factor (0-5), with darker shades indicating more tools. White indicates no factor-to-performance relationship.
4. Future Trends
In this study, we summrized the relevant research on Android performance analysis and illuminated real-world challenges in Sections 2 and 3. Drawing upon these findings, we now present discussion points regarding the research and practice of Android performance analysis, aiming to guide future researchers in this field.
4.1. Targeted Objective Imbalance
Through our literature review (Section 2) and real-world exploration (Section 3), we identified a total of six types of performance issues and 40 factors, as shown in Fig. 5, where colors and white distinguish whether there is a relationship between factors and performance issues. We observed that, apart from the performance issue “Data Consumption”, all the other five performance issues involve corresponding factors, suggesting data consumption is an important but untouched topic.
In addition, of the 40 discovered factors, 21 have been analyzed in the literature, while the remaining 19 (almost half) have not been examined by our researchers, highlighting a significant research gap. Moreover, it’s intriguing to note that there are 5 factors (namely, Multi-Threads, Obsolete Task, Redundant Frames, SQL Statement, and Wakelock) that solely appear in the literature without a corresponding issue or complaint in real-world scenarios. This suggests that these factors may stem more from the assumptions of academic researchers rather than being genuine concerns of users or developers in practical settings.
The presence of this imbalance underscores the need for further research into issues genuinely concerning users or developers. Indeed, more effort should be put into addressing the untargeted real-world factors, aiming to bridge the existing gap between identified factors and practical concerns.
4.2. Automatic Performance Analysis Tools are Still in an Early Stage
Our review also revealed that only 7 out of 40 factors (17.50%) are covered by state-of-the-art (SOTA) tools. We utilize Fig. 5 to further highlight this significant research gap, where Activity is the most targeted factor, followed by Resource and AsyncTask, and finally Service, Database Operation, Image, and Loop. Also, the majority of these tools are focused on issue detection, while only 3 of them (30.00%) are dedicated to fixing such issues. Although static, dynamic, and DL techniques have been demonstrated to be effective at analyzing performance issues in a series of research experiments, applying these approaches to reveal performance issues holistically in practice remains unsolved. Future research should not only focus on proposing a comprehensive framework to tackle many realistic challenges, such as API misuse, redundant objects, etc, which may require knowledge of code patterns as presented in Section 3.2.2, but also on finding solutions for fixing problems.
In addition, the adopted techniques are also highly imbalanced (most existing tools use traditional static or dynamic analysis, with only one work (Jabbarvand2020AutomatedCO, ) leveraging DL methods). The lack of variety in technique selection hinders the soundness, as static approaches are not proficient in handling complicated code features (e.g., reflection (sun2021taming, ), obfuscation (samhi2022jucify, ), and hardening (Schloegel2022LokiHC, )), and it’s non-trivial for dynamic tools to achieve satisfactory results to trigger sufficient paths in practice. Thus, developing effective tools against performance issues will be a noteworthy research topic in the future. In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e. code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Lu2021CodeXGLUEAM, ; Wang2022TestDrivenML, )). Thus, it is promising to apply advanced DL techniques to assist in solving realistic and specific performance issues. For example, convert code exhibiting the Unreleased References pattern mentioned in Section 3.2.2 into program analysis graphs (e.g., CFG and DFG). Use these graphs as features to train or finetune a GNN-based model (Zhou2018GraphNN, ; Xu2018HowPA, ; Velickovic2017GraphAN, ) to identify this anti-pattern. Leverage the network’s interpretability (Ying2019GNNExplainerGE, ; Yuan2021OnEO, ) to pinpoint problematic code segments and suggest optimizations, such as adding code to release references.
4.3. Constructing Benchmark Dataset is an Urgent Need
Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 11 (27.50%) of the 40 factors are covered by existing datasets. Even among the 11 factors, the data distribution is highly imbalanced. The number of activity samples is much larger than that of the other factors, which are more severe from a real-world perspective. In addition, the number of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools. Moreover, another major issue is the limited reproducibility. Our review indicates that only a small number of primary studies share their datasets, making it difficult for future researchers to validate and evaluate tools. Furthermore, these datasets lack real-time updates for the latest app, thus containing outdated samples.
Thus, it is urgent to construct up-to-date, comprehensive benchmark datasets to support fair comparisons, covering as many real-world cases as possible. The code examples we summarized from the exploratory study can all be used to construct a comprehensive dataset, covering 35 factors. Another possible solution involves leveraging the findings of our exploratory study, wherein we identify six specific code patterns from real-world scenarios. These patterns can act as seeds for mutation tools, enabling the generation of additional data to construct datasets.
5. THREATS TO VALIDITY
Our study has several factors that could affect the validity of our findings in terms of literature review and large-scale exploration.
For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works. To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 2.1.2, and used five well-known electronic databases. To minimize the risk of missing relevant works, we conducted backward and forward snowballing and cross-checked our results.
For real-world data collection, the data filtering process may miss some relevant information. This is because the sentiment classification model (hartmann2023, ) used to filter user reviews may not guarantee 100% accuracy in identifying negative reviews. Additionally, the Google Play scraper might fail to capture all user reviews. Lastly, there could be biases in our manual checking process. To minimize subjective differences, two authors independently checked the filtered data. In cases of disagreement, they discussed and resolved whether to include the data.
6. Conclusion
In this paper, we conducted a comprehensive empirical study to reveal the current state of Android performance analysis and the research gap between real-world challenges and existing techniques. Based on our in-depth analysis of 48 relevant studies and real-world issues across three major platforms (Google Play, Stack Overflow, GitHub), we discovered that (1) There are 6 types of performance issues and corresponding 40 factors; (2) Performance issues are sophisticated, with a single issue potentially stemming from multiple factors and a single factor causing multiple issues; (3) The root causes of data consumption issue are not clear; (4) The primary concerns of researchers, developers, and users are not aligned; (5) The majority of existing approaches primarily utilize static analysis; (6) There is a significant gap in existing tools’ ability to address all factors causing performance issues; (7) There is an urgent need to construct benchmark dataset to evaluate automatic tools; (8) DL-based and LLM-based approaches have shown promising results in addressing software engineering problems, suggesting their potential for tackling Android performance issues as well.

### v2

@@
 Abstract.
-Performance plays a critical role in ensuring the smooth operation of any mobile application, directly influencing user engagement and retention.
-Android applications are no exception.
-However, unlike functionality issues, performance issues are more challenging to discover as their root causes are sophisticated and typically emerge under specific payloads.
-To tackle this problem, researchers have dedicated substantial efforts to proposing automatic approaches for understanding, detecting, and resolving performance issues.
-Despite these endeavors, it still remains unknown what the status quo of Android performance analysis is, and whether existing approaches can indeed accurately reflect real performance issues.
-To fill this research gap, we conducted a systematic literature review followed by an explanatory study to explore relevant studies and real-world challenges.
-Our findings reveal that current tools have limited capabilities, covering only 17.50% of the performance issues.
-Additionally, existing datasets encompass only 27.50% of the issues and are very limited in size.
-We also show real-world issue patterns, underscoring the huge gap between the identified techniques and practical concerns.
-Furthermore, possible solutions are provided to guide future research towards achieving effective performance issue detection and resolution.
+Performance issues in Android applications significantly undermine users’ experience, engagement, and retention, which is a long-lasting research topic in academia.
+Unlike functionality issues, performance issues are more difficult to diagnose and resolve due to their complex root causes, which often emerge only under specific conditions or payloads.
+Although many efforts haven attempt to mitigate the impact of performance issues by developing methods to automatically identify and resolve them, it remains unclear if this objective has been fulfilled, and the existing approaches indeed targeted on the most critical performance issues encountered in real-world settings.
+To this end, we conducted a large-scale comparative study of Android performance issues in real-world applications and literature.
+Specifically, we started by investigating real-world performance issues, their underlying root causes (i.e., contributing factors), and common code patterns.
+We then took an additional step to empirically summarize existing approaches and datasets through a literature review, assessing how well academic research reflects the real-world challenges faced by developers and users.
+Our comparison results show a substantial divergence exists in the primary performance concerns of researchers, developers, and users.
+Among all the identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets.
+This stark contrast underscores a substantial gap in our understanding and management of performance issues.
+Consequently, it is crucial for our community to intensify efforts to bridge these gaps and achieve comprehensive detection and resolution of performance issues.
+Android, Software Maintenance, Performance Issues
 conference: ; ; copyright: none
@@
 The exponential growth of Android usage has seen an unprecedented expansion within the mobile app industry, with millions competing for users’ attention.
-In such a fiercely competitive circumstance, developers are under pressure to create new, fancy apps in the quickest manner.
-However, this emphasis on rapid development can sometimes come at the expense of app quality, particularly in terms of performance.
-Unlike functionality issues that directly cause app malfunctions or failures, performance issues occur when the app functions correctly but its efficiency is low, resource usage is unreasonable, or user interaction is poor, significantly reducing user satisfaction and software reliability (Li2019DetectingAD, ; Liu2014CharacterizingAD, ; Gao2017EveryPC, ; Liu2018NavyDroidAE, ).
-These issues include, but are not limited to, excessive energy consumption (Banerjee2018EnergyPatchRR, ; Pathak2012WhatIK, ; Carette2017InvestigatingTE, ), resource leaks (Wu2016LightWeightIA, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Yan2013SystematicTF, ), and responsiveness delays (Farooq2018RuntimeDroidRR, ; Kang2016DiagDroidAP, ; Lin2024AgingOG, ).
-For example, Liu et al.
-(Liu2014CharacterizingAD, ) demonstrate that wasteful energy consumption can occur in background threads that continue to update invisible GUIs or perform complex network computations, rapidly depleting battery life and degrading app performance.
-Similarly, Yan et al.
-(Yan2013SystematicTF, ) reveal resource leak defects associated with events such as screen rotation, app switching, pressing the back button, and file or database access.
+In such a fiercely competitive circumstance, developers face pressure to create new, fancy apps in the quickest manner (Nikzad2014APEAA, ).
+However, this emphasis on speed often leads to a focus on functional requirements, while non-functional requirements, particularly performance, are overlooked—even though they are crucial to the overall quality of the apps (Hort2021ASO, ).
+According to the classic SE textbook from Ian Sommerville, performance issues arise when software meets functional requirements but operates inefficiently or wastes resources, which can negatively affect its reliability (sommerville2011software, ).
+These issues are particularly critical in Android mobile devices, where resources like battery life, CPU, memory, and storage are limited (Guo2013CharacterizingAD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Lee2019ImprovingEE, ).
+In Android apps, performance issues affecting efficiency, resource utilization, or app smoothness can drastically reduce user satisfaction and software reliability (Li2019DetectingAD, ; Liu2014CharacterizingAD, ; Gao2017EveryPC, ; Liu2018NavyDroidAE, ).
+These issues include, but are not limited to, excessive energy consumption (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Pathak2012WhatIK, ; Carette2017InvestigatingTE, ), redundant memory consumption (Guo2013CharacterizingAD, ; Song2021IMGDroidDI, ; Habchi2019TheRO, ), and responsiveness delays (Farooq2018RuntimeDroidRR, ; Kang2016DiagDroidAP, ; Lin2024AgingOG, ).
 However, due to the non-functional nature of performance issues, their problems do not immediately manifest.
@@
 Furthermore, some performance issues only appear under specific usage scenarios or loads (Behrouz2016EnergyawareTM, ; Jabbarvand2019SearchBasedET, ; Yan2013SystematicTF, ; Farooq2018RuntimeDroidRR, ).
-Consequently, manually identifying performance issues is time-consuming and labor-intensive, making it difficult for developers to address all problems during the testing phase.
-Therefore, it is crucial to use effective automated tools to analyze, detect, and fix performance issues (Habchi2018OnAL, ; Li2019DetectingAD, ; Li2022CombattingEI, ).
-Unfortunately, at the current stage, researchers face difficulties in designing automated tools to comprehensively address Android performance issues due to: 1) Lack of clear classification of performance issues and their root causes, making it hard to fully understand the performance issues; 2) Uncertainty about the state quo of existing approaches, making it unclear to what extent these tools are reliable for detecting or fixing performance issues; 3) The absence of benchmark datasets in evaluating state-of-the-art tools; 4) Uncertainty about whether state-of-the-art tools and datasets accurately reflect real performance issues.
-To guide future research, we first conducted a systematic literature review to provide a holistic taxonomy for existing performance issues and their underlying root causes.
-We then identified research gaps between state-of-the-art tools and datasets relative to the targeted research objectives (i.e., performance issues).
-Subsequently, we analyzed real-world performance issues to pinpoint differences between existing literature and practical challenges.
-Through these investigations, our goal is to offer a thorough understanding of Android performance issues and to guide the development of more effective automated tools and comprehensive datasets for evaluation.
-In this paper, we aim to explore the following research questions:
-RQ1: What is the status quo of Android performance issues analysis in research field?
-  • What performance issues are being targeted by researchers?
-The lack of a universally accepted taxonomy leads to many papers targeting the same performance problems using different terms.
-Additionally, since one factor can lead to multiple performance issues and one issue can be caused by multiple factors, creating a clear classification system is challenging.
-Our work filled this gap by distinguishing between the consequences and factors of performance issues.
-We identified five categories of Android performance issues and 21 distinct root causes.
-  • What approaches have been developed for performance issues?
-To solve the above performance issues, researchers have proposed various kinds of techniques.
-Thus, in the second research question, we aim to understand the status quo of current techniques, seeking to comprehensively grasp the research gap between performance issues and automatic tools.
-In total, we classified them into three primary techniques: static-based, dynamic-based, and deep learning-based approaches.
-Static analysis is the most frequently adopted technique, while deep learning-based methods are much less commonly utilized for performance issues.
-Moreover, it is surprising to find that only 7 out of 21 factors are targeted by these tools.
-This evidence confirms a significant gap between performance issues and automated resolutions.
-Additionally, the majority of these tools focus on issue detection rather than fixing, highlighting another research gap concerning the need for proactive problem-solving.
-  • What datasets have been used in Android performance analysis?
-High-quality datasets are essential for evaluating performance analysis tools.
-We found that current datasets cover only 11 out of 21 factors, addressing a limited range of issues.
-Furthermore, the number of existing datasets is very small.
-This highlights an urgent need for a comprehensive public dataset.
-RQ2: What are gaps between cutting-edge studies and real-world scenarios?
-  • What real-world performance issues in the wild?
-Given that performance is an inherently practical concern, we proceeded to conduct a deeper analysis of real-world issues to pinpoint differences between existing literature and practical challenges.
-Our examination revealed one previously unrecognized performance issue and 19 previously unacknowledged factors across real-world platforms.
-This suggests that numerous factors contributing to performance issues in practical settings have not received attention in research.
-Furthermore, we uncovered a misalignment between the primary concerns of researchers, developers, and users.
-These discoveries underscore a substantial gap in fully comprehending and mitigating performance issues in real-world scenarios.
-  • What are the common patterns exhibited in real-world performance issues?
-Based on our systematic analysis for real-world data, we identified six common code patterns: API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
-These newly discovered patterns provide insights for designing automated tools to address performance issues and for creating datasets to evaluate the effectiveness of these tools.
+Consequently, manually identifying performance issues is time-consuming and labor-intensive, making it difficult for developers to address all problems during the coding and testing phase.
+Therefore, it is essential for researchers to support developers in analyzing, detecting, and resolving Android performance issues (Habchi2018OnAL, ; Li2019DetectingAD, ; Li2022CombattingEI, ; Vsquez2015HowDD, ).
+Our research fellows have proposed several approaches to mitigate the impact of performance issues (Li2019CharacterizingAD, ; Banerjee2014DetectingEB, ; Song2021IMGDroidDI, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Vsquez2015HowDD, ; Hort2021ASO, ; Liu2019DroidLeaksAC, ; Habchi2019TheRO, ).
+For example, in the sense of static analysis, Li et al.
+(Li2019CharacterizingAD, ) developed TAPIR to analyze anti-patterns in app call graphs to detect inefficient image displays that can lead to memory consumption, energy consumption, and responsiveness issues.
+In addition, dynamic approaches, such as µDroid (Jabbarvand2017DroidAE, ), are proposed to detect energy waste due to resource under-utilization, wakelock misuse, and inefficient loops by monitoring abnormal energy consumption.
+However, given the scattered nature of current research in this field, it remains unclear whether performance issues have been comprehensively resolved.
+Specifically, it is unknown what the current state of Android app performance analysis is, whether current research is focused on key performance issues encountered in real-world settings, and to what extent existing methods effectively and comprehensively address these critical real-world challenges.
+Additionally, the capabilities of existing approaches are quite distinct, as confirmed by the fact of lacking a systematic taxonomy to map performance issues with their corresponding root causes (a.k.a, factors).
+Specifically, in this work, we formulate these concerns into two research questions (RQs) that we aim to answer through empirical evidence and explanatory results.
+The two RQs are summarized as follows.
+RQ1: What are the primary performance issues that developers and users are most concerned with in the real-world settings (RQ1.1)?
+And what are the root causes (RQ1.2) and common code patterns (RQ1.3) of these issues?
+RQ2: Do the performance issues investigated in academic research align with those that developers and users are most concerned with in real-world (RQ2.1)?
+To what extent do the tools developed in academic research address real-world performance issues (RQ2.2)?
+Are the datasets developed in academic research adequate for evaluating the effectiveness of these tools in tackling Android performance issues (RQ2.3)?
+To answer these RQs, we began by conducting a large-scale analysis of real-world discussions on Android performance issues.
+Specifically, we collected 60,684 negative reviews from Google Play, 360,372 question threads from Stack Overflow, and 35,278 issues with 550,873 commits from GitHub to identify performance issues experienced by users and developers, analyze their root causes (i.e., contributing factors), and examine common code patterns.
+Following this, we conducted a systematic literature review on Android performance analysis to investigate whether existing works indeed touch on the leading performance issues encountered in real-world settings.
+Through a comparison of academic findings with real-world data, we developed a comprehensive taxonomy covering seven categories of performance issues and 72 contributing factors (Table 1).
+This taxonomy allowed us to identify substantial gaps in current academic research on Android performance analysis.
+Specifically, it turns out there is a substantial divergence in primary concerns among researchers, developers, and users.
+Additionally, it is surprising to find that among all identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets.
+On top of these findings, we further propose future research directions to bridge these gaps.
+In this paper, we make the following major contributions:
+  • We conducted a large-scale exploratory study of real-world discussions on Android performance issues by analyzing crowd-sourced data from Google Play, Stack Overflow, and GitHub.
+This allowed us to identify the primary performance issues faced by users and developers, along with their root causes and common code patterns.
+  • We conducted a systematic literature review on Android performance analysis to examine the performance issues that researchers focus on, the tools developed to address these issues, and the datasets used to evaluate the effectiveness of these tools.
+  • By comparing literature with real-world data, we developed a taxonomy that categorizes seven types of performance issues and identifies 72 contributing factors, enabling us to assess how well current research objectives, tools, and datasets address real-world challenges.
+Based on these insights, we proposed future research directions to address the identified gaps.
+2. A Large-scale Exploration of Real-world Discussions on Android Performance Issues
+While performance issues frequently arise in real-world scenarios, it remains unclear which specific issues concern users and developers, along with their root causes and common code patterns.
+To address this, we conducted an exploratory study to collect, identify, and analyze real-world performance issues as discussed on popular online platforms.
+Figure 1. The process of real-world data collection.
+2.1. Data Collection
+Following common practices in analyzing large-scale online discussions (Guo2013CharacterizingAD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Banerjee2018EnergyPatchRR, ), we collected data from three popular crowd-sourced platforms: Google Play, Stack Overflow, and GitHub.
+Indeed, Google Play provides user feedback to assess the impact on end-users, Stack Overflow offers discussions that reveal common issues and solutions, and GitHub contains issues and commits histories that reflect the challenges developers encounter.
+Fig.
+1 shows our three-step pipeline to perform the real-world exploration.
+Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub.
+Then, we filtered this data using performance-related keywords.
+Lastly, we manually checked real-world performance issues, root causes, and their common code patterns.
+2.1.1. Data Crawling
+Here, we elaborate the details of data crawling as follows.
+[Google Play - User Reviews.] To collect user reviews, we focus on those from the Google Play Store, the largest Android app marketplace globally.
+This platform plays a crucial role in the Android ecosystem, enabling app distribution and discovery for millions of users worldwide.
+Users often report performance issues through their reviews on the Google Play Store, providing valuable insights for researchers (Liu2014CharacterizingAD, ).
+We collected user reviews from the apps on Google Play based on the AppBrain popularity ranking.
+To ensure representativeness, we collected the top 200 apps in five monetary categories (top free, top paid, top grossing, top new free, top new paid) across 30 countries, totaling 7,681 unique apps.
+We used the Google Play scraper library111google-play-scraper: https://pypi.org/project/google-play-scraper/ and successfully collected 909,430 user reviews from these apps.
+Please note that we excluded user reviews that only have a rating score (1-5) without comments.
+Intuitively, performance issues are embedded in the negative user reviews, so we used the sentiment analysis model SiEBERT (hartmann2023, ) to identify negative reviews.
+This model is a fine-tuned checkpoint of RoBERTa (Liu2019RoBERTaAR, ), with fine-tuning datasets from diverse text sources such as reviews and tweets.
+As a result, we obtained 60,684 negative user reviews for further processing.
+[Stack Overflow - Question Threads.] We also collected performance-related question threads from Stack Overflow, a widely recognized Q&A platform for programmers and developers, which plays a crucial role in the Android ecosystem by facilitating knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ).
+Specifically, we explored the performance issues developers posted and discussed in the development process.
+Each question thread on Stack Overflow typically reflects a real-world programming challenge, with the follow-up discussions offering additional insights, perspectives, and potential solutions.
+To understand developers’ perspectives on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers with code, resulting in 360,372 questions.
+[Github - Issues & Commits.] Additionally, we collected performance issues and commits for open-source apps from GitHub, the most widely used platform for open-source development.
+GitHub issues provide structured information of reported performance issues, while commits create a clear and traceable history of code fixes made as solutions.
+Consequently, both issues and commits offer valuable insights into how developers effectively resolve these issues during the maintenance process.
+To access each app’s corresponding GitHub repository, we utilized F-Droid (FDroid, ), an open-source Android app repository.
+We chose F-Droid because it facilitates easier inspection of code patterns in open-source applications compared to closed-source ones.
+To gather a representative sample of Android apps, we crawled 17 categories of open-source Android apps along with their GitHub links from F-Droid.
+As a result, we successfully identified 4,228 Android apps and their GitHub repositories, which contained 35,278 issues and 550,973 commits.
+2.1.2. Keyword Filtering
+As we focus on Android performance issues, we implemented a keyword-based filtering approach to select relevant data.
+To compile an accurate keyword list, we conducted a Google search for “Android performance”, reviewing the top 10 most relevant pages, which included 100 websites.
+Additionally, we examined several foundational papers on Android performance issues (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ).
+From these sources, we obtained a list of 87 keywords about Android performance issues.
+The keyword list includes terms like “UI freeze”, “battery drain”, “memory leak”, “cpu utilization”, “app size” and so on.
+As a previously unknown performance issue might be caused by an existing factor and vice versa, existing keywords do not restrict us from discovering new performance issues and new factors.
+Using these keywords, we filtered 60,684 user reviews to 165, 360,372 Stack Overflow discussions to 893, 35,278 GitHub issues down to 317, and 550,973 GitHub commits to 780.
+2.1.3. Manual Checking.
+After filtering the data using performance-related keywords, we noticed some data entries were false positive, meaning they contained keywords but were not actually related to performance issues.
+Therefore, we further conducted a manual verification to ensure the data quality and relevance.
+User Reviews.
+We removed user reviews that were unrelated to performance issues, such as criticisms of developers.
+For example, one review stated: “The app works well, but only supports SMBv1, which has been deprecated and known to be very insecure.
+There have been numerous complaints about this but the author has been stubbornly unresponsive about it…”
+Question Threads.
+For Stack Overflow, we excluded question threads that are wrongly labeled on the code tag as well as low-quality threads where the number of negative votes exceeded the number of positive votes.
+GitHub - Issues.
+We first checked for and removed issues from inactive (dead) projects, such as those where developers do not respond or have explicitly stated that the project is no longer maintained.
+Additionally, we excluded issues that do not involve performance issues, such as “Show me … ports mac address host name mDNS response time, perhaps option to ping again and option to traceroute.
+GitHub - Commits.
+We manually checked all remaining commits and removed those that lacked clear messages.
+Additionally, we excluded commits with multiple changes, such as fixes for various bugs or features where performance was a minor aspect, as this made identifying the root cause difficult.
+Ultimately, we obtained 114 user reviews, 336 stack overflow discussions, 47 GitHub issues, and 295 GitHub commits.
+Two of the authors performed the aforementioned manual checking individually, for any disagreement, they discussed their findings until reaching a consensus.
+For any persisting disagreement, a senior author made the final decision.
+We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement, yielding scores of 0.944 for user reviews, 0.885 for GitHub issues, 0.895 for GitHub commits, and 0.871 for Stack Overflow posts, all of which signify substantial agreement.
+We then manually identified the Android performance issues within this data and present our findings in the following section.
+[…]
+[…]
+[…]
+2.2. Result Analysis
+To gain a thorough understanding of prevalent performance issues in real-world apps, we begin by analyzing the distribution of performance issues encountered by users and developers (Section 2.2.1).
+Next, we categorize the underlying root causes contributing to these issues (Section 2.2.2).
+Finally, we identify code patterns commonly associated with these performance issues (Section 2.2.3).
+2.2.1. Performance Issues in Real-world
+We identified seven key performance issues that are of concern to users and frequently discussed by developers in real-world scenarios.
+Fig.
+2 shows the distribution of these issues across different platforms.
+Figure 2. Distribution of performance issues across various sources.
+[Google Play - User Reviews.] User reviews typically focus on how performance issues negatively impact their experience without delving into the specific causes.
+As shown in the User Review pie chart in Fig.
+2, users are mainly concerned about responsiveness, energy consumption, storage consumption, memory consumption, and internet data usage.
+Responsiveness issues, such as slow responsiveness or UI freezes, have the most direct impact on user experience, appearing in 62.3% of all reported reviews.
+This is not surprising as responsiveness is the most noticeable aspect of performance that users can directly perceive.
+For example, one comment says: “… the software appears to be sluggish in map updating and does not properly locate my vehicle… I am hopeful these issues will be resolved in the future but if not then I will stop discontinue my premium subscription service and find another solution.” This indicates that unresolved performance issues can lead to users canceling subscriptions, causing companies to lose market share.
+Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries.
+These issues appeared in 21.93% of all reported reviews.
+E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes.
+drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app.
+Storage Consumption is often complained by users that the size of the app is too large, or the cache space consumed is too large when using the app.
+These issues appeared in 9.65% of all reported reviews.
+As mobile devices have limited storage, excessive storage consumption can lead to problems like a lack of available space for other apps or files, and potential app crashes.
+E.g., one comment mentions: “..Every time a full size version of a Google app gets updated it eats up even more internal storage space… if there’s not enough storage space in your internal storage it won’t install certain updates or new apps from Google Play store, you have to uninstall updates or apps before reinstalling updates and apps…” Finally, Internet Data Usage issues are reported by users when apps consume excessive data, especially when they aren’t using Wi-Fi, making them more sensitive to data usage.
+For example, one user mentioned: “…Not sure why, but my data usage has skyrocketed.
+I’m on a limited plan, now I have to buy extra data every month…”
+[Stack Overflow - Question Threads.] From the 336 question threads, we identified six categories of performance issues, as illustrated in the Stack Overflow pie chart in Fig.
+2.
+Interestingly, two performance issues (i.e., CPU usage and GPU usage) are not mentioned in user reviews.
+Additionally, while responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions.
+This discrepancy indicates that the performance issues developers focus on and those users are concerned about do not align.
+Since user satisfaction directly impacts app market share, developers should prioritize addressing the issues most important to users.
+[Github - Issues & Commits.] The GitHub Issues and GitHub Commits pie charts in Fig.
+2 show a similar distribution of performance issues across both data sources, with memory consumption, responsiveness, and storage consumption being the most frequently mentioned.
+Notably, while internet data usage is a niche issue primarily highlighted in user reviews and absent from Stack Overflow and GitHub commits, it also appears in GitHub issues, providing insights into its underlying causes.
+In summary, the five performance issues that users mention in reviews are also discussed by developers on GitHub and Stack Overflow.
+Developers further address two additional issues, CPU usage and GPU usage, which are rarely mentioned by users, likely because they are technical concerns not directly perceived in user experience.
+Answers to RQ1.1 We identified seven categories of performance issues: responsiveness, energy consumption, storage consumption, memory consumption, internet data usage, CPU usage, and GPU usage.
+A substantial divergence exists between the focus areas of performance issues for users and developers: while users primarily emphasize concerns around responsiveness, developers’ discussions are predominantly oriented toward memory consumption.
+Table 1. Taxonomy of seven categories of performance issues and their contributing factors. ‘R-W.’ indicates real-world sources, while ‘Lit.’ refers to literature sources.
+Category Factor R-W. Lit. Problem Example
+Memory Consumption Loop Repeatedly reading the same value into the array in an infinite loop
+Object Creation Repeatedly creating new instances instead of reusing existing ones
+Massive Data Outputting thumbnail URLs containing a large amount of information to the log
+Database Operation Each database query loads new data into memory without clearing the previous data
+Activity An anonymous inner class strongly references the Activity, preventing it from being garbage collected
+File Handling Loading an entire large file into memory at once, rather than processing it in chunks
+Fragment The previous fragment state isn’t properly cleared when switching between fragments
+Dynamic UI Loading Recreating the view and bitmap for each list item during scrolling, without effectively reusing the views
+Background Thread Failing to stop the refresh when the app enters the background causes the UI component to remain active
+Singleton A singleton holds a reference to a UI component after its Activity is destroyed, preventing garbage collection
+Override Failing to call super.onDestroy() after overriding a method prevents proper release of system resources
+Configuration Gradle JVM heap size is insufficient during project builds, causing OOM errors
+Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
+Service Failing to destroy a dead service  (Song2019ServDroidDS, )
+Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Liu2019DroidLeaksAC, )
+AsyncTask Failing to destroy AsyncTask when it holds active GUI components (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
+View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, )
+Third-Party Library The API usage in the new version of the third-party library caused excessive memory usage (Rua2023ALE, )
+UI Rendering Reinflating views on each “getView” call without reusing them increases memory usage (Das2020CharacterizingTE, )
+Data Type Declaring a variable as static holds a strong reference to the previous task, preventing garbage collection (Das2020CharacterizingTE, )
+Obsolete Task Retaining obsolete parameters after updating view layouts  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
+Responsiveness Service Failing to move the notification to the foreground in time during service start, causing an ANR
+Inefficient Method Selecting a less efficient method introduces unnecessary delays in processing UI actions
+Background Thread Rapidly creating numerous threads in the background to perform I/O on the same data
+Cache Large cache causes frequent I/O operations, leading to lag
+Text Handling Repeatedly calling EditText.setText() to highlight all search results at once leads to unnecessary UI updates
+Resource Insufficient thread resource allocation causes slow response during simultaneous operations (Bhatt2020AutomatedRO, )
+Image 1) Decoding large images in the UI thread, 2) Loading large image, 3) dynamically adjusting content during activity-switching events (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
+Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
+UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Vsquez2015HowDD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Janssen2022OnTI, )
+Loop An infinite while loop in the button’s onClick() event makes the app unresponsive after clicking (Lyu2018RemoveRF, )
+Main Thread Performing heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ; Gmez2016MiningTR, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
+Activity Loading new resources during activity restarts may freeze the UI  (Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
+AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
+View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, ; Das2020CharacterizingTE, )
+Obsolete Task Downloading obsolete content occurs when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
+SQL Statement Retrieving more columns/rows than necessary in SQL queries  (Lyu2019QuantifyingTP, )
+Advertisements Loading ads while maintaining responsiveness to user actions (Petalotis2024AnES, )
+Message Backlogging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
+Storage Consumption File Handling Temp files from URIs aren’t cleared after upload
+Redundant Code Similar functional pages and activities that could be reused, or unnecessary code that only occupies storage
+Configuration Including unnecessary third-party library dependencies in the Gradle file
+Image PNG images compress less effectively than JPG, resulting in larger file sizes
+Energy Consumption Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Jabbarvand2017DroidAE, ; Nikzad2014APEAA, ; Wu2016Relda2AE, ; Vsquez2015HowDD, ; Bhatt2020AutomatedRO, ; Rua2023ALE, ; Liu2019DroidLeaksAC, )
+Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, )
+Database Operation Frequently triggering database queries for auto-complete (Li2019DetectingAD, ; Lyu2018RemoveRF, ; Li2022CombattingEI, ; Lyu2017AnES, )
+Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, ; Jabbarvand2017DroidAE, ; Li2014AnES, ; Palomba2019OnTI, )
+Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, )
+Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, )
+HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, ; Li2014AnES, )
+Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
+Image Decoding identical image multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
+Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, )
+Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, ; Jabbarvand2017DroidAE, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
+Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
+UI Rendering 1) Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Janssen2022OnTI, )
+AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
+Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, ; Palomba2019OnTI, )
+Display Screen The OLED display consumes more energy when displaying light colors than dark ones (Li2014MakingWA, ; LinaresVsquez2018MultiObjectiveOO, ; Vsquez2015OptimizingEC, ; Wan2015DetectingDE, )
+Program Language Developing web applications using JavaScript consumes more energy than WebAssembly (Hasselt2022ComparingTE, ; Oliveira2017ASO, )
+Advertisements Frequent advertisement and analytics (A&A) service requests lead to energy waste (Cito2016BatteryawareTI, ; Petalotis2024AnES, )
+Machine Learning The inefficient machine learning algorithm on mobile devices is overly complex (McIntosh2018WhatCA, )
+View Drawing all items in a ListView without reusing existing data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
+CPU Usage Multi-Threads The music player mishandles metadata extraction, spawning thousands of threads, overloading the CPU.
+Loop An infinite loop repeatedly calls methods to handle images
+Image Glide consumes significant CPU resources to sequentially display images when playing GIFs
+UI Rendering Unnecessarily using transparency to render UI components (Lee2019ImprovingEE, )
+GPU Usage Third-Party Library Third-party library’s bug causes continuous redrawing when using GPU to render UI components
+UI Rendering Spin animation has unnecessary frequent refreshes, forcing GPU rendering on every frame (Lee2019ImprovingEE, )
+Internet Data Usage Resource Downloading unwanted content without user permission, wasting internet resources
+Image Automatically loading animated GIFs without Wi-Fi enabled
+Advertisements Automatically downloading embedded advertising services in the background
+2.2.2. Root Causes of Performance Issues
+To address performance issues, it is crucial to investigate the underlying root causes that lead to these issues.
+Based on our real-world analysis, we manually summarized the root causes (i.e., contributing factors) for each performance category, as indicated by the checkmarks (✓) in the R-W.
+column in Table 1.
+Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic.
+The absence of a universally accepted taxonomy results in many discussions targeting the same performance problems but using different descriptive terms (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ).
+Additionally, since one factor can lead to multiple performance issues consequences, and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system.
+To address this gap, our work distinguishes between the consequences and factors of performance issues.
+Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors.
+To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner.
+We now elaborate on the seven major categories, each with the corresponding 49 factors along with an example presented in Table 1.
+Memory Consumption.
+As discussed in Section 2.2.1, memory consumption is the most frequently highlighted performance issue among developers.
+Our real-world analysis revealed 20 factors that contribute to these memory consumption issues, which are listed under the Memory Consumption category in Table 1.
+These issues arise from a variety of factors, often resulting from inefficient coding practices that lead to excessive or unnecessary memory use.
+For example, repeatedly reading the same value into an array within an infinite loop can cause memory to be continuously occupied without being released, ultimately leading to an out-of-memory (OOM) error.
+Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks.
+Responsiveness.
+Another important aspect of Android performance issues is responsiveness.
+Poor responsiveness leads to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc.
+In our analysis, we identified 14 key factors contributing to this issue.
+Many real-world scenarios show that heavy operations associated with UI threads can slow responsiveness, such as decoding large images in the UI threads.
+Besides, inefficient code logic can also slow down the response time.
+For example, improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
+Storage Consumption.
+As the third most concerning performance issue for developers, inefficient storage consumption often leads to apps occupying more storage space than expected, which can limit the device’s ability to install other apps and store data.
+Through our real-world exploration, we identified four key factors contributing to storage consumption issues.
+These factors typically arise from inefficient coding practices and poor data management, resulting in excessive or unnecessary storage usage.
+For instance, creating a large number of similar functional pages and activities that could have been reused, or including redundant code that never gets executed during runtime, wastes storage space.
+Moreover, adding unnecessary third-party library dependencies to the Gradle file further increases the app’s size after building, leading to larger APK files.
+Energy Consumption.
+As outlined in Section 2.2.1, energy consumption ranks as the second highest concern among users, as apps that drain battery faster than expected can frustrate users, especially given the limited battery resources on mobile devices.
+However, developers have not prioritized addressing this issue to the same extent.
+Consequently, in our real-world analysis, we identified only two key factors contributing to energy consumption problems.
+The root causes generally stem from inefficient resource usage and unnecessary background operations that provide no tangible benefits to users.
+For instance, updating unneeded sensor data in the background or running background services at unnecessarily high frequencies both result in excessive energy consumption.
+It is worth noting that, although developers do not prioritize energy efficiency, many research papers have extensively studied this issue, bridging the gap in developer attention.
+We will further discuss the performance issues targeted by researchers in Section 3.2.1.
+CPU Usage.
+Excessive CPU usage can cause the device to overheat, triggering the thermal management system to automatically lower the processor’s frequency to prevent overheating.
+This leads to a decline in overall system performance, affecting the normal operation of other apps and background services, and may even result in crashes, while also shortening the lifespan of the hardware and battery.
+Based on our analysis, we identified four distinct factors contributing to CPU usage issues.
+These issues typically stem from unnecessary or inefficient operations that waste CPU resources.
+For instance, using transparency in canvases where it’s not needed forces the browser to perform extra work, increasing CPU load.
+Similarly, auto-playing GIFs by sequentially displaying images instead of using a static preview image results in unnecessary CPU consumption.
+GPU Usage.
+The GPU is essential for handling graphics-intensive tasks like UI rendering, animations, 3D graphics, and video/image processing, reducing the CPU’s workload.
+Improper or excessive GPU use can lead to issues like UI freeze, frame drops, device overheating, rapid battery drain, and inefficient system resource allocation.
+Our real-world analysis identified two key factors contributing to GPU usage issues.
+These issues typically arise from overly frequent rendering requests or third-party libraries’ bugs.
+For example, a spin animation with excessively frequent refreshes forces the GPU to render every frame, unnecessarily consuming resources.
+Besides, a bug in a third-party library could cause an app’s misbehavior such as continuous redrawing of UI components, resulting in excessive GPU usage without providing any tangible benefits.
+Internet Data Usage.
+Network resources on mobile devices are limited, especially when Wi-Fi is unavailable, requiring users to purchase cellular data.
+Inefficient data usage can lead to unnecessary expenses and create a negative user experience.
+From our analysis, we identified three key factors contributing to this issue.
+The root causes often involve the unnecessary consumption of network resources.
+For example, downloading unwanted content without user consent or automatically downloading embedded advertisements in the background can lead to excessive data usage.
+Answers to RQ1.2: Performance issues are inherently complex, often arising from an interplay of multiple contributing factors, where a single issue may stem from various causes, and, conversely, a single factor may amplify multiple issues.
+We identified 49 distinct factors contributing to the seven categories of performance issues.
+2.2.3. Common Code Patterns in Real-World
+After identifying key factors of performance issues, we analyzed Stack Overflow discussions, GitHub issues, and commits to uncover common code patterns associated with these issues.
+We summarized six types of common patterns as follows:
+1const CopyTextOption = ({bottomSheetId, postMessage, sourceScreen}: Props) => {
+2 const handleCopyText = useCallback(async () => {
+3 ...
+4 await dismissBottomSheet(bottomSheetId);
+5 Clipboard.setString(postMessage);
+6 };};
+Listing 1: Code Example of API Order Misuse
+API Misuse.
+API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API orders, and misuse of API parameters.
+These errors can result in severe performance consequences, such as ANR and delays.
+For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error.
+Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5).
+In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.
+1public class MySingleton() {
+2 Private Context mContext;
+3 Private static MySingleton mInstance;
+4 ...
+5 public void myMethod(Context context)
+6 {this.mContext = context;...}}
+Listing 2: Code Example of Unreleased Reference
+Unreleased References.
+Another significant pattern can be characterized by the failure to release references (or have them garbage collected).
+For example, if a Fragment is held as an implicit reference, it won’t be garbage collected until the object holding the reference completes its execution.
+Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak.
+Listing 2 demonstrates a code snippet from Stack Overflow, where a singleton object holds an mContext reference (line 2).
+Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object.
+This prevents it from being garbage collected, leading to a memory leak.
+1public class MyTextView extends TextView {
+2 public void setTypeface(Typeface tf, int style) {
+3 if (style == Typeface.BOLD) {
+4 super.setTypeface(Typeface.createFromAsset(...));
+5 } else if (style == Typeface.ITALIC) {
+6 super.setTypeface(Typeface.createFromAsset(...));
+7 } else {
+8 super.setTypeface(Typeface.createFromAsset(...));}}}
+Listing 3: Code Example of Redundant Objects
+Redundant Objects.
+The creation of redundant objects is a common pattern that causes performance issues.
+Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption.
+Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation.
+Moreover, creating new objects with the same functionality can also unnecessarily degrade performance.
+Listing 3 demonstrates a code snippet from Stack Overflow, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8).
+This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
+1 ...
+2 int length = 0;
+3 newFile.createNewFile();
+4 InputStream inputStream = ctx.getAssets().open("myBigFile.apk");
+5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
+6 byte[] buffer = new byte[inputStream.available()];
+7 while ((length = inputStream.read(buffer)) > 0) {
+8 fOutputStream.write(buffer, 0, length);}
+Listing 4: Code Example of Large-Scale Data
+Large-Scale Data.
+Large-scale data is also a culprit for performance issues.
+Loading large images without proper caching can cause excessive memory usage and slowdowns.
+Uploading large files without separating them into smaller chunks can lead to high memory consumption.
+As shown in a code snippet from Stack Overflow, where the buffer size is the same as the uploaded file.
+If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
+1public class ConcentricCircularView extends View {...
+2 protected void onDraw(Canvas canvas) {
+3 super.onDraw(canvas);
+4 ...
+5 for (int i=0;i<10;i++){...}}}
+Listing 5: Drawing the Same Content in the Loop
+UI Operations.
+UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration.
+For example, in the code snippet 5 from Stack Overflow, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw is invoked, even though the UI content doesn’t change.
+This causes the application to slow down and increases the computational load.
+Other Patterns.
+Besides the previously mentioned patterns, other types of patterns also warrant attention, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager on the main thread, etc.
+Answers to RQ1.3 We identified six prevalent code patterns that commonly contribute to performance issues: API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
 3. Studies on Performance Issues
-In this work, we are interested in how far we are towards automating performance issues analysis in Android apps.
-Thus, in this section, we first resort to a systematic literature review (SLR) to understand the status quo of Android app performance analysis.
+In the above Section 2, we systematically examined real-world Android performance issues, their root causes, and common code patterns.
+However, it remains unclear whether cutting-edge literature targets the same practical issues, whether existing techniques can effectively resolve them, and whether current datasets can comprehensively evaluate the techniques’ effectiveness.
+To answer these questions, we conducted a systematic literature review (SLR) to assess the current state of Android performance analysis.
+By comparing the performance issues, techniques, and datasets discussed in research papers with those observed in real-world discussions, we aim to uncover gaps between academic research and real-world challenges.
 3.1. Literature Review
-To conduct systematic literature review, we followed the methodologies outlined by Brereton et al.
+To conduct a systematic literature review, we followed the methodologies outlined by Brereton et al.
 (brereton2007lessons, ) and Kitchenham et al.
@@
 3 illustrates the working process which comprises 5 steps.
+Figure 3. The working process of our systematic literature review.
 3.1.1. Keywords Identification
@@
 Each group is made up of several keywords.
+G1 includes terms related to Android mobile phone-related keywords.
+G2 includes terms related to performance issues identified in Section 2.
+Please note that to ensure comprehensive coverage and avoid missing relevant studies, we also included broader terms like “performance” and “issue”.
 The search string is then constructed by combining them, i.e., g1 AND g2, where g1 and g2 represent the disjunction of keywords from groups G1 and G2, respectively.
@@
 G1: android, mobile, *phone*
-G2: performance, resource, energy, responsiveness, issue*
-Figure 1. The working process of our systematic literature review.
+G2: resource, response, energy, memory, launch, freeze, lag, battery, responsiveness, cpu, gpu, storage, latency, performance, issue*
 Table 2. CORE A/A* ranked software engineering and mobile venues.
 Type Source Research Codes Venues
-Journals CORE2020 0803, 0805 TSE, TOSEM, JSS, IST, TMC
-Conferences CORE2023 4606, 4612 MOBICOM, Mobisys, EuroSys, MOBIHOC, MSWIM, ASE, ESEC/FSE, ICSE, ISCA, EASE, ECSA, ESEM, FOSSACS, ICSA, ICSME, ICST, ISSRE, ISSTA, MSR, SANER, SEAMS
+Journals CORE2020 0803, 0805 TSE, TOSEM, EMSE, JSS, IST, TMC
+Conferences CORE2023 4606, 4612 MOBICOM, Mobisys, EuroSys, MOBIHOC, MSWIM, ASE, ESEC/FSE, ICSE, EASE, ECSA, ESEM, FOSSACS, ICSA, ICSME, ICST, ISSRE, ISSTA, MSR, SANER, SEAMS
 3.1.2. Repository Search
 To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library.
-From this initial search, we retrieved a total of 10,766 papers.
-However, not all of them were of high quality, and many were outside our area of focus.
+From this initial search, we retrieved a total of 11,442 papers.
+However, many of these studies were outside the scope of our focus.
 To narrow down to the most relevant papers, we concentrated on the fields of software engineering and mobile, selecting research codes 4606 and 4612 for conferences, and 0805 and 0803 for journals, specifically targeting publications containing the keywords “software” and “mobile”.
-In addition, to further guarantee the quality of papers, we exclusively chose those published from conferences rated CORE A*/A  [^1] CORE provides conference and journal ratings.
-https://www.core.edu.au/home.
-We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the title of relevant publications.
-As a result, we successfully identified 274 publications distributed across 22 venues as listed in Table 1 (i.e., there is no relevant paper identified in 4 of the venues).
+In addition, we exclusively chose those published from conferences and journals rated CORE A*/A [^2] CORE provides conference and journal ratings.
+For conference: https://portal.core.edu.au/conf-ranks/.
+For journal: https://portal.core.edu.au/jnl-ranks/ from 2012 to 2024.
+We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the titles of relevant publications.
+As a result, we successfully identified 441 publications distributed across 23 venues as listed in Table 2.
+Notably, there is no relevant paper identified in ECSA, FOSSACS, and SEAMS.
 3.1.3. Paper Exclusion
 To gather relevant papers as possible, we initially considered all returned results from our search.
-However, not every paper pertains to Android performance issues.
+However, not all papers aligned with our focus on performance issues in mobile apps caused by poor coding practices, as discussed in Section 1.
 To refine our collection, we conducted a thorough review of the abstracts (and full content when necessary) of these papers, aiming to retain only those closely related to our research focus.
@@
   2. (2)
-    Papers targeting on non-Android mobile devices (e.g., Windows Phone, IOS devices) were excluded.
+    Papers targeting on non-Android devices (e.g., Windows Phone, IOS devices) were excluded.
   3. (3)
-    Papers using Android performance issues as metrics/indicators for comparison were excluded. For example, Mantis (Kwon2013MantisAP, ) predicts app execution time based on app implementations.
+    Papers using Android performance issues as metrics/indicators for comparison were excluded.
+For instance, Mantis (Kwon2013MantisAP, ) predicts app execution time based on app implementations, and Eprof (Pathak2012WhereIT, ) is a profiler used to monitor energy consumption in apps.
   4. (4)
-    Papers targeting on Android performance issues but not focused on the mobile app level were excluded.
-For example, Hypnos (Jindal2013HypnosUA, ) is used for detecting sleep conflicts in smartphone device drivers, which is outside our scope focused on mobile app issues stemming from poor coding practices by developers.
-After applying these exclusion criteria, we retained 27 papers that are closely related to the topic of performance issues analysis at Android mobile level.
+    Papers targeting Android performance issues but not focused on the mobile app level were excluded.
+For example, Hypnos (Jindal2013HypnosUA, ) is used for detecting sleep conflicts in smartphone device drivers, which is outside our scope and focused on mobile app issues stemming from poor coding practices by developers.
+After applying these exclusion criteria, we retained 49 papers that are closely related to the topic of performance issues analysis at the Android mobile level.
 3.1.4. Back/Forward Snowballing
-After initially identifying relevant papers (i.e., 27 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature.
+After initially identifying relevant papers (i.e., 49 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature.
 The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources.
@@
 We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified.
-Eventually, we were able to locate a total of 48 papers.
-Please note that during snowballing, we also collected 82 keywords related to performance issues for real-world data collections in Section 3.1.2.
+Eventually, we were able to locate a total of 66 papers.
 3.1.5. Cross Check
-Even with the relevant selection criteria, it is not guaranteed that the quality of all identified papers is desirable.
-Therefore, we evaluated the quality of each paper by reading its full text.
 Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies.
-Ultimately, we collected 48 papers to answer our research questions defined in Section 1.
-2.2. Result
-In the following paragraphs of Section 2, we summarize the state quo from thorough reading of these papers and attempt to understand: 1) Targeted performance issues, which involve key factors that lead to problems needing resolution, 2) Automatic techniques designed to address mobile performance issues, and 3) Datasets designed to evaluate the effectiveness of these automatic techniques.
-Table 2. The top problems targeted by the examined papers.
-Category Factor Problem Example
-Energy Consumption Database Operation Frequently triggering database queries for auto-complete (Li2019DetectingAD, ; Lyu2018RemoveRF, ; Li2022CombattingEI, ; Lyu2017AnES, )
-Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, )
-Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, )
-Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, )
-HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, )
-Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, )
-Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
-Image Decoding identical image in multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
-Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, )
-Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, )
-Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
-Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
-UI Rendering Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
-AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
-Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, )
-View Drawing all items in a ListView without reusing exising data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
-HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
-Responsiveness Obsolete Task Downloading obsolete content occurs when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
-Image 1) Decoding large images in the UI thread, 2) Loading large image, 3) dynamically adjusting content during activity-switching events (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
-Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
-UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
-SQL Statement Retrieving more columns/rows than necessary in SQL queries  (Lyu2019QuantifyingTP, )
-Message Backloging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
-Main Thread Performing Heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, )
-Activity Loading new resources during activity restarts may freeze the UI  (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
-AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
-View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, )
-Memory Consumption Obsolete Task Retaining obsolete parameters after updating view layouts  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
-Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
-Service Failing to destroy a dead service  (Song2019ServDroidDS, )
-Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
-AsyncTask Failing to destroy AsyncTask when it holds active GUI components (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
-View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, )
-HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
-Data Loss Activity Failing to save and restore GUI data during activity restarts from runtime changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
-AsyncTask Updating destroyed GUI components when it has been recreated using AsyncTask (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
-Distorted UI Display Activity Overlapping fragments occur when a new activity is restarted with a new fragment attached while the old fragment is still retained by the system (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
-2.2.1. Result - Targeted Problem
-We now highlight the representative problems targeted by our fellow researchers (cf.
-Table 2).
-Based on the literature review results, we manually summarized and identified five major categories of Android mobile performance issues: Energy Consumption, Memory Consumption, Responsiveness, Data Loss, and Distorted UI Display.
-Then we reviewed the root causes (a.k.a factors) for each category thoroughly and demonstrate them in Table 2 along with the specific performance problems our researcher fellows are working on.
-Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic.
-The lack of a universally accepted taxonomy results in many papers targeting the same performance problems but using different descriptive terms.
-Additionally, since one factor can lead to multiple performance issues (consequences), and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system.
-Therefore, our work addresses this gap by distinguishing between the consequences and factors of performance issues.
-Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors.
-To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner.
-We now elaborate on the five major categories, each with the corresponding factors along with an example presented in Table 2.
-Energy Consumption.
+Ultimately, we collected 66 papers to assess the current state of research and identify gaps in addressing real-world performance issues.
+3.2. Results
+In this Section, we summarize the current state quo of Android performance analysis by thoroughly reviewing these selected papers.
+Our aim is to assess whether academic research objectives, tools, and datasets align with the real-world challenges faced by developers and users.
+Specifically, we evaluate: 1) Targeted performance issues and key factors that researchers focus on (Section 3.2.1), 2) Automatic techniques designed to address Android performance issues (Section 3.2.2), and 3) Datasets designed to evaluate the effectiveness of these automatic techniques (Section 3.2.3).
+3.2.1. Result - Targeted Performance Issues
+In our literature review, we categorized the performance issues discussed in the collected papers and identified the contributing factors to each issue.
+This approach allowed us to supplement the performance issues and factors not uncovered in our real-world exploration to obtain a holistic taxonomy.
+Besides, by comparing these targeted issues and factors in academic research with those observed in real-world discussions, we evaluated how well researchers’ focus aligns with the practical challenges faced by developers and users.
+Holistic Taxonomy.
+Through reviewing 66 research papers, we identified five types of performance issues and 44 contributing factors, marking each in the Lit.
+column of Table 1 with a checkmark (✓).
+By incorporating insights from both real-world discussions and research papers, we developed a final taxonomy encompassing seven types of performance issues and 72 factors.
+By examining these issues and their contributing factors from both real-world data and academic findings, we identified significant gaps between research focus and real-world challenges as follows.
+Figure 4. Annual publications on each performance issue.
+Gaps between research focus and real-world performance issues.
 As shown in Fig.
-2, energy consumption receives the most research attention, with 34 out of the examined papers (70.83%) and its increasing trend is expected to continue.
-This result is not surprising given that the most urgent task at the moment is to ensure that the app remains user-friendly and sustainable in the rapidly evolving app market, which is why many previous studies focused on this research topic.
-Specifically, we found that most of the energy consumption issues stem from a common pattern: frequent energy-consuming operations occurring together with loops/lifecycle functions.
-For example, Lyu et al.
-(Lyu2018RemoveRF, ) found that database writes within loops trigger inefficient autocommit behaviors, leading to significant energy waste.
-Liu et al.
-(Liu2014CharacterizingAD, ) discovered that even when the GUI is in the background and not visible, threads continue updating the UI, causing unnecessary energy waste.
-In addition, invoking APIs for requiring resources (e.g., wakelock, sensor) when they are not needed can also results in a substantial increase in energy consumption (Liu2014GreenDroidAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Banerjee2018EnergyPatchRR, ).
-Responsiveness.
-Another important aspect of Android performance issues is responsiveness.
-Fig.
-2 shows that 18 reviewed articles (37.50%) are crafted for poor responsiveness, leading to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc.
-Specifically, we identified 10 distinct factors that contribute to this category.
-Most primary studies (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ) discovered that heavy operations associated with UI threads can slow responsiveness.
-Additionally, Gao et al.
-(Gao2017EveryPC, ) and Sarra et al.
-(Habchi2019TheRO, ) show that improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
-Memory Consumption.
-There are 14 papers (29.17%) related to memory consumption issues in the collected studies.
-These memory consumption issues have diverse factors, often stemming from inefficient coding that leads to excessive memory use.
-For example, decoding images without resizing them results in extensive memory consumption (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ).
-Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks (Lin2015StudyAR, ; Pan2020StaticAC, ).
-Data Loss With regard to the data loss problem, Fig.
-2 indicates that 7 papers (14.58%) attempt to analyze data loss issues, which significantly degrade user experience when data fails to be saved or restored in UI threads.
-For instance, studies have revealed that failing to save and restore GUI data during activity restarts can cause existing data to vanish instantly when the screen orientation changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ).
-Distorted UI Display.
-Another aspect of performance issues relates to distorted UI display, which consists of only 2.08% (i.e., 1 study) of the collected studies.
-Specifically, Umar et al.
-(Farooq2018RuntimeDroidRR, ) shows that overlapping fragments occur when initializing a new fragment while retaining the old one, causing a poor user experience.
-Figure 2. Annual publications on each performance issue.
-Discussion.
-Among all the collected studies on performance issues for Android, it appears that energy consumption is the most frequently studied topic, followed by responsiveness analysis.
-Additionally, it is noteworthy that an increasing number of recent studies have started to address specific issues related to memory consumption and data loss (e.g., runtime change handling (Farooq2018RuntimeDroidRR, ; Chen2023TransparentRC, )), although the total number of relevant studies remains small.
-Moreover, the amount of literature focusing on distorted UI display is relatively small, indicating a need for further in-depth research.
-While this review categorizes the existing literature’s research objectives into five categories, the scope of Android performance issues can be much broader.
-Therefore, future research should go beyond these categories and propose new approaches tailored to emerging app scenarios.
-Answers to RQ1.1: Current research studies focus on five categories of performance issues, prioritized by their prevalence: Energy Consumption (70.83%), Responsiveness (37.50%), Memory Consumption (29.17%), Data Loss (14.58%), and Distorted UI Display (2.08%).
-We provide a holistic taxonomy of Android performance issues, encompassing 5 major categories and 21 distinct root causes.
-Performance issues are sophisiticated, as a single issue may stem from multiple factors and a single factor can cause multiple issues.
+4, we identified five key performance issues from literature: energy consumption, responsiveness, memory consumption, CPU usage, and GPU usage.
+These performance issues are also prevalent in real-world scenarios, indicating that researchers have made significant efforts to address concerns that are relevant to both developers and users.
+Despite these efforts, there remains a noticeable imbalance in focus.
+Among the five issues, energy consumption received the most attention, with 53 of the examined papers (80.30%) focusing on this issue, and this trend is likely to continue.
+However, despite this focus on energy consumption, Section 2.2 revealed that users are more concerned with responsiveness, while developers prioritize memory consumption.
+This highlights a discrepancy between the performance issues prioritized by researchers, developers, and users.
+Furthermore, two performance issues prevalent in real-world scenarios have not been explored in literature: storage consumption and internet data usage.
+These gaps suggest that researchers should place greater emphasis on the performance issues that are of primary concern to both users and developers in real-world apps.
+Gaps in academic research on real-world performance issue factors.
+By analyzing the cases where both the R-W.
+and Lit.
+columns are marked with a checkmark (✓), we found that only 21 out of 49 factors discussed in real-world settings (42.86%) have been studied in literature.
+It indicates that the majority of performance issues encountered in real-world scenarios remain underexplored in academic research.
+For instance, memory consumption, which is the top concern for developers, real-world data shows 20 contributing factors, but only 8 (40.00%) have been explored in literature.
+Similarly, for responsiveness, the top concern for users, 14 real-world factors were identified, but only nine (64.29%) have been studied in literature.
+Less common issues like CPU usage and GPU usage have seen even less attention, with only 25.00% and 50.00% of their respective real-world root causes being studied in literature.
+These gaps suggest that many root causes of real-world performance issues still remain unexplored by researchers.
+Notably, all identified root causes of energy consumption have been addressed in literature, indicating that this area has been extensively studied.
+This observation aligns with our finding that the majority of research efforts are concentrated on energy consumption.
+The presence of research-only factors not seen in real-world discussions.
+By examining rows where the R-W.
+column is unmarked (✗) but the Lit.
+column is marked (✓), we identified 23 factors that have been explored in literature but not observed in the real world.
+Specifically, we found one unseen factor in memory consumption, four in responsiveness, and 18 in energy consumption.
+This suggests that literature may uncover factors that did not appear in our real-world exploration, thereby further enriching the taxonomy.
+Notably, most of these unseen factors pertain to energy consumption, accounting for 90.00% of the factors explored in literature but not observed in the real-world.
+For example, Li et al.
+(Li2014MakingWA, ) found that OLED displays consume more energy when displaying light colors than dark colors.
+Similarly, Cito et al.
+(Cito2016BatteryawareTI, ) observed that frequent advertisement and analytics (A&A) service requests lead to significant energy waste.
+On the other hand, there is a substantial number of energy consumption factors that appear only in the literature, yet rarely manifest in real-world applications.
+This indicates a pronounced disconnection between academic research and the practical challenges faced by practitioners.
+Answers to RQ2.1: Through an extensive integration of real-world explorations and thorough literature review, we present a comprehensive taxonomy encapsulating seven primary types of performance issues and 72 contributing factors.
+A noteworthy finding shows that only 42.86% factors contributing to performance issues have been examined in academic research.
+A significant misalignment exists in the primary concerns of researchers, developers, and users: researchers concentrate on energy consumption, developers prioritize memory usage, and users are chiefly concerned with responsiveness.
 3.2.2. Result - Technique
-To address the above performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches.
-Table 3 presents the the list of available automatic approaches collected from our examined papers.
-Despite numerous claims of proposed automated tools, only a small fraction (i.e., 10 out of the 48 examined papers (20.83%)) works have made their tool publicly available.
+To address performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches.
+Table 3 presents the list of available automatic approaches collected from our examined papers.
+Among the 66 papers examined, eight (12.12%) provided publicly accessible tools designed to detect specific factors contributing to particular performance issues.
+This small fraction can be attributed to several reasons: some papers lack accessible artifacts, others provide only datasets without the corresponding tool code, and in some cases, the links to the artifacts are invalid.
 This indicates a significant gap between automatic tools with their practical usage to within our community.
-In addition, among the identified factors leading to performance issues, only 7 out of 21 are targeted by existing tools.
-The fact that the majority of the factors are overlooked by SOTA tools indicates a huge gap in addressing performance issues.
-Moreover, most automatic tools are designed with a primary focus on simply detecting performance issues rather than actively fixing them.
-This emphasizes a need for a shift towards developing tools that not only identify but also offer solutions to enhance performance.
-To better understand how these available tools are implemented, we now summarize the techniques adopted in these approaches.
+In addition, we found that existing tools target only 17 out of the 72 factors in our taxonomy (Table 1), covering just 23.61%.
+This indicates that the majority of factors contributing to the seven primary types of performance issues are overlooked by SOTA tools, highlighting a significant gap in addressing these issues comprehensively.
+Moreover, most automated tools are primarily designed to detect performance issues rather than actively resolve them.
+Given that these detection tools rely on specific code patterns, we can anticipate that the identified patterns can also be applied to remediate performance issues.
+We then summarize the approaches as follows:
 Table 3. A summary of automatic approaches used for Android mobile performance analysis. Static denotes Static Analysis, Dynamic denotes Dynamic Analysis, and DL denotes Deep Learning methods.
-Tool/Reference Approach Technique Stage Target Factor Link
-DDLDroid (Zhou2023DDLDroidED, ) Static Reverse Engineering Taint analysis Detect Activity  (DDLDroidTool, )
-DLD (Riganelli2020DataLD, ) Dynamic Runtime Monitoring Model-based Testing Detect Activity  (Riganelli2020DataLDTool, )
-iFixDataloss (Guo2022DetectingAF, ) Static Dynamic Fuzzing Search-based Testing Static Program Repair Detect Fix Activity  (Guo2022DetectingAFTool, )
-RCHDroid (Chen2023TransparentRC, ) Dynamic Runtime Management Garbage Collection Fix Activity  (Chen2023TransparentRCTool, )
-VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Activity Resource  (Lu2022DetectingRUTool, )
-ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Resource  (Jabbarvand2020AutomatedCOTool, )
-AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect AsyncTask  (Pan2020StaticACTool, )
-TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Image  (Li2019CharacterizingADTool, )
-Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Database Loop  (Lyu2018RemoveRFTool, )
-ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural- Code Analysis Detect Service  (Song2019ServDroidDSTool, )
+Tool Approach Technique Stage Performance Issue Target Factor Link
+VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Memory Consumption Energy Consumption Activity Resource  (Lu2022DetectingRUTool, )
+AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
+TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
+ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural Code Analysis Detect Memory Consumption Energy Consumption Service  (Song2019ServDroidDSTool, )
+Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Energy Consumption Responsiveness Database Operation Loop  (Lyu2018RemoveRFTool, )
+PlumbDroid (Bhatt2020AutomatedRO, ) Static Control-Flow Analysis Patch Generation Detect Fix Responsiveness Resource  (Bhatt2020AutomatedROTool, )
+µDroid (Jabbarvand2017DroidAE, ) Dynamic Mutation Testing Power Trace Analysis Detect Energy Consumption Resource Wakelock Loop  (Jabbarvand2017DroidAETool, )
+ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Energy Consumption Resource  (Jabbarvand2020AutomatedCOTool, )
 Static Analysis.
-Static approaches are the art of reasoning about the behavior of programs without actually running them.
-It has been widely used (70.00%) to either detect or fix performance issues, as shown in Table 3.
-Specifically, taint analysis, symbolic execution, code pattern analysis and data flow analysis are the most commonly adopted techniques for rule-based static analysis as many performance issues arise from common code patterns.
+Most existing approaches are based on static analysis to detect or fix performance issues, as shown in Table 3.
+This is intuitive since static analysis involves reasoning about program behavior without execution, making it lightweight to deploy Specifically, taint analysis, symbolic execution (Lu2022DetectingRU, ), code pattern analysis (Pan2020StaticAC, ; Li2019CharacterizingAD, ; Song2019ServDroidDS, ), and data flow analysis (Lyu2018RemoveRF, ) are commonly employed in rule-based static analysis as many performance issues arise from common code patterns.
 Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ).
 Moreover, they can also provide fundamental information for dynamic fuzzing tools to trigger performance bugs on the fly (Guo2022DetectingAF, ).
-Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Guo2022DetectingAF, ).
+Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Lyu2018RemoveRF, ; Bhatt2020AutomatedRO, ).
 Dynamic Analysis.
-There are only four (i.e.
-40.00%) primary studies that use dynamic analysis approaches to detect or fix performance bugs (especially for data loss issues).
-This result is align with expectation because dynamic analysis involves executing apps in a controlled environment and unable to cover all possible paths (dong2020timemachine, ).
+There are only two studies that primarily employ dynamic analysis approaches to detect performance bugs.
+This result aligns with expectations because dynamic analysis involves executing apps in a controlled environment and is unable to cover all possible paths (dong2020timemachine, ).
 However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ).
-Table 3 shows that three dynamic tools are designed to address data loss issues in the context of activity restarts associated with sophisticated lifecycle methods, which are challenging to be detected statically.
+As shown in Table 3, both dynamic analysis tools target energy consumption issues, particularly unnecessary energy waste such as resource under-utilization, wakelock misuse, and inefficient loops, which are detectable through abnormal energy consumption monitoring.
 We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
 Deep Learning.
-With the increasing popularity of deep learning (DL) techniques, we are surprised to find that only one paper adopted DL techniques for performance analysis.
+Despite the growing popularity of deep learning (DL) in software engineering tasks like code analysis (Allamanis2017LearningTR, ), bug detection (Pradel2018DeepBugsAL, ), and program repair (zhao4254659autopatch, ), surprisingly, only one study applied DL to Android performance analysis.
 Specifically, Jabbarvand et al.
 (Jabbarvand2020AutomatedCO, ) observed runtime features dynamically and leveraged DL models to locate energy anomalies.
-Although the number of DL-related studies is relatively small, DL techniques have been widely employed in various software engineering tasks, such as code analysis (Allamanis2017LearningTR, ), bug detection (Pradel2018DeepBugsAL, ), and program repair (zhao4254659autopatch, ).
-It is promising to see a garnered considerable research attention to employ ML/DL techniques to thwart Android performance issues.
-Answers to RQ1.2: The majority of existing approaches primarily utilize static analysis (70.00%), with dynamic analysis (40.00%) coming next, and deep learning (10.00%).
-Only 10 out of the examined papers (20.83%) offer publicly accessible automated tools, highlighting a significant accessibility gap.
-Only 7 factors (33.33%) are addressed by existing automated tools, indicating incomplete coverage and potential oversight.
-Most tools prioritize issue detection over resolution, indicating a need for proactive problem-solving.
-Table 4. A list of datasets used for Android mobile performance analysis.
-Reference Year # APP Type Target Factor Link
-Guo et al. (Guo2022DetectingAF, ) 2022 66 Real-World Activity  (Guo2022DetectingAFTool, )
-Chen et al. (Chen2023TransparentRC, ) 2023 27 Real-World Activity  (Chen2023TransparentRCTool, )
-Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Activity Resource  (Lu2022DetectingRUTool, )
-Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) 2020 9 Micro Resource  (Jabbarvand2020AutomatedCOTool, )
-AsyncBench (Pan2020StaticAC, ) 2020 69 Micro AsyncTask  (Pan2020StaticACTool, )
-Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Image  (Li2019CharacterizingADTool, )
-Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Database Operations Loop  (Lyu2018RemoveRFTool, )
-Song et al. (Song2019ServDroidDS, ) 2019 45 Real-World Service  (Song2019ServDroidDSTool, )
-Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World UI Rendering View Inefficient Method Hashmap  (Habchi2019TheROData, )
+The limited adoption of applying DL to Android performance analysis may due to challenges in collecting large datasets that are sufficient to train DL models on performance issues and their corresponding fixes.
+Answers to RQ2.2: We identified eight publicly available tools, primarily using static analysis, dynamic analysis, and deep learning approaches.
+Existing tools target only 17 of the 72 identified factors to performance issues (23.61%), leaving a substantial number of factors unexamined.
+Most tools focus on detecting performance issues while neglecting their resolution.
+Table 4. A list of datasets used for Android mobile performance analysis, based on our literature review.
+Dataset Year # APP Type Performance Issue Target Factor Link
+Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Energy Consumption Responsiveness Database Operations Loop  (Lyu2018RemoveRFTool, )
+Vásquez et al. (LinaresVsquez2018MultiObjectiveOO, ) 2018 5 Real-World Energy Consumption Display Screen  (LinaresVsquez2018MultiObjectiveOOTool, )
+Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World Memoy Consumption Energy Consumption Responsiveness UI Rendering View Inefficient Method  (Habchi2019TheROData, )
+Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
+AsyncBench (Pan2020StaticAC, ) 2020 69 Micro Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
+Liu et al. (Liu2019DroidLeaksAC, ) 2020 34 Real-World Memory Consumption Energy Consumption Resource  (Liu2019DroidLeaksACTool, )
+Das et al. (Das2020CharacterizingTE, ) 2020 316 Real-World Memory Consumption Energy Consumption UI Rendering Data Type Inefficient Method View Wakelock  (Das2020CharacterizingTETool, )
+Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Memory Consumption Energy Consumption Resource  (Lu2022DetectingRUTool, )
+Janssen et al. (Janssen2022OnTI, ) 2022 40 Real-World Energy Consumption Responsiveness UI Rendering  (Janssen2022OnTITool, )
+Petalotis et al. (Petalotis2024AnES, ) 2024 9 Real-World Energy Consumption Responsiveness Advertisements  (Petalotis2024AnESTool, )
 3.2.3. Result - Dataset
-To evaluate the aforementioned automatic tools, researchers have proposed several datasets targeting unique performance factors.
-After examining all the public links or websites within the 48 identified studies, we were surprised to find only 9 (18.75%) unique open-sourced datasets, as shown in Table 4.
+To collect datasets, several types of papers need to be included: 1) Tool papers that construct datasets to evaluate their designed tools; 2) Empirical studies that collect specific datasets to analyze performance issues, where labeled data can be used to support evaluation.
+After examining all the public links or websites within the 66 identified studies, we identified 10 (15.15%) unique, available, and open-sourced datasets, as shown in Table 4.
+Notably, we conducted the examination in October 2024.
 Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison.
-Additionally, the number of factors covered by existing datasets is quite limited, w.r.t 11 factors in total.
-The shortage of dataset suggests an urgent need of comprehensive dataset construction.
-Moreover, there are a limited number (i.e., 89 apps per dataset on average) of apps in each dataset, making it unrealistic to cover all potential performance issues.
-In other words, even though some tools might effectively detect issues, they would still suffer from soundness issues due to the existence of false negatives.
-Overall, more effort should be dedicated to creating more representative and comprehensive datasets.
-Answers to RQ1.3: Available datasets covers only 11 out of 21 performance factors, highlighting a significant gap in tool evaluation.
-Existing datasets lack the representation needed to cover all performance issues adequately.
-3. A Large-scale Exploration of Real-world Apps
-In RQ1 (Section 2), we systematically summarised the Android performance issues in academia.
-However, there may still be gaps between cutting-edge studies and real-world scenarios.
-Thus, we conducted an explanatory study by examining real-world performance issues discussions.
-Specifically, we collected reported performance issues from three popular platforms of Android ecosystem: Google Play, Stack Overflow, and GitHub.
-We 1) reviewed user feedback on Google Play to assess the impact on end-users; 2) explored the discussions on Stack Overflow to gain insights into common issues, discussions, and solutions shared within the developer community; and 3) analysed issues and commit histories on GitHub to understand the challenges developers are facing.
-While detailing research gaps, we also present code patterns that we believe could inspire future works, contributing to our broader community.
-Figure 3. The process of real-world data collection.
-3.1. Data Collection
-Fig.
-3 shows our three-step pipeline to obtain real-world performance issues, discussions, and solutions.
-Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub.
-Next, we filtered collected information by keywords obtained from RQ1.
-Lastly, we manually verified real-world performance issues, discussions, and their common patterns.
-3.1.1. Data Crawling
-To understand real-world performance issues, we collect data from three popular platforms in the Android development ecosystem as below.
-[Google Play - User Reviews.] The Google Play Store is the biggest Android app store in the world.
-It constitutes an essential component of the Android ecosystem, facilitating the distribution and discovery of apps for millions of users worldwide.
-Users frequently complain about the performance issues of the Android apps by leaving reviews in the Google Play Store, providing useful information for researchers (Liu2014CharacterizingAD, ).
-We collected user reviews from the apps on Google Play based on the AppBrain popularity ranking [^2] AppBrain website https://www.appbrain.com/stats/google-play-rankings.
-To hold the representativeness, we collected the top 200 apps in five monetary categories (top free, top paid, top grossing, top new free, top new paid) across 30 countries, totaling 7,681 unique apps.
-We used the Google Play scraper library333google-play-scraper: https://pypi.org/project/google-play-scraper/ and successfully collected 909,430 user reviews from these apps.
-Please note that we excluded user reviews which only have a rating score (1-5) without comments.
-Intuitively, performance issues are embedded in the negative user reviews, so we used the sentiment analysis model SiEBERT (hartmann2023, ) to identify negative reviews.
-This model is a fine-tuned checkpoint of RoBERTa (Liu2019RoBERTaAR, ), with fine-tuning datasets from diverse text sources such as reviews and tweets.
-As a result, we roughly obtained 60,684 negative user reviews for further process.
-[Stack Overflow - Question Threads.] Stack Overflow is a popular Q&A website for programmers and developers.
-It plays a crucial role in the Android ecosystem by providing a platform for knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ).
-Thus, we explored the performance issues developers consider and discuss in the development process.
-Each question threads on Stack Overflow normally reflects one real-world programming challenges, and the follow-up discussion in the thread provide additional thoughts, insights, and potential solutions.
-To understand developers’ opinions on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers, resulting in 448,977 questions.
-We further narrowed down to 360,372 threads containing the code tag for further refinement.
-[Github - Issues & Commits.] GitHub is a critical platform in the Android ecosystem.
-Issues and commits are essential tools on GitHub that help manage the development process in Android development projects.
-Issues provide a structured way to handle reported performance issues, while commits ensure a clear and traceable history of code changes as solutions.
-Therefore, issues and commits can provide valuable insights about how developers actually solve those issues in the maintenance process.
-We used F-Droid (FDroid, ), an open-source Android app repository, to access each app’s corresponding GitHub repository.
-Here, we chose F-Droid because it is way more easier to inspect code patterns in open-source apps compared to closed-source ones.
-To gather representative Android apps, we crawled 17 categories of open-source Android apps and their GitHub links from F-Droid.
-Based on these links, we successfully identify 4,228 Android apps and their GitHub repositories, containing 35,278 issues and 550,973 commits.
-3.1.2. Keyword Filtering
-As we focus on Android performance issues, we adapted a keyword-based filtering to select relevant data only.
-Alongside of the systematical literature review of 48 academic research papers in Section 2.1.4, we also snowballed a list of keywords about Android performance issues.
-The keyword list includes terms like “energy waste”, “response time”, “responsive”, and so on.
-As a previously unknown performance issue might be caused by an existing factor and vice versa, keywords from existing literature do not restrict us from discovering new performance issues and new factors.
-Using these keywords, we filtered 60,684 user reviews to 132, 360,372 Stack Overflow discussions to 773, 35,278 GitHub issues down to 297, and 550,973 GitHub commits to 645.
-3.1.3. Manual Checking.
-After filtering data by performance-related keywords, we notice some data entries are false positive, meaning they contained keywords but were not actually related to performance issues or their root causes.
-Therefore, we further conducted a manual verification to ensure the data quality and relevance.
-User Reviews.
-We removed those unrelated to performance issues, such as criticisms of developers.
-For example, “The app works well, but only supports SMBv1, which has been deprecated and known to be very insecure.
-There have been numerous complaints about this but the author has been stubbornly unresponsive about it…”
-Question Threads. For Stack Overflow, we excluded question threads that are wrongly labelled on the code tag and those low quality threads that their negative votes outnumber positive votes.
-GitHub - Issues.
-We first checked and removed issues from inactive (dead) projects, such as those where developers do not respond or have explicitly stated that the project is no longer maintained.
-Additionally, we excluded issues that do not involve performance issues, such as “Show me … ports mac address host name mDNS response time, perhaps option to ping again and option to traceroute.
-GitHub - Commits.
-We manually checked all remaining commits, and removed those without clear commit messages.
-Additionally, we excluded commits with multiple changes, such as fixing various bugs or features, where performance was only a minor part, making it difficult to identify the root causes.
-Above all, we obtained 104 user reviews, 332 stack overflow discussions, 78 GitHub issues, and 289 GitHub commits.
-Two of authors performed the aforementioned manual checking individually, for any disagreement, they discussed and reached the same answer.
-For persisting disagreement, a senior author will then involve and decide.
-We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement.
-The results indicate a Cohen’s kappa of 0.953 for user reviews, 0.897 for commits, 0.923 for GitHub issues, and 0.885 for Stack Overflow posts, all of which signify substantial agreement.
-We then manually identified the Android performance issues in these data and present our findings in the next section.
-[…]
-[…]
-[…]
-Figure 4.
-Distribution of performance issues and factors in user reviews, stack overflow, GitHub issues, and GitHub commits.
-Other Factors: Polling, Syntax, Hasmap, Cookies, Override, Thread Pool, Cache, Copy Text, Background Thread, HTTP Request.
-3.2. Result Analysis
-Fig.
-4 illustrates the distribution of performance issues and factors across the three data sources.
-Overall, we identified six categories of performance issues (including one previously unseen) and 35 factors (including 19 previously unseen) from the three real-world data sources.
-Next, we analyzed our findings from the perspectives of User Reviews, Stack Overflow, and GitHub Issues & Commits.
-3.2.1. Performance Issues and Factors in Real-world
-We first identified performance issues in real-world apps and highlight the gaps between these issues and those discussed in research communities.
-[Google Play - User Reviews.] User reviews typically focus on how performance issues negatively impact their experience without delving into the specific causes.
-Some user reviews mention more than one performance issue, so the percentages add up to more than 100%.
-We totally identified six performance issues, five of which were known from the literature review, while one performance issue was previously unseen.
-Responsiveness issues, such as slow responsiveness or UI freezes, have the most direct impact on user experience, appearing 79 times (75.96%).
-For example, one comment says: “… the software appears to be sluggish in map updating and does not properly locate my vehicle… I am hopeful these issues will be resolved in the future but if not then I will stop discontinue my premium subscription service and find another solution.” This indicates that unresolved performance issues can lead to users canceling subscriptions, causing companies to lose market share.
-Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries.
-These issues appeared 25 times (24.04%).
-E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes.
-drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app.
-Additionally, a performance issue not covered in existing literature is Data Consumption.
-E.g., one comment mentions: “..
-Still cant even disable it, much less delete it…And worst of all, it still runs, using up data, memory and running down battery.
-Shame on You XXX…”
-[Stack Overflow - Question Threads.] From the 332 question threads, we identified five categories of performance issues, all of which are already mentioned in the literature review (Section 2).
-Developers most frequently discussed Memory Consumption (80.72%), which is mainly caused by Image factor.
-However, the prevalence of Memory Consumption only ranks forth in existing research papers according to Fig.
-2.
-Additionally, while Responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions.
-This discrepancy indicates that the performance issues developers focus on, those users are concerned about, and the ones researchers prioritize, do not align.
-Moreover, we identified a total of 30 factors contributing to these performance issues, and only 16 factors appeared in the literature review.
-Among the 14 newly discovered factors, the most frequent one is File Operation, which typically involves inefficiently downloading, uploading, or saving large files, leading to memory shortages.
-The next most frequent previously unseen factor is Data Type, which often lead to Memory Consumption and Responsiveness issues.
-[Github - Issues & Commits.] Among the 289 commits, we identified four categories of performance issues, all of which are already mentioned in the literature review (Section 2).
-The most frequent performance issue is Memory Consumption (84.43%), followed by Responsiveness (11.07%).
-This distribution closely resembles that found in Stack Overflow discussions, indicating that developers are most concerned with memory consumption.
-In addition, we identified 22 factors, 11 of which are previously unseen.
-For example, the factor Parameter can lead to data loss or memory consumption.
-In Commit  [^4] https://github.com/ente-io/ente/commit/0165b7569b4acee1699de59c7c6ec1a8b91a7b4f, the change was to remove the fileUrl from console.log parameters to prevent a memory leak.
-Regarding GitHub issues, we identified four performance issues and nine factors (including three unseen factors).
-The most frequently discussed issue is Data Loss (51.28%), followed by Memory Consumption (38.46%) and Responsiveness (8.97%).
-Interestingly, Data Loss is the most common issue in GitHub Issues, while Memory Consumption is the most frequently addressed in commits.
-This discrepancy indicates that the issues reported in GitHub issues are not always the ones developers prioritize for fixing.
-Answers to RQ2.1 We identified an unseen performance issue related to Data Consumption and 19 previously unrecognized factors such as Data Type, File Operation, and Object Creation, highlighting gaps in research coverage.
-The primary concerns of researchers, developers, and users are not aligned: researchers mainly focus on energy consumption, developers prioritize memory consumption, and users are most concerned with responsiveness.
-3.2.2. Common Code Patterns in Real-World
-After discovering numerous new factors, we then wanted to delve deeper into the prevalent code patterns associated with these issues.
-Through analysis of Stack Overflow discussions, GitHub issues, and commits, we aim to identify the performance issues commonly faced by developers during development and examine their code patterns.
-In total, we summarized six common patterns as follows:
-API Misuse.
-API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API order, and misuse of API parameters.
-These errors can result in severe performance consequences, such as ANR and delays.
-For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error.
-Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5).
-In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.
-1const CopyTextOption = ({bottomSheetId, postMessage, sourceScreen}: Props) => {
-2 const handleCopyText = useCallback(async () => {
-3 ...
-4 await dismissBottomSheet(bottomSheetId);
-5 Clipboard.setString(postMessage);
-6 };};
-Listing 1: Code Example of API Order Misuse
-Unreleased References Another significant pattern can be characterized by the failure to release references (or have them garbage collected).
-For example, if a Fragment is held as an implicit reference, it will not be garbage collected until the object holding the reference completes its execution.
-Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak.
-Listing 2 demonstrates a code snippet from Stack Overflow  [^5] https://stackoverflow.com/questions/56102382, where a singleton object holds an mContext reference (line 2).
-Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object.
-This prevents it from being garbage collected, leading to a memory leak.
-1public class MySingleton() {
-2 Private Context mContext;
-3 Private static MySingleton mInstance;
-4 ...
-5 public void myMethod(Context context)
-6 {this.mContext = context;...}}
-Listing 2: Code Example of Unreleased Reference
-Redundant Objects.
-The creation of redundant objects is a common pattern that causes performance issues.
-Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption.
-Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation.
-Moreover, creating new objects with the same functionality can also unnecessarily degrades performance.
-Listing 3 demonstrates a code snippet from Stack Overflow  [^6] https://stackoverflow.com/questions/20942671, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8).
-This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
-1public class MyTextView extends TextView {
-2 public void setTypeface(Typeface tf, int style) {
-3 if (style == Typeface.BOLD) {
-4 super.setTypeface(Typeface.createFromAsset(...));
-5 } else if (style == Typeface.ITALIC) {
-6 super.setTypeface(Typeface.createFromAsset(...));
-7 } else {
-8 super.setTypeface(Typeface.createFromAsset(...));
-9 }}}
-Listing 3: Code Example of Redundant Objects
-Large-Scale Data.
-Large-scale data is also a culprit for performance issues.
-Loading large images without proper caching can cause excessive memory usage and slowdowns.
-Uploading large files without separating them into smaller chunks can lead to high memory consumption.
-As shown in a code snippet from Stack Overflow  [^7] https://stackoverflow.com/questions/8820837, where the buffer size is the same as the uploaded file.
-If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
-1 ...
-2 int length = 0;
-3 newFile.createNewFile();
-4 InputStream inputStream = ctx.getAssets().open("myBigFile.apk");
-5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
-6 byte[] buffer = new byte[inputStream.available()];
-7 while ((length = inputStream.read(buffer)) > 0) {
-8 fOutputStream.write(buffer, 0, length);
-9 }
-Listing 4: Code Example of Large-Scale Data
-UI Operations.
-UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration.
-For example, in the code snippet 5 from Stack Overflow  [^8] https://stackoverflow.com/questions/42944376, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw in invoked, even though the UI content doesn’t change.
-This causes the application to slow down and increases the computational load.
-1public class ConcentricCircularView extends View {...
-2 protected void onDraw(Canvas canvas) {
-3 super.onDraw(canvas);
-4 ...
-5 for (int i=0;i<10;i++){...}}}
-Listing 5: Drawing the Same Content in the Loop
-Other Patterns.
-Besides the previously mentioned patterns, there are several other types of patterns to be aware of, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager in the main thread, etc.
-Answers to RQ2.2 We have meticulously identified six common code patterns that contribute to performance issues, which encompass API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
-Figure 5.
-Performance issues and corresponding factors from literature review and real-world data.
-Colors represent the number of automated tools capable of analyzing each factor (0-5), with darker shades indicating more tools.
-White indicates no factor-to-performance relationship.
+In addition, the number of factors covered by these datasets is quite limited.
+Out of the 72 factors identified in our taxonomy, only 24 (33.33%) are addressed by existing datasets.
+This indicates a significant need for constructing more comprehensive datasets.
+Moreover, there is a considerable imbalance in dataset size.
+The smallest dataset contains only five apps, while the largest has 324, with an average of 105.7 apps per dataset.
+This imbalance makes it unrealistic for any single dataset to cover all potential issues comprehensively.
+Even if some tools are effective at detecting performance issues, they may still face soundness challenges due to false negatives.
+Therefore, more effort is needed to create datasets that are not only comprehensive but also more representative of the diverse range of performance issues.
+Answers to RQ2.3: We identified 10 open-source datasets for evaluating performance issues; however, they encompass only 24 out of 72 factors (33.33%), highlighting a significant gap in tool evaluation.
+These datasets suffer from inadequate sample sizes, with the smallest containing only five apps and the largest consisting of 324.
 4. Future Trends
-In this study, we summrized the relevant research on Android performance analysis and illuminated real-world challenges in Sections 2 and 3.
-Drawing upon these findings, we now present discussion points regarding the research and practice of Android performance analysis, aiming to guide future researchers in this field.
+Through our real-world exploration (Section 2) and literature review (Section 3), we have revealed significant gaps in current academic research: it addresses only a fraction of the performance issues that developers and users prioritize, existing tools lack the capacity to comprehensively tackle these issues, and current datasets are insufficient for effectively evaluating tool performance.
+Drawing upon these findings, we present discussion points regarding the research and practice of Android performance analysis, aiming to guide future researchers in this field.
 4.1. Targeted Objective Imbalance
-Through our literature review (Section 2) and real-world exploration (Section 3), we identified a total of six types of performance issues and 40 factors, as shown in Fig.
-5, where colors and white distinguish whether there is a relationship between factors and performance issues.
-We observed that, apart from the performance issue “Data Consumption”, all the other five performance issues involve corresponding factors, suggesting data consumption is an important but untouched topic.
-In addition, of the 40 discovered factors, 21 have been analyzed in the literature, while the remaining 19 (almost half) have not been examined by our researchers, highlighting a significant research gap.
-Moreover, it’s intriguing to note that there are 5 factors (namely, Multi-Threads, Obsolete Task, Redundant Frames, SQL Statement, and Wakelock) that solely appear in the literature without a corresponding issue or complaint in real-world scenarios.
-This suggests that these factors may stem more from the assumptions of academic researchers rather than being genuine concerns of users or developers in practical settings.
-The presence of this imbalance underscores the need for further research into issues genuinely concerning users or developers.
-Indeed, more effort should be put into addressing the untargeted real-world factors, aiming to bridge the existing gap between identified factors and practical concerns.
+In Section 3.2.1, we found that there is a misalignment in the primary concerns of researchers, developers, and users: researchers primarily focus on energy consumption, developers prioritize memory consumption, while users are most concerned with responsiveness.
+This suggests that future research should increase focus on responsiveness and memory issues, as these have a more immediate impact on both end-users and developers.
+We also found two categories of performance issues—storage consumption and internet data usage—that remain unexplored in academic research, indicating potential areas for further investigation.
+Additionally, only 42.86% of the factors contributing to performance issues observed in real-world settings have been explored in academic research, remaining nearly 60% still awaiting for researchers to investigate.
+Interestingly, there are 23 factors among the 72 factors lead to performance issues are only been explored in literature but not observed in the real-world scenarios.
+This suggests that the issues currently discussed in academic research may not be of real concern to users or developers in real-world settings.
+Overall, this imbalance between academic research objectives and real-world challenges underscores the need for greater focus on the real-world performance issues and contributing factors that are most relevant to end-users and developers.
 4.2. Automatic Performance Analysis Tools are Still in an Early Stage
-Our review also revealed that only 7 out of 40 factors (17.50%) are covered by state-of-the-art (SOTA) tools.
-We utilize Fig.
-5 to further highlight this significant research gap, where Activity is the most targeted factor, followed by Resource and AsyncTask, and finally Service, Database Operation, Image, and Loop.
-Also, the majority of these tools are focused on issue detection, while only 3 of them (30.00%) are dedicated to fixing such issues.
-Although static, dynamic, and DL techniques have been demonstrated to be effective at analyzing performance issues in a series of research experiments, applying these approaches to reveal performance issues holistically in practice remains unsolved.
-Future research should not only focus on proposing a comprehensive framework to tackle many realistic challenges, such as API misuse, redundant objects, etc, which may require knowledge of code patterns as presented in Section 3.2.2, but also on finding solutions for fixing problems.
-In addition, the adopted techniques are also highly imbalanced (most existing tools use traditional static or dynamic analysis, with only one work (Jabbarvand2020AutomatedCO, ) leveraging DL methods).
+Our study revealed that only 17 out of 72 factors (23.61%) are addressed by state-of-the-art (SOTA) tools, with most tools focusing on detection rather than resolution.
+In fact, only two out of the existing tools are dedicated to fixing performance issues.
+Although static, dynamic, and deep learning (DL) techniques have shown effectiveness in research experiments, applying these methods to holistically detect and fix performance issues in practice remains an unsolved challenge.
+Future research should focus not only on building comprehensive frameworks to detect real-world performance issues, but also on providing solutions for fixing these issues.
 The lack of variety in technique selection hinders the soundness, as static approaches are not proficient in handling complicated code features (e.g., reflection (sun2021taming, ), obfuscation (samhi2022jucify, ), and hardening (Schloegel2022LokiHC, )), and it’s non-trivial for dynamic tools to achieve satisfactory results to trigger sufficient paths in practice.
@@
 In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e.
-code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Lu2021CodeXGLUEAM, ; Wang2022TestDrivenML, )).
-Thus, it is promising to apply advanced DL techniques to assist in solving realistic and specific performance issues.
-For example, convert code exhibiting the Unreleased References pattern mentioned in Section 3.2.2 into program analysis graphs (e.g., CFG and DFG).
-Use these graphs as features to train or finetune a GNN-based model (Zhou2018GraphNN, ; Xu2018HowPA, ; Velickovic2017GraphAN, ) to identify this anti-pattern.
-Leverage the network’s interpretability (Ying2019GNNExplainerGE, ; Yuan2021OnEO, ) to pinpoint problematic code segments and suggest optimizations, such as adding code to release references.
+code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Wang2022TestDrivenML, )).
+Applying advanced DL techniques to detect and fix Android performance issues is promising, although significant challenges remain in collecting sufficient performance issue dataset for model training.
+The emergence of large language models (LLMs) also presents new possibilities.
+LLMs, with their in-context learning capabilities, could address the limitations of traditional DL approaches by enabling models to detect and resolve performance issues with fewer examples and greater adaptability (Hou2023LargeLM, ; Liao2023A3CodGenAR, ; Si2024AST, ; Fan2023LargeLM, ).
+This opens an exciting direction for future research in Android performance analysis, promising to bridge the current gaps in existing techniques.
 4.3. Constructing Benchmark Dataset is an Urgent Need
-Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 11 (27.50%) of the 40 factors are covered by existing datasets.
-Even among the 11 factors, the data distribution is highly imbalanced.
-The number of activity samples is much larger than that of the other factors, which are more severe from a real-world perspective.
+Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 24 (33.33%) of the 72 factors are covered by existing datasets.
+Even among the 24 factors, the data distribution is highly imbalanced, as the smallest dataset contains only five apps, while the largest has 324.
 In addition, the number of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools.
-Moreover, another major issue is the limited reproducibility.
-Our review indicates that only a small number of primary studies share their datasets, making it difficult for future researchers to validate and evaluate tools.
+Another critical issue is reproducibility.
+Our study indicates that only a small number of primary studies share valid datasets, making it difficult for future researchers to validate and evaluate tools.
 Furthermore, these datasets lack real-time updates for the latest app, thus containing outdated samples.
 Thus, it is urgent to construct up-to-date, comprehensive benchmark datasets to support fair comparisons, covering as many real-world cases as possible.
-The code examples we summarized from the exploratory study can all be used to construct a comprehensive dataset, covering 35 factors.
-Another possible solution involves leveraging the findings of our exploratory study, wherein we identify six specific code patterns from real-world scenarios.
-These patterns can act as seeds for mutation tools, enabling the generation of additional data to construct datasets.
 5. THREATS TO VALIDITY
-Our study has several factors that could affect the validity of our findings in terms of literature review and large-scale exploration.
-For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works.
-To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 2.1.2, and used five well-known electronic databases.
-To minimize the risk of missing relevant works, we conducted backward and forward snowballing and cross-checked our results.
+Our study has several factors that could affect the validity of our findings in terms of large-scale exploration and literature review.
 For real-world data collection, the data filtering process may miss some relevant information.
@@
 In cases of disagreement, they discussed and resolved whether to include the data.
+For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works.
+To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 3.1.2, and used five well-known electronic databases.
+To minimize the risk of missing relevant work, we conducted backward and forward snowballing and cross-checked our results.
 6. Conclusion
-In this paper, we conducted a comprehensive empirical study to reveal the current state of Android performance analysis and the research gap between real-world challenges and existing techniques.
-Based on our in-depth analysis of 48 relevant studies and real-world issues across three major platforms (Google Play, Stack Overflow, GitHub), we discovered that (1) There are 6 types of performance issues and corresponding 40 factors; (2) Performance issues are sophisticated, with a single issue potentially stemming from multiple factors and a single factor causing multiple issues; (3) The root causes of data consumption issue are not clear; (4) The primary concerns of researchers, developers, and users are not aligned; (5) The majority of existing approaches primarily utilize static analysis; (6) There is a significant gap in existing tools’ ability to address all factors causing performance issues; (7) There is an urgent need to construct benchmark dataset to evaluate automatic tools; (8) DL-based and LLM-based approaches have shown promising results in addressing software engineering problems, suggesting their potential for tackling Android performance issues as well.
+In this paper, we conducted a comparative study to examine real-world Android performance issues encountered by users and developers, and to evaluate the extent to which current academic research addresses these practical challenges.
+By analyzing large-scale discussions from three popular crowd-sourced platforms on Android performance issues, alongside a review of 66 high-quality research papers, we identified several gaps between real-world challenges and the focus of academic research.
+Our main findings are as follows: (1) A taxonomy categorizes seven types of performance issues and 72 contributing factors.
+(2) Six common code patterns of these performance issues.
+(3) There is a misalignment in the primary concerns of researchers, developers, and users.
+(4) Only 42.86% of the real-world factors have been explored in academic research.
+(5) The majority of existing tools primarily utilize static analysis, dynamic analysis, and deep learning, with most tools detecting issues rather than fixing them.
+(6) Existing available tools target only 23.61% of identified factors.
+(7) Existing available datasets cover only 33.33% factors and lack sufficient data to support effective tool evaluation.
+Finally, we proposed future research directions to bridge the identified gaps.

### v3

@@
-Automatically Analyzing Performance Issues in Android Apps: How Far Are We?
+A Comparative Study of Android Performance Issues in Real-world Applications and Literature
 Abstract.
@@
 Unlike functionality issues, performance issues are more difficult to diagnose and resolve due to their complex root causes, which often emerge only under specific conditions or payloads.
-Although many efforts haven attempt to mitigate the impact of performance issues by developing methods to automatically identify and resolve them, it remains unclear if this objective has been fulfilled, and the existing approaches indeed targeted on the most critical performance issues encountered in real-world settings.
-To this end, we conducted a large-scale comparative study of Android performance issues in real-world applications and literature.
+Although many efforts have attempted to mitigate the impact of performance issues by developing methods to automatically identify and resolve them, it remains unclear if this objective has been fulfilled, and the existing approaches indeed targeted on the most critical performance issues encountered in real-world settings.
+To this end, we conduct a large-scale comparative study of Android performance issues in real-world applications and literature.
 Specifically, we started by investigating real-world performance issues, their underlying root causes (i.e., contributing factors), and common code patterns.
@@
 Our comparison results show a substantial divergence exists in the primary performance concerns of researchers, developers, and users.
-Among all the identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets.
+Among all the identified factors, 57.14% have not been examined in academic research, while a substantial 63.41% remain unaddressed by existing tools, and 70.73% lack corresponding datasets.
 This stark contrast underscores a substantial gap in our understanding and management of performance issues.
 Consequently, it is crucial for our community to intensify efforts to bridge these gaps and achieve comprehensive detection and resolution of performance issues.
-Android, Software Maintenance, Performance Issues
-conference: ; ; copyright: none
+Android, Software Maintenance, Performance Issues, Comparative Study
+$c$s: Software and its engineering
 1. Introduction
@@
 Therefore, it is essential for researchers to support developers in analyzing, detecting, and resolving Android performance issues (Habchi2018OnAL, ; Li2019DetectingAD, ; Li2022CombattingEI, ; Vsquez2015HowDD, ).
-Our research fellows have proposed several approaches to mitigate the impact of performance issues (Li2019CharacterizingAD, ; Banerjee2014DetectingEB, ; Song2021IMGDroidDI, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Vsquez2015HowDD, ; Hort2021ASO, ; Liu2019DroidLeaksAC, ; Habchi2019TheRO, ).
-For example, in the sense of static analysis, Li et al.
-(Li2019CharacterizingAD, ) developed TAPIR to analyze anti-patterns in app call graphs to detect inefficient image displays that can lead to memory consumption, energy consumption, and responsiveness issues.
-In addition, dynamic approaches, such as µDroid (Jabbarvand2017DroidAE, ), are proposed to detect energy waste due to resource under-utilization, wakelock misuse, and inefficient loops by monitoring abnormal energy consumption.
-However, given the scattered nature of current research in this field, it remains unclear whether performance issues have been comprehensively resolved.
-Specifically, it is unknown what the current state of Android app performance analysis is, whether current research is focused on key performance issues encountered in real-world settings, and to what extent existing methods effectively and comprehensively address these critical real-world challenges.
-Additionally, the capabilities of existing approaches are quite distinct, as confirmed by the fact of lacking a systematic taxonomy to map performance issues with their corresponding root causes (a.k.a, factors).
+To address performance issues in Android applications, researchers have explored the topic from multiple perspectives.
+On one hand, several studies aim to understand real-world performance issues by analyzing practical data sources (Kumari2024AnES, ; Noor2022EndUP, ; Das2016AQA, ; Vsquez2015HowDD, ).
+For instance, Noor et al.
+(Noor2022EndUP, ) examined 368,704 low-rated Google Play reviews and identified eight categories of user-reported performance issues, such as responsiveness delays and graphical glitches.
+Kumari et al.
+(Kumari2024AnES, ) analyzed 385 Stack Overflow posts to summarize developer-discussed performance issues and their causes, including UI/UX inefficiencies and problematic development practices.
+On the other hand, academic research has focused on performance-related code smells and design flaws (LeGoaer2022ecoCodeAS, ; Wu2023ASL, ; Fawad2024AndroidSC, ; Palomba2019OnTI, ).
+For example, Wu et al.
+(Wu2023ASL, ) reviewed 35 studies and categorized Android code smells into five types, including energy inefficiency and memory overuse.
+Similarly, Fawad et al.
+(Fawad2024AndroidSC, ) examined 79 papers and identified 237 Android code smell types, many of which are closely related to performance concerns, such as wakelock misuse and inefficient database queries.
+In addition to identifying performance issues, researchers have proposed various techniques to mitigate them (Li2019CharacterizingAD, ; Banerjee2014DetectingEB, ; Song2021IMGDroidDI, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Vsquez2015HowDD, ; Hort2021ASO, ; Liu2019DroidLeaksAC, ; Habchi2019TheRO, ).
+For example, in the area of static analysis, Li et al.
+(Li2019CharacterizingAD, ) developed TAPIR, which analyzes anti-patterns in app call graphs to detect inefficient image displays that can lead to memory consumption, energy consumption, and responsiveness issues.
+Meanwhile, dynamic approaches such as µDroid (Jabbarvand2017DroidAE, ) have been proposed to detect energy waste caused by resource underutilization, wakelock misuse, and inefficient loops by monitoring abnormal energy consumption.
+However, despite extensive research efforts, the study of Android performance issues remains fragmented.
+Most existing studies focus on a single data source or perspective, such as the user perspective from Google Play (Noor2022EndUP, ), the developer perspective from GitHub issues (Das2016AQA, ; Vsquez2015HowDD, ; Liu2014CharacterizingAD, ), or the researcher perspective from academic literature (Wu2023ASL, ; Fawad2024AndroidSC, ; Hort2021ASO, ).
+Moreover, these studies examine performance issues from different dimensions: some focus on symptoms (e.g., memory leaks (Banerjee2018EnergyPatchRR, ; Liu2016FixingRL, )), others on root causes (e.g., improper wakelock management (Pathak2012WhatIK, )), or specific code-level patterns (e.g., inefficient image loading (Song2021IMGDroidDI, )).
+Terminological inconsistencies further exacerbate this fragmentation, as similar issues are defined differently across studies, such as green bugs (Goar2020EnforcingGC, ), energy leaks (Carette2017InvestigatingTE, ), and energy smells (Iannone2020RefactoringAE, ).
+As a result, it remains unclear what the current state of Android app performance analysis is, whether current research is focused on key performance issues encountered in real-world settings, and to what extent existing methods effectively and comprehensively address these critical real-world challenges.
 Specifically, in this work, we formulate these concerns into two research questions (RQs) that we aim to answer through empirical evidence and explanatory results.
@@
 To answer these RQs, we began by conducting a large-scale analysis of real-world discussions on Android performance issues.
-Specifically, we collected 60,684 negative reviews from Google Play, 360,372 question threads from Stack Overflow, and 35,278 issues with 550,873 commits from GitHub to identify performance issues experienced by users and developers, analyze their root causes (i.e., contributing factors), and examine common code patterns.
+Specifically, we collected 60,684 negative reviews from Google Play, 749,067 question threads from Stack Overflow, and 16,977 issues with 344,922 commits from GitHub to identify performance issues experienced by users and developers, analyze their root causes (i.e., contributing factors), and examine common code patterns.
 Following this, we conducted a systematic literature review on Android performance analysis to investigate whether existing works indeed touch on the leading performance issues encountered in real-world settings.
-Through a comparison of academic findings with real-world data, we developed a comprehensive taxonomy covering seven categories of performance issues and 72 contributing factors (Table 1).
-This taxonomy allowed us to identify substantial gaps in current academic research on Android performance analysis.
+Through a comparison of academic findings with real-world data, we developed a comprehensive taxonomy covering seven categories of performance issues and 82 contributing factors (Table 1).
+This taxonomy (1) unifies inconsistent terminologies that fragmented prior studies, (2) explicitly distinguishes consequences from causes, providing a clearer mapping between real-world problems and their contributing factors, and (3) integrates user, developer, and researcher perspectives within a single framework.
+Together, these features enable systematic cross-perspective analysis and allow us to identify substantial gaps in current academic research on Android performance analysis.
 Specifically, it turns out there is a substantial divergence in primary concerns among researchers, developers, and users.
-Additionally, it is surprising to find that among all identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets.
+Additionally, it is surprising to find that among all identified factors, 57.14% have not been examined in academic research, while a substantial 63.41% remain unaddressed by existing tools, and 70.73% lack corresponding datasets.
 On top of these findings, we further propose future research directions to bridge these gaps.
-In this paper, we make the following major contributions:
-  • We conducted a large-scale exploratory study of real-world discussions on Android performance issues by analyzing crowd-sourced data from Google Play, Stack Overflow, and GitHub.
-This allowed us to identify the primary performance issues faced by users and developers, along with their root causes and common code patterns.
-  • We conducted a systematic literature review on Android performance analysis to examine the performance issues that researchers focus on, the tools developed to address these issues, and the datasets used to evaluate the effectiveness of these tools.
-  • By comparing literature with real-world data, we developed a taxonomy that categorizes seven types of performance issues and identifies 72 contributing factors, enabling us to assess how well current research objectives, tools, and datasets address real-world challenges.
-Based on these insights, we proposed future research directions to address the identified gaps.
+In this paper, we made the following key contributions:
+  • Cross-perspective analysis of real-world and academic sources.
+We collected and analyzed performance-related data from Google Play (user perspective), Stack Overflow and GitHub (developer perspective), and academic literature (researcher perspective), providing the first unified investigation across real-world and academic sources from multiple stakeholder perspectives.
+  • A unified consequence–cause taxonomy.
+We proposed a hierarchical taxonomy that systematically links observable performance issues (e.g., sluggish UI, memory bloat) to their contributing factors (e.g., inefficient image handling, unreleased wakelocks).
+The taxonomy covers seven types of performance issues and identifies 82 contributing factors, enabling consistent classification and comparison across sources.
+  • Quantitative gap analysis between research and practice.
+Using the taxonomy, we compared literature with real-world data to assess how well current research, tools, and datasets align with real-world concerns, and proposed directions to address the identified gaps.
+  • Artifact release for reproducibility. We have released our artifacts (our_repo, ), including source code, dataset, and summarized results, to support future research and community development.
 2. A Large-scale Exploration of Real-world Discussions on Android Performance Issues
@@
 Following common practices in analyzing large-scale online discussions (Guo2013CharacterizingAD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Banerjee2018EnergyPatchRR, ), we collected data from three popular crowd-sourced platforms: Google Play, Stack Overflow, and GitHub.
-Indeed, Google Play provides user feedback to assess the impact on end-users, Stack Overflow offers discussions that reveal common issues and solutions, and GitHub contains issues and commits histories that reflect the challenges developers encounter.
+Indeed, Google Play provides user feedback to assess the impact on end-users, Stack Overflow offers discussions that reveal common issues and solutions, and GitHub contains issues and commit histories that reflect the challenges developers encounter.
 Fig.
@@
 Each question thread on Stack Overflow typically reflects a real-world programming challenge, with the follow-up discussions offering additional insights, perspectives, and potential solutions.
-To understand developers’ perspectives on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers with code, resulting in 360,372 questions.
+To understand developers’ perspectives on Android performance issues, we crawled all question threads labeled with “android” with code, resulting in 749,067 questions.
 [Github - Issues & Commits.] Additionally, we collected performance issues and commits for open-source apps from GitHub, the most widely used platform for open-source development.
@@
 We chose F-Droid because it facilitates easier inspection of code patterns in open-source applications compared to closed-source ones.
-To gather a representative sample of Android apps, we crawled 17 categories of open-source Android apps along with their GitHub links from F-Droid.
-As a result, we successfully identified 4,228 Android apps and their GitHub repositories, which contained 35,278 issues and 550,973 commits.
+To gather a representative sample of Android apps, we crawled 17 categories of open-source Android apps along with their GitHub links from F-Droid, resulting in an initial dataset of 4,228 open-source Android apps.
+To ensure that the selected apps are not dummy projects, we introduced an additional validation step.
+Specifically, we verified whether each app also had a corresponding link to the Google Play Store.
+This yielded 1,643 apps that were both open-source and publicly distributed.
+From the repositories of these 1,643 apps, we extracted a total of 16,977 GitHub issues and 344,922 commits for our subsequent analysis.
 2.1.2. Keyword Filtering
@@
 To compile an accurate keyword list, we conducted a Google search for “Android performance”, reviewing the top 10 most relevant pages, which included 100 websites.
-Additionally, we examined several foundational papers on Android performance issues (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ).
+Additionally, we examined several foundational papers on Android performance issues (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Das2016AQA, ).
 From these sources, we obtained a list of 87 keywords about Android performance issues.
-The keyword list includes terms like “UI freeze”, “battery drain”, “memory leak”, “cpu utilization”, “app size” and so on.
-As a previously unknown performance issue might be caused by an existing factor and vice versa, existing keywords do not restrict us from discovering new performance issues and new factors.
-Using these keywords, we filtered 60,684 user reviews to 165, 360,372 Stack Overflow discussions to 893, 35,278 GitHub issues down to 317, and 550,973 GitHub commits to 780.
+The keyword list includes terms like “slow response”, “memory leak”, “battery drain”, “cpu utilization”, “app size” and so on.
+The full list of keywords is available in (our_repo, ).
+These keywords were used to match the titles of user reviews, Stack Overflow posts, GitHub issues, and commits.
+It is important to note that title matching served only as an initial filtering step.
+In subsequent manual analysis, we further examined the associated descriptions and code snippets to extract actual performance issues and discover contributing factors, which may not have been explicitly mentioned in the original keyword list.
+This deeper analysis allowed us to go beyond the scope of the predefined keywords and uncover previously undocumented insights.
+Using these keywords, we filtered 60,684 user reviews to 165, 749,067 Stack Overflow discussions to 2,158, 16,977 GitHub issues down to 149, and 344,922 GitHub commits to 558.
 2.1.3. Manual Checking.
@@
 Additionally, we excluded commits with multiple changes, such as fixes for various bugs or features where performance was a minor aspect, as this made identifying the root cause difficult.
-Ultimately, we obtained 114 user reviews, 336 stack overflow discussions, 47 GitHub issues, and 295 GitHub commits.
+Ultimately, we obtained 114 user reviews, 1,484 Stack Overflow discussions, 69 GitHub issues, and 222 GitHub commits.
 Two of the authors performed the aforementioned manual checking individually, for any disagreement, they discussed their findings until reaching a consensus.
 For any persisting disagreement, a senior author made the final decision.
-We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement, yielding scores of 0.944 for user reviews, 0.885 for GitHub issues, 0.895 for GitHub commits, and 0.871 for Stack Overflow posts, all of which signify substantial agreement.
+We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement, yielding scores of 0.944 for user reviews, 0.868 for GitHub issues, 0.885 for GitHub commits, and 0.873 for Stack Overflow posts, all of which signify substantial agreement.
 We then manually identified the Android performance issues within this data and present our findings in the following section.
-[…]
-[…]
-[…]
+$\text{Google Play - User Reviews: 60,684}\rightarrow\text{165}$ $\rightarrow\text{114}$
+$\text{Stack Overflow - Question Threads: 749,067}\rightarrow\text{2,158}$ $\rightarrow\text{1,484}$
+$\text{GitHub - Issues \& Commits: 16,977/344,922}\rightarrow\text{149/558}$ $\rightarrow\text{69/222}$
 2.2. Result Analysis
@@
 Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries.
-These issues appeared in 21.93% of all reported reviews.
+These issues appeared in 20.5% of all reported reviews.
 E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes.
@@
 Storage Consumption is often complained by users that the size of the app is too large, or the cache space consumed is too large when using the app.
-These issues appeared in 9.65% of all reported reviews.
+These issues appeared in 9.0% of all reported reviews.
 As mobile devices have limited storage, excessive storage consumption can lead to problems like a lack of available space for other apps or files, and potential app crashes.
@@
 I’m on a limited plan, now I have to buy extra data every month…”
-[Stack Overflow - Question Threads.] From the 336 question threads, we identified six categories of performance issues, as illustrated in the Stack Overflow pie chart in Fig.
+[Stack Overflow - Question Threads.] From the 1,484 question threads, we identified seven categories of performance issues, as illustrated in the Stack Overflow pie chart in Fig.
 2.
@@
 This discrepancy indicates that the performance issues developers focus on and those users are concerned about do not align.
+This misalignment can be attributed to the differing perspectives of users and developers when evaluating performance.
+From a user-centric perspective, performance is often assessed through perceptible outcomes such as smooth responsiveness, battery consumption, storage consumption, and internet data usage.
+In contrast, a developer-centric perspective typically involves technical diagnostics, focusing on issues such as memory usage, CPU utilization, and GPU rendering efficiency.
 Since user satisfaction directly impacts app market share, developers should prioritize addressing the issues most important to users.
@@
 2 show a similar distribution of performance issues across both data sources, with memory consumption, responsiveness, and storage consumption being the most frequently mentioned.
-Notably, while internet data usage is a niche issue primarily highlighted in user reviews and absent from Stack Overflow and GitHub commits, it also appears in GitHub issues, providing insights into its underlying causes.
+Notably, while internet data usage is a niche issue mainly mentioned in user reviews and absent from GitHub commits, it also appears in GitHub issues, which offer valuable insights into the underlying causes discussed by developers.
 In summary, the five performance issues that users mention in reviews are also discussed by developers on GitHub and Stack Overflow.
@@
 A substantial divergence exists between the focus areas of performance issues for users and developers: while users primarily emphasize concerns around responsiveness, developers’ discussions are predominantly oriented toward memory consumption.
-Table 1. Taxonomy of seven categories of performance issues and their contributing factors. ‘R-W.’ indicates real-world sources, while ‘Lit.’ refers to literature sources.
-Category Factor R-W. Lit. Problem Example
+Table 1. Taxonomy of seven categories of performance issues and their contributing factors.
+Category Factor Real-World Literature Problem Example
 Memory Consumption Loop Repeatedly reading the same value into the array in an infinite loop
@@
 Database Operation Each database query loads new data into memory without clearing the previous data
+Recursive Call Recursively sending requests rapidly creates excessive request queues and objects
 Activity An anonymous inner class strongly references the Activity, preventing it from being garbage collected
@@
 Configuration Gradle JVM heap size is insufficient during project builds, causing OOM errors
-Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
+Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, ; Das2016AQA, ; MazueraRozo2020InvestigatingTA, )
 Service Failing to destroy a dead service  (Song2019ServDroidDS, )
-Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Liu2019DroidLeaksAC, )
+Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Liu2019DroidLeaksAC, ; Das2016AQA, ; Wu2023ASL, )
 AsyncTask Failing to destroy AsyncTask when it holds active GUI components (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
-View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, )
+View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, ; Habchi2020AndroidCS, )
 Third-Party Library The API usage in the new version of the third-party library caused excessive memory usage (Rua2023ALE, )
 UI Rendering Reinflating views on each “getView” call without reusing them increases memory usage (Das2020CharacterizingTE, )
-Data Type Declaring a variable as static holds a strong reference to the previous task, preventing garbage collection (Das2020CharacterizingTE, )
+Data Type Declaring static variables retains strong references to previous tasks, preventing garbage collection (Das2020CharacterizingTE, ; Palomba2017LightweightDO, ; Habchi2020AndroidCS, )
 Obsolete Task Retaining obsolete parameters after updating view layouts  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
 Responsiveness Service Failing to move the notification to the foreground in time during service start, causing an ANR
-Inefficient Method Selecting a less efficient method introduces unnecessary delays in processing UI actions
+Inefficient Method Selecting a less efficient method introduces unnecessary delays in processing UI actions (Feng2019LearningPO, )
 Background Thread Rapidly creating numerous threads in the background to perform I/O on the same data
+Multi-Threads A thread gets blocked while waiting for a shared mutex held by another thread
 Cache Large cache causes frequent I/O operations, leading to lag
 Text Handling Repeatedly calling EditText.setText() to highlight all search results at once leads to unnecessary UI updates
-Resource Insufficient thread resource allocation causes slow response during simultaneous operations (Bhatt2020AutomatedRO, )
-Image 1) Decoding large images in the UI thread, 2) Loading large image, 3) dynamically adjusting content during activity-switching events (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
+Resource Insufficient thread resource allocation causes slow response during simultaneous operations (Bhatt2020AutomatedRO, ; Prestat2024DynAMICSAT, )
+Image Decoding large images in the UI thread (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, ; Das2016AQA, )
 Dynamic UI Loading Frequent layout inflations during scrolling (Liu2014CharacterizingAD, )
-UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Vsquez2015HowDD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Janssen2022OnTI, )
-Loop An infinite while loop in the button’s onClick() event makes the app unresponsive after clicking (Lyu2018RemoveRF, )
-Main Thread Performing heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ; Gmez2016MiningTR, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
-Activity Loading new resources during activity restarts may freeze the UI  (Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
-AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
-View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, ; Das2020CharacterizingTE, )
+HTTP Request Overlong connection timeout keeps page in loading state (Zhao2018LeveragingPA, ; Das2016AQA, ; Zhao2019SystematicallyTA, ; Das2016AQA, )
+UI Rendering Improper layout stacking causes redundant pixel redraws in the same frame (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Vsquez2015HowDD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Janssen2022OnTI, ; Nistor2014SunCatHD, ; MazueraRozo2020InvestigatingTA, ; Wu2023ASL, )
+Loop An infinite while loop in the button’s onClick() event makes the app unresponsive after clicking (Lyu2018RemoveRF, ; Palomba2017LightweightDO, ; MazueraRozo2020InvestigatingTA, )
+Main Thread Performing Heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ; Gmez2016MiningTR, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, ; Feng2019LearningPO, ; Das2016AQA, ; Wu2023ASL, )
+Activity Loading new resources during activity restarts may freeze the UI (Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, ; Callan2022ImprovingRO, )
+AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, ; Prestat2024DynAMICSAT, ; Feng2019LearningPO, ; Das2016AQA, )
+View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, ; Das2020CharacterizingTE, ; Prestat2024DynAMICSAT, ; Wu2023ASL, )
 Obsolete Task Downloading obsolete content occurs when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
@@
 Advertisements Loading ads while maintaining responsiveness to user actions (Petalotis2024AnES, )
-Message Backlogging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
-Storage Consumption File Handling Temp files from URIs aren’t cleared after upload
+Message Backloging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
+Storage Consumption Cache Residual data remains after the application is uninstalled
+Third-Party Library Unused third-party dependencies are packaged into the APK
+Resource Unused resources (such as audio files and assets) are packaged into the APK
+File Handling Temp files from URIs aren’t cleared after upload
 Redundant Code Similar functional pages and activities that could be reused, or unnecessary code that only occupies storage
@@
 Image PNG images compress less effectively than JPG, resulting in larger file sizes
-Energy Consumption Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Jabbarvand2017DroidAE, ; Nikzad2014APEAA, ; Wu2016Relda2AE, ; Vsquez2015HowDD, ; Bhatt2020AutomatedRO, ; Rua2023ALE, ; Liu2019DroidLeaksAC, )
-Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, )
+Energy Consumption Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Jabbarvand2017DroidAE, ; Nikzad2014APEAA, ; Wu2016Relda2AE, ; Vsquez2015HowDD, ; Bhatt2020AutomatedRO, ; Rua2023ALE, ; Liu2019DroidLeaksAC, ; Hort2021ASO, ; LeGoaer2022ecoCodeAS, ; Cruz2019CatalogOE, ; BernalCrdenas2015ImprovingEC, ; Goar2020EnforcingGC, ; Vsquez2017GEMMAMO, ; Morales2018EARMOAE, ; Das2016AQA, )
+Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, ; BernalCrdenas2015ImprovingEC, )
+Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, ; Jabbarvand2017DroidAE, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, ; LeGoaer2022ecoCodeAS, ; Cruz2019CatalogOE, ; Goar2020EnforcingGC, ; Iannone2020RefactoringAE, ; Vsquez2017GEMMAMO, ; Prestat2024DynAMICSAT, ; Palomba2017LightweightDO, ; MazueraRozo2020InvestigatingTA, ; Wu2023ASL, )
+HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, ; Li2014AnES, ; Goar2020EnforcingGC, ; Vsquez2017GEMMAMO, ; Das2016AQA, ; Wu2023ASL, )
+Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, ; Jabbarvand2017DroidAE, ; Li2014AnES, ; Palomba2019OnTI, ; Goar2020EnforcingGC, ; Das2016AQA, )
+Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, ; BernalCrdenas2015ImprovingEC, ; Goar2020EnforcingGC, ; Das2016AQA, )
+Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, ; LeGoaer2022ecoCodeAS, )
 Database Operation Frequently triggering database queries for auto-complete (Li2019DetectingAD, ; Lyu2018RemoveRF, ; Li2022CombattingEI, ; Lyu2017AnES, )
-Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, ; Jabbarvand2017DroidAE, ; Li2014AnES, ; Palomba2019OnTI, )
-Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, )
-Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, )
-HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, ; Li2014AnES, )
-Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
-Image Decoding identical image multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
-Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, )
-Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, ; Jabbarvand2017DroidAE, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
+Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, ; Feng2019LearningPO, ; Das2016AQA, )
+Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, ; Palomba2019OnTI, ; Vsquez2014MiningEA, ; Goar2020EnforcingGC, ; Iannone2020RefactoringAE, ; Vsquez2017GEMMAMO, ; Bangash2023CosteffectiveSF, ; Prestat2024DynAMICSAT, ; Morales2018EARMOAE, ; Palomba2017LightweightDO, ; Feng2019LearningPO, ; Das2016AQA, ; Wu2023ASL, )
+Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, ; Cruz2019CatalogOE, )
+Image Decoding identical image in multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, ; Das2016AQA, )
 Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
-UI Rendering 1) Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Janssen2022OnTI, )
-AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
-Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, ; Palomba2019OnTI, )
-Display Screen The OLED display consumes more energy when displaying light colors than dark ones (Li2014MakingWA, ; LinaresVsquez2018MultiObjectiveOO, ; Vsquez2015OptimizingEC, ; Wan2015DetectingDE, )
+UI Rendering Loading CSS for below-the-fold content during webpage loading wastes energy  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Janssen2022OnTI, ; Cruz2019CatalogOE, ; Wu2023ASL, )
+AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, ; BernalCrdenas2015ImprovingEC, ; Iannone2020RefactoringAE, )
+Display Screen The OLED display consumes more energy when displaying light colors than dark ones (Li2014MakingWA, ; LinaresVsquez2018MultiObjectiveOO, ; Vsquez2015OptimizingEC, ; Wan2015DetectingDE, ; Cruz2019CatalogOE, ; BernalCrdenas2015ImprovingEC, ; Goar2020EnforcingGC, ; Vsquez2017GEMMAMO, )
 Program Language Developing web applications using JavaScript consumes more energy than WebAssembly (Hasselt2022ComparingTE, ; Oliveira2017ASO, )
-Advertisements Frequent advertisement and analytics (A&A) service requests lead to energy waste (Cito2016BatteryawareTI, ; Petalotis2024AnES, )
+Advertisements Frequent advertisement and analytics (A&A) service requests lead to energy waste (Cito2016BatteryawareTI, ; Petalotis2024AnES, ; Vsquez2017GEMMAMO, )
 Machine Learning The inefficient machine learning algorithm on mobile devices is overly complex (McIntosh2018WhatCA, )
-View Drawing all items in a ListView without reusing existing data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
-CPU Usage Multi-Threads The music player mishandles metadata extraction, spawning thousands of threads, overloading the CPU.
+View Drawing all items in a ListView without reusing exising data  (Habchi2019TheRO, ; Couto2020EnergyRF, ; Feng2019LearningPO, )
+CPU Usage Multi-Threads The music player mishandles metadata extraction, spawning thousands of threads, overloading the CPU
 Loop An infinite loop repeatedly calls methods to handle images
 Image Glide consumes significant CPU resources to sequentially display images when playing GIFs
+Third-Party Library Third-party library’s bug causes a failure to properly release the resources used to display GIFs
+Background Thread Frequently perform read and write operations in a background thread
+Resource Use high sampling rate sensors such as the gyroscope
 UI Rendering Unnecessarily using transparency to render UI components (Lee2019ImprovingEE, )
 GPU Usage Third-Party Library Third-party library’s bug causes continuous redrawing when using GPU to render UI components
-UI Rendering Spin animation has unnecessary frequent refreshes, forcing GPU rendering on every frame (Lee2019ImprovingEE, )
+UI Rendering Spin animation has unnecessary frequent refreshes, forcing GPU rendering on every frame (Lee2019ImprovingEE, ; Habchi2020AndroidCS, ; Wu2023ASL, )
 Internet Data Usage Resource Downloading unwanted content without user permission, wasting internet resources
 Image Automatically loading animated GIFs without Wi-Fi enabled
+HTTP Request High-speed movement triggers frequent GPS updates, causing redundant API uploads
 Advertisements Automatically downloading embedded advertising services in the background
@@
 The absence of a universally accepted taxonomy results in many discussions targeting the same performance problems but using different descriptive terms (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ).
-Additionally, since one factor can lead to multiple performance issues consequences, and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system.
+Additionally, since one factor can lead to multiple performance issue consequences, and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system.
 To address this gap, our work distinguishes between the consequences and factors of performance issues.
-Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors.
-To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner.
-We now elaborate on the seven major categories, each with the corresponding 49 factors along with an example presented in Table 1.
+Specifically, we applied the thematic analysis methodology (Cruzes2011RecommendedSF, ; Das2016AQA, ) to analyze all the real-world data we collected, including GitHub issues and commits, Google Play Store user comments, and Stack Overflow posts.
+Using this approach, we classified performance issues according to their observed consequences, identified the underlying contributing factors, and established explicit links between them to construct a unified cross-source taxonomy.
+This process involved the following steps:
+  • (i) Familiarization with the data: Two authors independently reviewed all the real-world records to gain a thorough understanding of the content.
+  • (ii) Initial coding: After familiarization, the two authors conducted open labeling by assigning descriptive tags to each record.
+In total, they analyzed 114 Google Play user reviews, 1,484 Stack Overflow question threads, 69 GitHub issues, and 222 GitHub commits, as described in Section 2.1.3.
+Each fragment was labeled independently by both authors.
+At this stage, we identified 226 distinct labels.
+For example, a review, GitHub issue, Stack Overflow post, or commit might be annotated with tags such as memory usage inefficiency, memory leak, out of memory, image decoding without resizing, or image decoding in the main thread.
+  • (iii) Categorization of issues and factors: The initial labels were then organized into higher-level performance issues and contributing factors.
+For instance, memory usage inefficiency, memory leak, and out of memory were grouped under the broader issue category Memory Consumption.
+In contrast, tags such as decoding images without resizing, inefficient image format, and decoding identical images multiple times were consolidated into the broader Image factor.
+At this stage, we obtained 15 categories of performance issues and 96 categories of contributing factors.
+  • (iv) Review and refinement: All authors collaboratively reviewed and refined the labeled data to ensure accurate classification.
+During this iterative process, we held multiple discussions to merge semantically similar items, split overly broad categories, and discard irrelevant or redundant tags.
+For example, we combined the factors Drawing View and Bottom Sheet View into the broader View category.
+Finally, through this systematic and iterative process, we developed a taxonomy of seven major performance issues and 63 contributing factors.
+An illustrative example is presented in Table 1, where the Real-World column is marked with (✓).
 Memory Consumption.
 As discussed in Section 2.2.1, memory consumption is the most frequently highlighted performance issue among developers.
-Our real-world analysis revealed 20 factors that contribute to these memory consumption issues, which are listed under the Memory Consumption category in Table 1.
+Our real-world analysis revealed 21 factors that contribute to these memory consumption issues, which are listed under the Memory Consumption category in Table 1.
 These issues arise from a variety of factors, often resulting from inefficient coding practices that lead to excessive or unnecessary memory use.
@@
 Poor responsiveness leads to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc.
-In our analysis, we identified 14 key factors contributing to this issue.
+In our analysis, we identified 16 key factors contributing to this issue.
 Many real-world scenarios show that heavy operations associated with UI threads can slow responsiveness, such as decoding large images in the UI threads.
@@
 As the third most concerning performance issue for developers, inefficient storage consumption often leads to apps occupying more storage space than expected, which can limit the device’s ability to install other apps and store data.
-Through our real-world exploration, we identified four key factors contributing to storage consumption issues.
+Through our exploration, we identified seven key factors contributing to storage consumption issues.
 These factors typically arise from inefficient coding practices and poor data management, resulting in excessive or unnecessary storage usage.
@@
 However, developers have not prioritized addressing this issue to the same extent.
-Consequently, in our real-world analysis, we identified only two key factors contributing to energy consumption problems.
+Consequently, in our real-world analysis, we identified only six key factors contributing to energy consumption problems.
 The root causes generally stem from inefficient resource usage and unnecessary background operations that provide no tangible benefits to users.
@@
 This leads to a decline in overall system performance, affecting the normal operation of other apps and background services, and may even result in crashes, while also shortening the lifespan of the hardware and battery.
-Based on our analysis, we identified four distinct factors contributing to CPU usage issues.
+Based on our analysis, we identified seven distinct factors contributing to CPU usage issues.
 These issues typically stem from unnecessary or inefficient operations that waste CPU resources.
@@
 Internet Data Usage.
-Network resources on mobile devices are limited, especially when Wi-Fi is unavailable, requiring users to purchase cellular data.
-Inefficient data usage can lead to unnecessary expenses and create a negative user experience.
-From our analysis, we identified three key factors contributing to this issue.
-The root causes often involve the unnecessary consumption of network resources.
+This category primarily refers to the unnecessary or unauthorized use of network resources, regardless of whether the device is connected via Wi-Fi or cellular data, and regardless of whether the network behavior is caused by the app’s own logic or by integrated third-party APIs.
+Although such issues may occur under various network conditions, their negative impact is often more significant when users are not connected to Wi-Fi, as inefficient data usage can lead to unexpected expenses.
+From our analysis, we identified four key factors contributing to this issue.
+The root causes often involve unnecessary or unauthorized consumption of network resources, regardless of whether the device is connected via Wi-Fi or cellular data, and irrespective of whether the behavior is initiated by the app itself or by third-party libraries (e.g., advertising SDKs).
 For example, downloading unwanted content without user consent or automatically downloading embedded advertisements in the background can lead to excessive data usage.
 Answers to RQ1.2: Performance issues are inherently complex, often arising from an interplay of multiple contributing factors, where a single issue may stem from various causes, and, conversely, a single factor may amplify multiple issues.
-We identified 49 distinct factors contributing to the seven categories of performance issues.
+We identified 63 distinct factors contributing to the seven categories of performance issues.
 2.2.3. Common Code Patterns in Real-World
-After identifying key factors of performance issues, we analyzed Stack Overflow discussions, GitHub issues, and commits to uncover common code patterns associated with these issues.
-We summarized six types of common patterns as follows:
+After identifying the key contributing factors of performance issues, we analyzed Stack Overflow discussions, GitHub issues, and commits to uncover common code patterns associated with these issues.
+We found that the relationships among code patterns, performance issues, and contributing factors are inherently many-to-many: a single code pattern may involve multiple contributing factors and lead to various types of performance issues; similarly, a single factor may appear across multiple different code patterns.
+It is worth noting that a number of prior studies have also focused on identifying and summarizing performance-related code patterns in Android applications (Das2020CharacterizingTE, ; cruz2018using, ; Liu2014CharacterizingAD, ; Nistor2014SunCatHD, ; Guo2013CharacterizingAD, ; Song2021IMGDroidDI, ; Habchi2019TheRO, ; Li2019CharacterizingAD, ; Liu2016UnderstandingAD, ; Cruz2019CatalogOE, ).
+For example, the well-known DrawAllocation issue typically refers to performing memory allocation within UI rendering methods such as onDraw(), and has been widely discussed in the literature (Das2020CharacterizingTE, ; cruz2018using, ) and flagged by the Android Lint tool.
+This specific issue often involves contributing factors such as Loop, Object Creation, and UI Rendering.
+In our study, however, DrawAllocation is considered a concrete instance of a broader code pattern category we term UI Operations.
+More generally, each of our identified code pattern categories encompasses a broader set of real-world cases, which may be composed of different combinations of contributing factors.
+Many specific patterns reported in prior research can be viewed as subtypes under these broader categories.
+We present six representative categories of common code patterns below.
 1const CopyTextOption = ({bottomSheetId, postMessage, sourceScreen}: Props) => {
@@
 These errors can result in severe performance consequences, such as ANR and delays.
-For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error.
+For example, passing a null context (a bad practice in parameter value passing) to PendingIntent may result in an Application Not Responding (ANR) error.
+This pattern may involve multiple contributing factors.
+Specifically, if the context is derived from an Activity or Service with an invalid lifecycle state, it can prevent the intent from being properly dispatched or executed, thereby blocking the Main Thread and eventually triggering an ANR.
 Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5).
 In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.
+This case mainly involves the Main Thread and UI Rendering factors.
 1public class MySingleton() {
@@
 This prevents it from being garbage collected, leading to a memory leak.
+This case mainly involves the Singleton, Activity, and Service factors.
 1public class MyTextView extends TextView {
@@
 This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
+This case mainly involves the Object Creation factor.
 1 ...
@@
 Large-Scale Data.
-Large-scale data is also a culprit for performance issues.
+Large-scale data is also a common cause of performance issues, since operations on such data often exceed the capacity of mobile devices with limited resources.
 Loading large images without proper caching can cause excessive memory usage and slowdowns.
@@
 If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
+This case mainly involves the Massive Data and File Handling factors.
 1public class ConcentricCircularView extends View {...
@@
 This causes the application to slow down and increases the computational load.
+This case mainly involves contributing factors such as Loop and UI Rendering.
 Other Patterns.
 Besides the previously mentioned patterns, other types of patterns also warrant attention, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager on the main thread, etc.
+We made the full list publicly available in our artifact: (our_repo, ).
 Answers to RQ1.3 We identified six prevalent code patterns that commonly contribute to performance issues: API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
+2.3. Practitioner Validation of the Taxonomy
+To further validate whether the performance issues and contributing factors identified in our taxonomy reflect practical concerns, we conducted a survey with Android practitioners.
+2.3.1. Method
+We designed a questionnaire consisting of the following parts:
+  • Q1a: What is your current role related to Android software development?
+  • Q1b: What is your age range? (23 or below, 24–27, 28 or above)
+  • Q1c: How many years of experience do you have in Android development or related research?
+  • Q2: This section presents seven common categories of performance issues.
+For each category, please indicate the extent to which it concerns you during development.
+Responses were recorded on a 5-point Likert scale: 1 = not concerned at all, 2 = slightly concerned, 3 = moderately concerned, 4 = concerned, 5 = very concerned.
+  • Q3: This section presents each performance issue category along with its contributing factors and an example illustrating how each factor can lead to the issue.
+For each category, you are asked to review the associated factors and select all those that they consider a concern or believe may potentially affect performance during development.
+This question used a multiple-choice format, allowing participants to select zero or more factors that reflect their concerns or are potentially relevant to their development experience.
+Among them, Q2 was designed to validate the relevance of the seven performance issue categories by measuring how strongly participants with Android development experience perceive each category as a practical concern.
+Q3 was designed to validate the comprehensiveness of the contributing factors by examining whether participants recognize these factors as relevant to their own development experience, thereby assessing the alignment between our taxonomy and real-world practice.
+The survey was conducted online and remained open for two weeks.
+We collected anonymous responses to encourage honest feedback, and all participants provided informed consent before answering the questions.
+Figure 3. Average concern scores for different categories of Android performance issues.
+2.3.2. Results
+We received 21 valid responses.
+Participants’ ages fell into three ranges: 23 or below (4.8%), 24–27 (76.2%), and 28 or above (19.0%), with the majority in the 24–27 range.
+Most participants (81%) are casual developers with Android development experience, while the remaining 19% are professional developers.
+On average, participants reported 4.19 years of experience in Android development or related research.
+For Q2, the median concern score across all seven performance issue categories was 4, indicating that participants widely recognize these issues as important in practice.
+Fig.
+3 shows the average concern level for each performance issue category.
+The most concerning issue was Memory Consumption (mean: 4.95), while the least concerning was Internet Data Usage, which aligns with the patterns we observed in real-world discussions.
+For Q3, participants selected an average of 54.19 out of 63 contributing factors, resulting in an overall alignment rate of 86.16%.
+Notably, 18 factors were selected by all participants, such as the UI Rendering factor contributing to GPU Usage issues and the Resource factor contributing to Energy Consumption problems.
+This suggests strong agreement between our taxonomy and developers’ real-world experiences.
+In contrast, some factors received fewer selections.
+For example, the Third-party Library factor contributing to CPU Usage issues and the Singleton factor contributing to Memory Consumption issues were each selected by only 11 participants.
+This may be because these factors occur less frequently or are less visible during development, leading to them being unintentionally overlooked by developers.
+Overall, the survey results support the validity of our taxonomy and confirm that the performance issues and contributing factors identified in our study accurately reflect real-world development concerns, thereby reinforcing the findings presented in RQ1.
 3. Studies on Performance Issues
 In the above Section 2, we systematically examined real-world Android performance issues, their root causes, and common code patterns.
-However, it remains unclear whether cutting-edge literature targets the same practical issues, whether existing techniques can effectively resolve them, and whether current datasets can comprehensively evaluate the techniques’ effectiveness.
+However, it remains unclear whether cutting-edge literature targets the same practical issues (RQ2.1), whether existing techniques can effectively resolve them (RQ2.2), and whether current datasets can comprehensively evaluate the techniques’ effectiveness (RQ2.3).
 To answer these questions, we conducted a systematic literature review (SLR) to assess the current state of Android performance analysis.
@@
 4 illustrates the working process which comprises 5 steps.
+Here, we clarify that this study focuses on performance issues resulting from inefficient coding practices within Android applications, rather than problems related to Android device hardware or the operating system.
 Figure 4. The working process of our systematic literature review.
@@
 To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library.
-From this initial search, we retrieved a total of 11,442 papers.
+From this initial search, we retrieved a total of 6,649 papers from IEEE, 1,507 from ACM, 537 from Springer, 1,969 from ScienceDirect, and 1,008 from Wiley.
+Since many papers appeared across multiple libraries, we removed duplicates based on paper titles, resulting in 11,574 unique papers.
 However, many of these studies were outside the scope of our focus.
@@
 3.1.3. Paper Exclusion
-To gather relevant papers as possible, we initially considered all returned results from our search.
+To gather as many relevant papers as possible, we initially considered all returned results from our search.
 However, not all papers aligned with our focus on performance issues in mobile apps caused by poor coding practices, as discussed in Section 1.
@@
 Specifically, we applied the following exclusion criteria to streamline this process:
-  1. (1)
-    Short papers (i.e., less than six pages in double-column format or 11 pages in single-column format) were excluded.
   1. (1)
@@
 For example, Hypnos (Jindal2013HypnosUA, ) is used for detecting sleep conflicts in smartphone device drivers, which is outside our scope and focused on mobile app issues stemming from poor coding practices by developers.
-After applying these exclusion criteria, we retained 49 papers that are closely related to the topic of performance issues analysis at the Android mobile level.
+After applying these exclusion criteria, we retained 60 papers that are closely related to the topic of performance issues analysis at the Android mobile level.
 3.1.4. Back/Forward Snowballing
-After initially identifying relevant papers (i.e., 49 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature.
+After initially identifying relevant papers (i.e., 60 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature.
 The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources.
@@
 We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified.
-Eventually, we were able to locate a total of 66 papers.
+Eventually, we were able to locate a total of 85 papers.
 3.1.5. Cross Check
 Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies.
-Ultimately, we collected 66 papers to assess the current state of research and identify gaps in addressing real-world performance issues.
+Ultimately, we collected 85 papers to assess the current state of research and identify gaps in addressing real-world performance issues.
+As shown in Fig.
+5(a), the majority of these studies were published in conference proceedings (62 out of 85, accounting for 72.94%).
+Moreover, we observe a consistent research interest in this topic, with papers published every year since 2012.
+All papers involved in the SLR process are available in our replication package under the Literature_Review/data directory (our_repo, ).
 3.2. Results
@@
 Holistic Taxonomy.
-Through reviewing 66 research papers, we identified five types of performance issues and 44 contributing factors, marking each in the Lit.
-column of Table 1 with a checkmark (✓).
-By incorporating insights from both real-world discussions and research papers, we developed a final taxonomy encompassing seven types of performance issues and 72 factors.
+Through reviewing 85 research papers, we identified five types of performance issues and 45 contributing factors, marking each in the Literature column of Table 1 with a checkmark (✓).
+By incorporating insights from both real-world discussions and research papers, we developed a final taxonomy encompassing seven types of performance issues and 82 factors.
 By examining these issues and their contributing factors from both real-world data and academic findings, we identified significant gaps between research focus and real-world challenges as follows.
-Figure 4. Annual publications on each performance issue.
+(a) Number of publications per year by venue type.
+(b) Annual publications on each performance issue.
+Figure 5. Distribution of selected studies by venue type and performance issue.
 Gaps between research focus and real-world performance issues.
 As shown in Fig.
-4, we identified five key performance issues from literature: energy consumption, responsiveness, memory consumption, CPU usage, and GPU usage.
+5(b), we identified five key performance issues from literature: energy consumption, responsiveness, memory consumption, GPU usage, and CPU usage.
 These performance issues are also prevalent in real-world scenarios, indicating that researchers have made significant efforts to address concerns that are relevant to both developers and users.
 Despite these efforts, there remains a noticeable imbalance in focus.
-Among the five issues, energy consumption received the most attention, with 53 of the examined papers (80.30%) focusing on this issue, and this trend is likely to continue.
+Among the five issues, energy consumption received the most attention, with 69 of the examined papers (81.18%) focusing on this issue, and this trend is likely to continue.
 However, despite this focus on energy consumption, Section 2.2 revealed that users are more concerned with responsiveness, while developers prioritize memory consumption.
 This highlights a discrepancy between the performance issues prioritized by researchers, developers, and users.
-Furthermore, two performance issues prevalent in real-world scenarios have not been explored in literature: storage consumption and internet data usage.
+Moreover, two performance issues that are prevalent in real-world applications, storage consumption and internet data usage, have been relatively underexplored in the academic literature (Bijlani2021WhereDM, ; Zhang2012HowEA, ; Liu2016WhoMM, ), highlighting the need for more focused attention in future research.
 These gaps suggest that researchers should place greater emphasis on the performance issues that are of primary concern to both users and developers in real-world apps.
 Gaps in academic research on real-world performance issue factors.
-By analyzing the cases where both the R-W.
-and Lit.
-columns are marked with a checkmark (✓), we found that only 21 out of 49 factors discussed in real-world settings (42.86%) have been studied in literature.
+By analyzing the cases where both the Real-World and Literature columns are marked with a checkmark (✓), we found that only 27 out of 63 factors discussed in real-world settings (42.86%) have been studied in literature.
 It indicates that the majority of performance issues encountered in real-world scenarios remain underexplored in academic research.
-For instance, memory consumption, which is the top concern for developers, real-world data shows 20 contributing factors, but only 8 (40.00%) have been explored in literature.
-Similarly, for responsiveness, the top concern for users, 14 real-world factors were identified, but only nine (64.29%) have been studied in literature.
-Less common issues like CPU usage and GPU usage have seen even less attention, with only 25.00% and 50.00% of their respective real-world root causes being studied in literature.
+For instance, memory consumption, which is the top concern for developers, real-world data shows 21 contributing factors, but only 8 (38.10%) have been explored in literature.
+Similarly, for responsiveness, the top concern for users, 16 real-world factors were identified, but only 10 (62.50%) have been studied in literature.
+Less common issues like CPU usage and GPU usage have seen even less attention, with only 14.29% and 50.00% of their respective real-world root causes being studied in literature.
 These gaps suggest that many root causes of real-world performance issues still remain unexplored by researchers.
@@
 The presence of research-only factors not seen in real-world discussions.
-By examining rows where the R-W.
-column is unmarked (✗) but the Lit.
-column is marked (✓), we identified 23 factors that have been explored in literature but not observed in the real world.
-Specifically, we found one unseen factor in memory consumption, four in responsiveness, and 18 in energy consumption.
+By examining rows where the Real-World column is unmarked (✗) but the Literature column is marked (✓), we identified 19 factors that have been explored in literature but not observed in the real world.
+Specifically, we found one unseen factor in memory consumption, four in responsiveness, and 14 in energy consumption.
 This suggests that literature may uncover factors that did not appear in our real-world exploration, thereby further enriching the taxonomy.
-Notably, most of these unseen factors pertain to energy consumption, accounting for 90.00% of the factors explored in literature but not observed in the real-world.
+Notably, most of these unseen factors pertain to energy consumption, accounting for 73.68% of the factors explored in literature but not observed in the real-world.
 For example, Li et al.
@@
 This indicates a pronounced disconnection between academic research and the practical challenges faced by practitioners.
-Answers to RQ2.1: Through an extensive integration of real-world explorations and thorough literature review, we present a comprehensive taxonomy encapsulating seven primary types of performance issues and 72 contributing factors.
-A noteworthy finding shows that only 42.86% factors contributing to performance issues have been examined in academic research.
+Answers to RQ2.1: Through an extensive integration of real-world explorations and thorough literature review, we present a comprehensive taxonomy encapsulating seven primary types of performance issues and 82 contributing factors.
+A noteworthy finding shows that only 42.86% of factors contributing to performance issues have been examined in academic research.
 A significant misalignment exists in the primary concerns of researchers, developers, and users: researchers concentrate on energy consumption, developers prioritize memory usage, and users are chiefly concerned with responsiveness.
@@
 Table 3 presents the list of available automatic approaches collected from our examined papers.
-Among the 66 papers examined, eight (12.12%) provided publicly accessible tools designed to detect specific factors contributing to particular performance issues.
+Among the 85 papers examined, 14 (16.47%) provided publicly accessible tools designed to detect specific factors contributing to particular performance issues.
 This small fraction can be attributed to several reasons: some papers lack accessible artifacts, others provide only datasets without the corresponding tool code, and in some cases, the links to the artifacts are invalid.
-This indicates a significant gap between automatic tools with their practical usage to within our community.
-In addition, we found that existing tools target only 17 out of the 72 factors in our taxonomy (Table 1), covering just 23.61%.
+This indicates a significant gap between automatic tools and their practical usability within the community.
+In addition, we found that existing tools target only 30 out of the 82 factors in our taxonomy (Table 1), covering just 36.59%.
 This indicates that the majority of factors contributing to the seven primary types of performance issues are overlooked by SOTA tools, highlighting a significant gap in addressing these issues comprehensively.
@@
 Table 3. A summary of automatic approaches used for Android mobile performance analysis. Static denotes Static Analysis, Dynamic denotes Dynamic Analysis, and DL denotes Deep Learning methods.
-Tool Approach Technique Stage Performance Issue Target Factor Link
-VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Memory Consumption Energy Consumption Activity Resource  (Lu2022DetectingRUTool, )
-AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
-TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
-ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural Code Analysis Detect Memory Consumption Energy Consumption Service  (Song2019ServDroidDSTool, )
-Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Energy Consumption Responsiveness Database Operation Loop  (Lyu2018RemoveRFTool, )
-PlumbDroid (Bhatt2020AutomatedRO, ) Static Control-Flow Analysis Patch Generation Detect Fix Responsiveness Resource  (Bhatt2020AutomatedROTool, )
-µDroid (Jabbarvand2017DroidAE, ) Dynamic Mutation Testing Power Trace Analysis Detect Energy Consumption Resource Wakelock Loop  (Jabbarvand2017DroidAETool, )
-ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Energy Consumption Resource  (Jabbarvand2020AutomatedCOTool, )
+Tool Approach Technique Stage
+Performance Issue
+Target Factor
+Link
+VALA (Lu2022DetectingRU, ) Static
+Dataflow Analysis
+Symbolic Execution
+Detect
+Memory Consumption
+Energy Consumption
+Activity
+Resource
+ (Lu2022DetectingRUTool, )
+AsyncChecker(Pan2020StaticAC, ) Static
+Code Pattern Analysis
+Interprocedural-
+Code Analysis
+Detect
+Memory Consumption
+Energy Consumption
+Responsiveness
+AsyncTask  (Pan2020StaticACTool, )
+TAPIR (Li2019CharacterizingAD, ) Static
+Code Pattern Analysis
+Detect
+Memory Consumption
+Energy Consumption
+Responsiveness
+Image  (Li2019CharacterizingADTool, )
+ServDroid (Song2019ServDroidDS, ) Static
+Code Pattern Analysis
+Interprocedural Code Analysis
+Detect
+Memory Consumption
+Energy Consumption
+Service  (Song2019ServDroidDSTool, )
+Lyu et al. (Lyu2018RemoveRF, ) Static
+Dataflow Analysis
+Program Transformation
+Detect
+Fix
+Energy Consumption
+Responsiveness
+Database Operation
+Loop
+ (Lyu2018RemoveRFTool, )
+PlumbDroid (Bhatt2020AutomatedRO, ) Static
+Control-Flow Analysis
+Patch Generation
+Detect
+Fix
+Responsiveness Resource  (Bhatt2020AutomatedROTool, )
+µDroid (Jabbarvand2017DroidAE, ) Dynamic
+Mutation Testing
+Power Trace Analysis
+Detect Energy Consumption
+Resource
+Wakelock
+Loop
+ (Jabbarvand2017DroidAETool, )
+ACETON (Jabbarvand2020AutomatedCO, )
+DL
+Dynamic
+Deep Learning
+Feature Engineering
+Runtime Monitoring
+Detect Energy Consumption Resource  (Jabbarvand2020AutomatedCOTool, )
+Leafactor (Cruz2017LeafactorIE, ) Static
+Code Pattern Analysis
+Program Transformation
+Detect
+Fix
+Energy Consumption
+Wakelock
+View
+UI Rendering
+ (LeafactorTool, )
+PAPRIKA (Hecht2015DetectingAI, ) Static
+Code Pattern Analysis
+Detect
+Memory Consumption
+Responsiveness
+Energy Consumption
+Inefficient Method
+Data Type
+Activity
+ (PAPRIKATool, )
+ecoCode (LeGoaer2022ecoCodeAS, ) Static
+Rule-Based Detection
+XML/Gradle Inspection
+Detect Energy Consumption
+Resource
+Wakelock
+Redundant Frames
+Service
+Loop
+HTTP Request
+ (ecoCodeTool, )
+aDoctor (Iannone2020RefactoringAE, )
+Static
+Code Pattern Analysis
+Program Transformation
+Detect
+Fix
+Energy Consumption
+Wakelock
+Inefficient Method
+Data Type
+ (adoctorTool, )
+Das et al. (Das2020CharacterizingTE, )
+Static
+Rule-Based Detection
+AST Analysis
+Detect
+Energy Consumption
+Memory Consumption
+Responsiveness
+UI Rendering
+Data Type
+Main Thread
+Inefficient Method
+View
+Wakelock
+Activity
+ (AndroidLintTool, )
+Vásquez et al. (Vsquez2015HowDD, )
+Dynamic
+Runtime Monitoring
+Detect
+Responsiveness
+Main Thread
+ (StrictModeTool, )
 Static Analysis.
 Most existing approaches are based on static analysis to detect or fix performance issues, as shown in Table 3.
-This is intuitive since static analysis involves reasoning about program behavior without execution, making it lightweight to deploy Specifically, taint analysis, symbolic execution (Lu2022DetectingRU, ), code pattern analysis (Pan2020StaticAC, ; Li2019CharacterizingAD, ; Song2019ServDroidDS, ), and data flow analysis (Lyu2018RemoveRF, ) are commonly employed in rule-based static analysis as many performance issues arise from common code patterns.
+This is intuitive since static analysis involves reasoning about program behavior without execution, making it lightweight to deploy.
+Specifically, taint analysis, symbolic execution (Lu2022DetectingRU, ), code pattern analysis (Pan2020StaticAC, ; Li2019CharacterizingAD, ; Song2019ServDroidDS, ), and data flow analysis (Lyu2018RemoveRF, ) are commonly employed in rule-based static analysis as many performance issues arise from common code patterns.
 Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ).
@@
 Dynamic Analysis.
-There are only two studies that primarily employ dynamic analysis approaches to detect performance bugs.
+There are only three studies that primarily employ dynamic analysis approaches to detect performance bugs.
 This result aligns with expectations because dynamic analysis involves executing apps in a controlled environment and is unable to cover all possible paths (dong2020timemachine, ).
 However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ).
-As shown in Table 3, both dynamic analysis tools target energy consumption issues, particularly unnecessary energy waste such as resource under-utilization, wakelock misuse, and inefficient loops, which are detectable through abnormal energy consumption monitoring.
+As shown in Table 3, two dynamic analysis tools target energy consumption issues, particularly unnecessary energy waste such as resource under-utilization, wakelock misuse, and inefficient loops, which are detectable through abnormal energy consumption monitoring.
+One dynamic analysis tool, StrictMode (StrictModeTool, ), targets responsiveness issues, particularly those caused by executing heavy operations on the main thread, which can lead to UI freezes and slow interactions.
 We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
@@
 The limited adoption of applying DL to Android performance analysis may due to challenges in collecting large datasets that are sufficient to train DL models on performance issues and their corresponding fixes.
-Answers to RQ2.2: We identified eight publicly available tools, primarily using static analysis, dynamic analysis, and deep learning approaches.
-Existing tools target only 17 of the 72 identified factors to performance issues (23.61%), leaving a substantial number of factors unexamined.
+Answers to RQ2.2: We identified 14 publicly available tools, primarily using static analysis, dynamic analysis, and deep learning approaches.
+Existing tools target only 30 of the 82 identified factors to performance issues (36.59%), leaving a substantial number of factors unexamined.
 Most tools focus on detecting performance issues while neglecting their resolution.
@@
 Dataset Year # APP Type Performance Issue Target Factor Link
-Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Energy Consumption Responsiveness Database Operations Loop  (Lyu2018RemoveRFTool, )
-Vásquez et al. (LinaresVsquez2018MultiObjectiveOO, ) 2018 5 Real-World Energy Consumption Display Screen  (LinaresVsquez2018MultiObjectiveOOTool, )
-Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World Memoy Consumption Energy Consumption Responsiveness UI Rendering View Inefficient Method  (Habchi2019TheROData, )
-Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
-AsyncBench (Pan2020StaticAC, ) 2020 69 Micro Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
-Liu et al. (Liu2019DroidLeaksAC, ) 2020 34 Real-World Memory Consumption Energy Consumption Resource  (Liu2019DroidLeaksACTool, )
-Das et al. (Das2020CharacterizingTE, ) 2020 316 Real-World Memory Consumption Energy Consumption UI Rendering Data Type Inefficient Method View Wakelock  (Das2020CharacterizingTETool, )
-Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Memory Consumption Energy Consumption Resource  (Lu2022DetectingRUTool, )
-Janssen et al. (Janssen2022OnTI, ) 2022 40 Real-World Energy Consumption Responsiveness UI Rendering  (Janssen2022OnTITool, )
-Petalotis et al. (Petalotis2024AnES, ) 2024 9 Real-World Energy Consumption Responsiveness Advertisements  (Petalotis2024AnESTool, )
+Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World
+Energy Consumption
+Responsiveness
+Database Operations
+Loop
+ (Lyu2018RemoveRFTool, )
+Vásquez et al. (LinaresVsquez2018MultiObjectiveOO, ) 2018 5 Real-World
+Energy Consumption
+Display Screen
+ (LinaresVsquez2018MultiObjectiveOOTool, )
+Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World
+Memory Consumption
+Energy Consumption
+Responsiveness
+UI Rendering
+View
+Inefficient Method
+ (Habchi2019TheROData, )
+Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World
+Memory Consumption
+Energy Consumption
+Responsiveness
+Image  (Li2019CharacterizingADTool, )
+AsyncBench (Pan2020StaticAC, ) 2020 69 Micro
+Memory Consumption
+Energy Consumption
+Responsiveness
+AsyncTask  (Pan2020StaticACTool, )
+Liu et al. (Liu2019DroidLeaksAC, ) 2020 34 Real-World
+Memory Consumption
+Energy Consumption
+Resource
+ (Liu2019DroidLeaksACTool, )
+Das et al. (Das2020CharacterizingTE, ) 2020 316 Real-World
+Memory Consumption
+Energy Consumption
+UI Rendering
+Data Type
+Inefficient Method
+View
+Wakelock
+ (Das2020CharacterizingTETool, )
+Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World
+Memory Consumption
+Energy Consumption
+Resource
+ (Lu2022DetectingRUTool, )
+Janssen et al. (Janssen2022OnTI, ) 2022 40 Real-World
+Energy Consumption
+Responsiveness
+UI Rendering
+ (Janssen2022OnTITool, )
+Petalotis et al. (Petalotis2024AnES, ) 2024 9 Real-World
+Energy Consumption
+Responsiveness
+Advertisements
+ (Petalotis2024AnESTool, )
+Cruz et al. (Cruz2019CatalogOE, ) 2019 1,027 Real-World Energy Consumption
+Display Screen
+UI Rendering
+Wakelock
+Resource
+Obsolete Task
+ (Cruz2019CatalogOEDataset, )
+Prestat et al. (Prestat2024DynAMICSAT, ) 2024 5 Real-World
+Energy Consumption
+Responsiveness
+Memory Consumption
+Wakelock
+Inefficient Method
+Resource
+AsyncTask
+View
+ (Prestat2024DynAMICSATDataset, )
 3.2.3. Result - Dataset
 To collect datasets, several types of papers need to be included: 1) Tool papers that construct datasets to evaluate their designed tools; 2) Empirical studies that collect specific datasets to analyze performance issues, where labeled data can be used to support evaluation.
-After examining all the public links or websites within the 66 identified studies, we identified 10 (15.15%) unique, available, and open-sourced datasets, as shown in Table 4.
-Notably, we conducted the examination in October 2024.
+After examining all the public links or websites within the 85 identified studies, we identified 12 (14.11%) unique, available, and open-sourced datasets, as shown in Table 4.
+Notably, we conducted the examination in Jun 2025.
 Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison.
 In addition, the number of factors covered by these datasets is quite limited.
-Out of the 72 factors identified in our taxonomy, only 24 (33.33%) are addressed by existing datasets.
+Out of the 82 factors identified in our taxonomy, only 24 (29.27%) are addressed by existing datasets.
 This indicates a significant need for constructing more comprehensive datasets.
 Moreover, there is a considerable imbalance in dataset size.
-The smallest dataset contains only five apps, while the largest has 324, with an average of 105.7 apps per dataset.
+The smallest dataset contains only five apps, while the largest has 1,027, with an average of 174.1 apps per dataset.
 This imbalance makes it unrealistic for any single dataset to cover all potential issues comprehensively.
@@
 Therefore, more effort is needed to create datasets that are not only comprehensive but also more representative of the diverse range of performance issues.
-Answers to RQ2.3: We identified 10 open-source datasets for evaluating performance issues; however, they encompass only 24 out of 72 factors (33.33%), highlighting a significant gap in tool evaluation.
-These datasets suffer from inadequate sample sizes, with the smallest containing only five apps and the largest consisting of 324.
+Answers to RQ2.3: We identified 12 open-source datasets for evaluating performance issues; however, they encompass only 24 out of 82 factors (29.27%), highlighting a significant gap in tool evaluation.
+These datasets suffer from inadequate sample sizes, with the smallest containing only five apps and the largest consisting of 1,027.
 4. Future Trends
@@
 This suggests that future research should increase focus on responsiveness and memory issues, as these have a more immediate impact on both end-users and developers.
-We also found two categories of performance issues—storage consumption and internet data usage—that remain unexplored in academic research, indicating potential areas for further investigation.
+We also found two categories of performance issues, storage consumption and internet data usage, that remain unexplored in academic research, indicating potential areas for further investigation.
 Additionally, only 42.86% of the factors contributing to performance issues observed in real-world settings have been explored in academic research, remaining nearly 60% still awaiting for researchers to investigate.
-Interestingly, there are 23 factors among the 72 factors lead to performance issues are only been explored in literature but not observed in the real-world scenarios.
+Interestingly, there are 19 factors among the 82 factors that lead to performance issues that have only been explored in literature but not observed in the real-world scenarios.
 This suggests that the issues currently discussed in academic research may not be of real concern to users or developers in real-world settings.
@@
 4.2. Automatic Performance Analysis Tools are Still in an Early Stage
-Our study revealed that only 17 out of 72 factors (23.61%) are addressed by state-of-the-art (SOTA) tools, with most tools focusing on detection rather than resolution.
-In fact, only two out of the existing tools are dedicated to fixing performance issues.
+Our study revealed that only 30 out of 82 factors (36.59%) are addressed by state-of-the-art (SOTA) tools, with most tools focusing on detection rather than resolution.
+In fact, only four out of the existing tools are dedicated to fixing performance issues.
 Although static, dynamic, and deep learning (DL) techniques have shown effectiveness in research experiments, applying these methods to holistically detect and fix performance issues in practice remains an unsolved challenge.
@@
 Thus, developing effective tools against performance issues will be a noteworthy research topic in the future.
-In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e.
-code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Wang2022TestDrivenML, )).
-Applying advanced DL techniques to detect and fix Android performance issues is promising, although significant challenges remain in collecting sufficient performance issue dataset for model training.
+In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ; pan2024don, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e., code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Wang2022TestDrivenML, ; zhang2025deployability, )).
+Applying advanced DL techniques to detect and fix Android performance issues is promising, although significant challenges remain in collecting a sufficient performance issue dataset for model training.
 The emergence of large language models (LLMs) also presents new possibilities.
-LLMs, with their in-context learning capabilities, could address the limitations of traditional DL approaches by enabling models to detect and resolve performance issues with fewer examples and greater adaptability (Hou2023LargeLM, ; Liao2023A3CodGenAR, ; Si2024AST, ; Fan2023LargeLM, ).
+LLMs, with their in-context learning capabilities, could address the limitations of traditional DL approaches by enabling models to detect and resolve performance issues with fewer examples and greater adaptability (Hou2023LargeLM, ; Liao2023A3CodGenAR, ; Si2024AST, ; Fan2023LargeLM, ; pan2024large, ; han2024chase, ; Huang2022SEFK, ; Su2024EnhancingET, ).
 This opens an exciting direction for future research in Android performance analysis, promising to bridge the current gaps in existing techniques.
 4.3. Constructing Benchmark Dataset is an Urgent Need
-Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 24 (33.33%) of the 72 factors are covered by existing datasets.
-Even among the 24 factors, the data distribution is highly imbalanced, as the smallest dataset contains only five apps, while the largest has 324.
-In addition, the number of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools.
+Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 24 (29.27%) of the 82 factors are covered by existing datasets.
+Even among the 24 factors, the data distribution is highly imbalanced, as the smallest dataset contains only five apps, while the largest has 1,027.
+In addition, the amount of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools.
 Another critical issue is reproducibility.
@@
 5. Threats to Validity
-Our study has several factors that could affect the validity of our findings in terms of large-scale exploration and literature review.
-For real-world data collection, the data filtering process may miss some relevant information.
+Our study is subject to several potential threats to validity, which we categorize into internal and external aspects.
+5.1. Internal Validity
+This concerns the soundness of our data collection and literature review processes.
+For real-world data filtering, the process may miss some relevant information.
 This is because the sentiment classification model (hartmann2023, ) used to filter user reviews may not guarantee 100% accuracy in identifying negative reviews.
@@
 To minimize the risk of missing relevant work, we conducted backward and forward snowballing and cross-checked our results.
+5.2. External Validity
+This concerns the generalizability of our findings beyond the studied datasets and contexts.
+Regarding the representativeness of data sources, our real-world data was derived from open-source and community-driven platforms such as GitHub, Stack Overflow, and Google Play.
+While these platforms reflect commonly reported issues from general developers and users, they may not capture the performance concerns of proprietary or enterprise-level applications developed by organizations such as Google or Meta.
+Furthermore, Stack Overflow discussions may be biased toward novice or intermediate developers, as experienced professionals may participate less frequently.
+As for the research scope, this study focuses exclusively on performance issues in Android mobile applications.
+Other Android-based platforms, such as Android TV, Android Auto, or IoT systems, may exhibit different performance behaviors due to variations in hardware, user interactions, and platform-specific constraints.
+These platforms lie beyond the scope of this work and represent promising avenues for future research.
+Another external threat lies in the lack of user-centric validation.
+While our developer survey provides evidence that the taxonomy reflects practical development concerns, we did not conduct a complementary study with end users.
+Consequently, our taxonomy may not fully capture performance issues that users subjectively perceive as most critical.
+Future work could incorporate user-centric validation, for example through controlled user studies or large-scale questionnaires, to further strengthen the external validity of our findings.
+6. Related Work
+To investigate Android performance issues, numerous studies have been conducted, which can be broadly categorized into two directions: (1) empirical studies based on real-world data sources, and (2) literature reviews that summarize the current state of performance-related research.
+6.1. Empirical Studies on Android Performance Issues
+Numerous studies have explored Android performance issues using real-world sources (Kumari2024AnES, ; Noor2022EndUP, ; Das2016AQA, ; Vsquez2015HowDD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Habchi2019TheRO, ; Habchi2020AndroidCS, ; MazueraRozo2020InvestigatingTA, ; Prestat2022AnES, ).
+These studies typically focus on a single type of data source and adopt either the user or developer perspective.
+For example, Noor et al.
+(Noor2022EndUP, ) focused on the user perspective by examining 368,704 low-rated Google Play reviews from 55 apps.
+They identified eight categories of performance issues reported by users, including responsiveness delays, graphical glitches, and image rendering problems.
+Kumari et al.
+(Kumari2024AnES, ) analyzed 385 Stack Overflow posts to understand performance issues discussed by developers.
+They proposed two taxonomies: one categorizing eight performance issue types (e.g., UI/UX, graphics, and development practices), and another describing their associated causes.
+Similarly, Das et al.
+(Das2016AQA, ), Vásquez et al.
+(Vsquez2015HowDD, ) and Liu et al.
+(Liu2014CharacterizingAD, ) analyzed GitHub issue reports and commit histories to categorize performance problems that developers frequently encounter.
+While these studies effectively identify performance issues and root causes from user and developer perspectives, they do not assess whether these real-world concerns are adequately addressed by academic research.
+Moreover, although some works propose taxonomies, the classification dimensions and terminology vary significantly, hindering comparison and integration across studies.
+In contrast, our study integrates data from multiple sources, including Google Play (user perspective), Stack Overflow and GitHub (developer perspective), and academic literature (researcher perspective).
+Building on this integration, we adopt a hierarchical consequence–cause structure that explicitly links observable performance issues to their underlying root causes, enabling direct mapping to existing tools and datasets for comprehensive gap analysis.
+6.2. Literature Reviews on Android Performance Issues
+To assess the current state of Android performance analysis, several systematic literature reviews have summarized the performance-related issues studied in academic research (Wu2023ASL, ; Fawad2024AndroidSC, ; Hort2021ASO, ).
+For example, Wu et al.
+(Wu2023ASL, ) reviewed 35 studies on Android code smells and identified five major categories, including performance-related smells such as energy inefficiency and memory usage.
+Similarly, Fawad et al.
+(Fawad2024AndroidSC, ) analyzed 79 papers and extracted 237 types of Android code smells, covering concerns such as durable wakelocks, member-ignoring methods, and inefficient database access, many of which are related to performance.
+In addition, Hort et al.
+(Hort2021ASO, ) reviewed 156 papers on non-functional performance optimization in Android apps.
+These studies focused on four key performance characteristics (e.g., responsiveness, launch time, memory consumption, and energy usage), and summarized common optimization strategies such as anti-pattern removal, refactoring, and preloading.
+While these reviews offer valuable insights into performance issues and proposed solutions, they do not evaluate whether academic efforts effectively address the most critical challenges in real-world development.
+They also overlook the extent to which existing tools and datasets support the detection and evaluation of such issues.
+Consequently, the actual alignment between academic research and practical concerns remains unclear.
+In contrast, our study first identifies the full spectrum of real-world performance issues.
+We then analyze how many of these issues are examined in the literature, which tools are available to address them, and whether corresponding datasets exist to support empirical evaluation.
+This enables a comprehensive understanding of research coverage and highlights gaps to guide future work.
 7. Conclusion
 In this paper, we conducted a comparative study to examine real-world Android performance issues encountered by users and developers, and to evaluate the extent to which current academic research addresses these practical challenges.
-By analyzing large-scale discussions from three popular crowd-sourced platforms on Android performance issues, alongside a review of 66 high-quality research papers, we identified several gaps between real-world challenges and the focus of academic research.
-Our main findings are as follows: (1) A taxonomy categorizes seven types of performance issues and 72 contributing factors.
+By analyzing large-scale discussions from three popular crowd-sourced platforms on Android performance issues, alongside a review of 85 high-quality research papers, we identified several gaps between real-world challenges and the focus of academic research.
+Our main findings are as follows: (1) A taxonomy categorizes seven types of performance issues and 82 contributing factors.
 (2) Six common code patterns of these performance issues.
@@
 (5) The majority of existing tools primarily utilize static analysis, dynamic analysis, and deep learning, with most tools detecting issues rather than fixing them.
-(6) Existing available tools target only 23.61% of identified factors.
-(7) Existing available datasets cover only 33.33% factors and lack sufficient data to support effective tool evaluation.
+(6) Existing available tools target only 36.59% of identified factors.
+(7) Existing available datasets cover only 29.27% of factors and lack sufficient data to support effective tool evaluation.
 Finally, we proposed future research directions to bridge the identified gaps.